Jefferson Lab HALL A Logbook for April 11 - Full Listing

  • Main INDEX

    Saturday

  • 357671 04/30/11 23:41 Fatiha Coda crash
  • 357670 04/30/11 23:38 a-onl Start_Run_4036,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=all prescalers = 1,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357669 04/30/11 22:58 a-onl Start_Run_4035,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=all prescalers = 1,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357668 04/30/11 22:42 a-onl End_of_Run_4034
  • 357667 04/30/11 22:12 yez R-HRS: ADC of Shower#66 is getting bad
  • 357666 04/30/11 21:54 Fatiha Spot++ for run 4030
  • 357665 04/30/11 21:46 a-onl Start_Run_4034,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=all prescalers = 1,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357664 04/30/11 21:46 a-onl End_of_Run_4033
  • 357663 04/30/11 21:08 Fatiha Beam Current StripChart
  • 357662 04/30/11 21:07 Fatiha Hall A General tools
  • 357661 04/30/11 21:04 auto_snap Radiation and Ozone snapshot
  • 357660 04/30/11 21:04 auto_snap Miscellaneous snapshot
  • 357659 04/30/11 21:04 auto_snap Hall A Tools snapshot
  • 357658 04/30/11 21:04 auto_snap Compton Polarimeter snapshot
  • 357657 04/30/11 21:04 auto_snap BigBox Magnet snapshot
  • 357656 04/30/11 21:03 auto_snap Beam Position and Current snapshot
  • 357655 04/30/11 20:38 a-onl Start_Run_4033,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=all prescalers = 1,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357654 04/30/11 20:37 a-onl End_of_Run_4032
  • 357653 04/30/11 19:39 a-onl Start_Run_4032,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=all prescalers = 1,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357652 04/30/11 19:38 a-onl End_of_Run_4031
  • 357651 04/30/11 18:34 a-onl Start_Run_4031,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=all prescalers = 1,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357650 04/30/11 18:33 a-onl End_of_Run_4030
  • 357649 04/30/11 18:08 efuchey Target screenshots
  • 357648 04/30/11 17:26 a-onl Start_Run_4030,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=all prescalers = 1,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357647 04/30/11 17:25 a-onl End_of_Run_4029
  • 357646 04/30/11 16:19 a-onl Start_Run_4029,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=all prescalers = 1,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357645 04/30/11 16:18 a-onl End_of_Run_4028
  • 357644 04/30/11 16:04 cisbani Day Shift Summary
  • 357643 04/30/11 16:02 adaq Beam Time Accounting
  • 357642 04/30/11 15:56 xqiu target status
  • 357641 04/30/11 15:16 a-onl Start_Run_4028,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=all prescalers = 1,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357640 04/30/11 15:15 a-onl End_of_Run_4027
  • 357639 04/30/11 14:18 a-onl Start_Run_4027,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=all prescalers = 1,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357638 04/30/11 14:16 a-onl End_of_Run_4026
  • 357637 04/30/11 13:20 cisbani 2nd bunch of runs with 3He-20cm / 120 uA (kin 6.5)
  • 357636 04/30/11 13:18 a-onl Start_Run_4026,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=all prescalers = 1,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357635 04/30/11 13:16 a-onl End_of_Run_4025
  • 357634 04/30/11 13:09 xqiu Moved target to He3 20cm.
  • 357633 04/30/11 13:07 a-onl Start_Run_4025,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text= - ps6=ps7=1, ps2 = 1, ps4 = 2.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357632 04/30/11 13:03 auto_snap Radiation and Ozone snapshot
  • 357631 04/30/11 13:03 auto_snap Miscellaneous snapshot
  • 357630 04/30/11 13:03 auto_snap Hall A Tools snapshot
  • 357629 04/30/11 13:03 auto_snap Compton Polarimeter snapshot
  • 357628 04/30/11 13:03 auto_snap BigBox Magnet snapshot
  • 357627 04/30/11 13:03 auto_snap Beam Position and Current snapshot
  • 357626 04/30/11 12:43 a-onl End_of_Run_4024
  • 357625 04/30/11 12:33 cisbani HallA Tools
  • 357624 04/30/11 12:26 adaq spot++ (spot_L)
  • 357623 04/30/11 11:30 a-onl Start_Run_4024,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Dummy 20cm - ps6=ps7=1, ps2 = 1, ps4 = 2.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357622 04/30/11 11:29 a-onl End_of_Run_4023
  • 357621 04/30/11 10:27 a-onl Start_Run_4023,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Dummy 20cm - ps6=ps7=1, ps2 = 1, ps4 = 2.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357620 04/30/11 10:26 a-onl End_of_Run_4022
  • 357619 04/30/11 09:17 a-onl Start_Run_4022,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Dummy 20cm - ps6=ps7=1, ps2 = 1, ps4 = 2.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357618 04/30/11 09:16 a-onl End_of_Run_4021
  • 357617 04/30/11 09:00 cisbani Low Methane Gas Alarm
  • 357616 04/30/11 08:49 cisbani Half wave plate is IN now
  • 357615 04/30/11 08:16 a-onl Start_Run_4021,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Dummy 20cm - ps6=ps7=1, ps2 = 1, ps4 = 2.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357614 04/30/11 08:14 a-onl End_of_Run_4020
  • 357613 04/30/11 08:02 efuchey 4/30/2011 Owl shift summary
  • 357612 04/30/11 08:02 adaq Beam Time Accounting
  • 357611 04/30/11 07:19 a-onl Start_Run_4020,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Dummy 20cm - ps6=ps7=1, ps2 = 1, ps4 = 2.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357610 04/30/11 07:18 a-onl End_of_Run_4019
  • 357609 04/30/11 06:47 xbai raster & spot++ for dummy 20cm
  • 357608 04/30/11 06:25 a-onl Start_Run_4019,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Dummy 20cm - ps6=ps7=1, ps2 = 1, ps4 = 2.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357607 04/30/11 06:24 a-onl End_of_Run_4018
  • 357606 04/30/11 05:31 xbai prescaled rate for the new target dummy 20 cm
  • 357605 04/30/11 05:25 a-onl Start_Run_4018,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Dummy 20cm - ps6=ps7=1, ps2 = 1, ps4 = 2.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357604 04/30/11 05:20 efuchey Target changed to dummy 20 cm
  • 357603 04/30/11 05:16 a-onl End_of_Run_4017
  • 357602 04/30/11 05:03 auto_snap Radiation and Ozone snapshot
  • 357601 04/30/11 05:03 auto_snap Miscellaneous snapshot
  • 357600 04/30/11 05:03 auto_snap Hall A Tools snapshot
  • 357599 04/30/11 05:03 auto_snap Compton Polarimeter snapshot
  • 357598 04/30/11 05:03 auto_snap BigBox Magnet snapshot
  • 357597 04/30/11 05:03 auto_snap Beam Position and Current snapshot
  • 357596 04/30/11 04:46 a-onl Start_Run_4017,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production C - ps6=ps7=2, ps2 = 2, ps4 = 3.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357595 04/30/11 04:45 a-onl End_of_Run_4016
  • 357594 04/30/11 04:35 efuchey Beamline HV crate
  • 357593 04/30/11 03:59 a-onl Start_Run_4016,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production C - ps6=ps7=2, ps2 = 2, ps4 = 3.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357592 04/30/11 03:58 a-onl End_of_Run_4015
  • 357591 04/30/11 03:35 efuchey Hall A tools screenshot
  • 357590 04/30/11 03:12 a-onl Start_Run_4015,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production C - ps6=ps7=2, ps2 = 2, ps4 = 3.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357589 04/30/11 03:09 a-onl End_of_Run_4014
  • 357588 04/30/11 02:51 xbai prescale after coda restart
  • 357587 04/30/11 02:26 a-onl Start_Run_4014,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production C - ps6=ps7=2, ps2 = 2, ps4 = 3.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357586 04/30/11 02:22 efuchey CODA problem
  • 357585 04/30/11 01:41 xbai prescales for carbon 5mm
  • 357584 04/30/11 01:36 xbai target changed
  • 357583 04/30/11 01:35 a-onl Start_Run_4013,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production C - ps6=ps7=2, ps2 = 2, ps4 = 3.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357582 04/30/11 01:26 efuchey Target move to Carbon 5mm
  • 357581 04/30/11 01:18 a-onl End_of_Run_4012
  • 357580 04/30/11 01:11 xbai raster spot++
  • 357579 04/30/11 00:33 a-onl Start_Run_4012,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357578 04/30/11 00:32 a-onl End_of_Run_4011
  • 357577 04/30/11 00:19 xbai target status snapshot
  • 357576 04/30/11 00:14 cisbani Swing Shift Summary
  • 357575 04/30/11 00:07 adaq Beam Time Accounting
  • Friday

  • 357574 04/29/11 23:45 a-onl Start_Run_4011,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357573 04/29/11 23:44 a-onl End_of_Run_4010
  • 357572 04/29/11 22:57 a-onl Start_Run_4010,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357571 04/29/11 22:56 a-onl End_of_Run_4009
  • 357570 04/29/11 22:04 Ramesh Half wave plate is OUT now
  • 357569 04/29/11 21:55 a-onl Start_Run_4009,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357568 04/29/11 21:54 a-onl End_of_Run_4008
  • 357567 04/29/11 21:03 auto_snap Radiation and Ozone snapshot
  • 357566 04/29/11 21:03 auto_snap Miscellaneous snapshot
  • 357565 04/29/11 21:03 auto_snap Hall A Tools snapshot
  • 357564 04/29/11 21:03 auto_snap Compton Polarimeter snapshot
  • 357563 04/29/11 21:03 auto_snap BigBox Magnet snapshot
  • 357562 04/29/11 21:02 auto_snap Beam Position and Current snapshot
  • 357561 04/29/11 20:43 a-onl Start_Run_4008,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357560 04/29/11 20:42 a-onl End_of_Run_4007
  • 357559 04/29/11 19:38 a-onl Start_Run_4007,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357558 04/29/11 19:37 a-onl End_of_Run_4006
  • 357557 04/29/11 19:10 cisbani Fake alarm in HV VDC
  • 357556 04/29/11 18:43 a-onl Start_Run_4006,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357555 04/29/11 18:41 a-onl End_of_Run_4005
  • 357554 04/29/11 18:21 Kalyan Target screens
  • 357553 04/29/11 18:20 cisbani HallA Tools screen
  • 357552 04/29/11 18:18 cisbani spot++ (spot_L)
  • 357551 04/29/11 17:49 a-onl Start_Run_4005,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357550 04/29/11 17:48 a-onl End_of_Run_4004
  • 357549 04/29/11 16:53 a-onl Start_Run_4004,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357548 04/29/11 16:52 a-onl End_of_Run_4003
  • 357547 04/29/11 16:04 adaq Beam Time Accounting
  • 357546 04/29/11 15:59 saw Day shift summary
  • 357545 04/29/11 15:39 saw Methane low pressure alarm
  • 357544 04/29/11 15:21 a-onl End_of_Run_4002
  • 357543 04/29/11 15:15 yez Set value of RHRS HV L15.7 to zero
  • 357542 04/29/11 14:44 a-onl Start_Run_4002,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357541 04/29/11 14:43 a-onl End_of_Run_4001
  • 357540 04/29/11 13:08 a-onl Start_Run_4001,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357539 04/29/11 13:07 a-onl End_of_Run_4000
  • 357538 04/29/11 13:02 auto_snap Radiation and Ozone snapshot
  • 357537 04/29/11 13:02 auto_snap Miscellaneous snapshot
  • 357536 04/29/11 13:02 auto_snap Hall A Tools snapshot
  • 357535 04/29/11 13:02 auto_snap Compton Polarimeter snapshot
  • 357534 04/29/11 13:02 auto_snap BigBox Magnet snapshot
  • 357533 04/29/11 13:02 auto_snap Beam Position and Current snapshot
  • 357532 04/29/11 12:06 yez RHRS HVs L15.7 is used for Lumi signal and is bad, so please don't turn it on
  • 357531 04/29/11 11:57 saw Beam interrupted for Faraday cup for Hall C setup
  • 357530 04/29/11 11:49 saw Spot++ screenshots after spot move
  • 357529 04/29/11 11:48 a-onl Start_Run_4000,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357528 04/29/11 11:46 a-onl End_of_Run_3999
  • 357527 04/29/11 11:44 a-onl Start_Run_3999,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357526 04/29/11 10:58 saw Beam down for spot move
  • 357525 04/29/11 10:58 a-onl End_of_Run_3998
  • 357524 04/29/11 10:32 saw IHWP changed from OUT to IN during run 3998
  • 357523 04/29/11 10:09 a-onl Start_Run_3998,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357522 04/29/11 10:07 a-onl End_of_Run_3997
  • 357521 04/29/11 09:22 yez Owl shift Runs summary:
  • 357520 04/29/11 09:11 yez Cumulative good electrons for owl shift runs were not calculated correctly. Update soon
  • 357519 04/29/11 09:10 saw RHMS S15.7 HV is supposed to be off
  • 357518 04/29/11 09:07 saw 3996 Crashed. 3995 was OK.
  • 357517 04/29/11 08:54 saw Run 3995 crashed. ROC10
  • 357516 04/29/11 08:53 a-onl Start_Run_3997,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357515 04/29/11 08:15 a-onl Start_Run_3996,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357514 04/29/11 08:14 a-onl End_of_Run_3995
  • 357513 04/29/11 08:04 C. Gu Target screenshots
  • 357512 04/29/11 08:02 adaq Beam Time Accounting
  • 357511 04/29/11 08:01 A. Camsonne Owl Shift Summary
  • 357510 04/29/11 06:58 a-onl Start_Run_3995,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357509 04/29/11 06:57 a-onl End_of_Run_3994
  • 357508 04/29/11 05:40 a-onl Start_Run_3994,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357507 04/29/11 05:40 a-onl End_of_Run_3993
  • 357506 04/29/11 05:02 auto_snap Radiation and Ozone snapshot
  • 357505 04/29/11 05:02 auto_snap Miscellaneous snapshot
  • 357504 04/29/11 05:02 auto_snap Hall A Tools snapshot
  • 357503 04/29/11 05:02 auto_snap Compton Polarimeter snapshot
  • 357502 04/29/11 05:02 auto_snap BigBox Magnet snapshot
  • 357501 04/29/11 05:02 auto_snap Beam Position and Current snapshot
  • 357500 04/29/11 04:39 gueye Note on runs 3991 & 3992
  • 357499 04/29/11 04:30 a-onl Start_Run_3993,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357498 04/29/11 04:28 a-onl End_of_Run_3992
  • 357497 04/29/11 04:28 a-onl Start_Run_3992,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357496 04/29/11 04:27 a-onl End_of_Run_3991
  • 357495 04/29/11 04:12 a-onl Start_Run_3991,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357494 04/29/11 04:12 a-onl End_of_Run_3990
  • 357493 04/29/11 03:55 gueye Pps6 & ps7 changed at run 3988
  • 357492 04/29/11 03:54 P. Aguilera Screenshots target
  • 357491 04/29/11 03:02 a-onl Start_Run_3990,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357490 04/29/11 03:01 a-onl End_of_Run_3989
  • 357489 04/29/11 01:57 a-onl Start_Run_3989,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357488 04/29/11 01:56 a-onl End_of_Run_3988
  • 357487 04/29/11 01:56 P. Aguilera tools screen
  • 357486 04/29/11 01:52 P. Aguilera screenshots spot++
  • 357485 04/29/11 00:48 a-onl Start_Run_3988,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Production He3 - ps6=ps7=2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357484 04/29/11 00:47 a-onl End_of_Run_3987
  • 357483 04/29/11 00:03 adaq Beam Time Accounting
  • 357482 04/29/11 00:01 johna Swing shift summary
  • Thursday

  • 357481 04/28/11 23:49 cisbani run 3986 terminated abnormally (EB2 disconnected)
  • 357480 04/28/11 23:47 a-onl Start_Run_3987,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=coda restarted,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357479 04/28/11 23:06 a-onl Start_Run_3986,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357478 04/28/11 23:05 a-onl End_of_Run_3985
  • 357477 04/28/11 22:08 cisbani DAQ screen
  • 357476 04/28/11 22:06 a-onl Start_Run_3985,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357475 04/28/11 22:05 a-onl End_of_Run_3984
  • 357474 04/28/11 21:07 a-onl Start_Run_3984,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357473 04/28/11 21:02 auto_snap Radiation and Ozone snapshot
  • 357472 04/28/11 21:02 auto_snap Miscellaneous snapshot
  • 357471 04/28/11 21:02 auto_snap Hall A Tools snapshot
  • 357470 04/28/11 21:02 auto_snap Compton Polarimeter snapshot
  • 357469 04/28/11 21:02 auto_snap BigBox Magnet snapshot
  • 357468 04/28/11 21:01 auto_snap Beam Position and Current snapshot
  • 357467 04/28/11 20:57 a-onl End_of_Run_3983
  • 357466 04/28/11 20:38 mhuang Target screens
  • 357465 04/28/11 20:06 johna Start data taking, 90uA
  • 357464 04/28/11 20:00 a-onl Start_Run_3983,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357463 04/28/11 19:59 a-onl End_of_Run_3982
  • 357462 04/28/11 19:46 a-onl Start_Run_3982,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357461 04/28/11 19:43 a-onl End_of_Run_3981
  • 357460 04/28/11 19:41 johna Beam recovery update
  • 357459 04/28/11 19:41 a-onl Start_Run_3981,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357458 04/28/11 19:38 a-onl End_of_Run_3980
  • 357457 04/28/11 19:33 a-onl Start_Run_3980,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=current not yet stable,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357456 04/28/11 19:32 a-onl End_of_Run_3979
  • 357455 04/28/11 19:28 a-onl Start_Run_3979,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=current is rumping up,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357454 04/28/11 18:54 a-onl End_of_Run_3978
  • 357453 04/28/11 18:47 a-onl End_of_Run_3977
  • 357452 04/28/11 18:47 a-onl Start_Run_3978,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=production (first after beam study),bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357451 04/28/11 18:42 johna Problems establishing high current
  • 357450 04/28/11 18:18 a-onl Start_Run_3977,Run_type=Spot++,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=1,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=raster check,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357449 04/28/11 18:17 mhuang Target moved to He3, 20cm
  • 357448 04/28/11 18:13 johna Beam startup
  • 357447 04/28/11 18:07 A. Camsonne Buffering - Busy signal issue
  • 357446 04/28/11 17:57 cisbani Spot_L
  • 357445 04/28/11 17:55 a-onl End_of_Run_3976
  • 357444 04/28/11 17:53 a-onl Start_Run_3976,Run_type=Spot++,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=raster check,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357443 04/28/11 17:52 a-onl End_of_Run_3975
  • 357442 04/28/11 17:45 johna Sync problem fixed, we believe
  • 357441 04/28/11 17:42 a-onl Start_Run_3975,Run_type=Cosmics,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=cosmic, no beam - test epicRunstart - check sync signal,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357440 04/28/11 17:41 mhuang Target moved to BeO for beam recovery
  • 357439 04/28/11 17:41 a-onl End_of_Run_3974
  • 357438 04/28/11 17:40 a-onl Start_Run_3974,Run_type=Cosmics,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=cosmic, no beam - test epicRunstart - check sync signal,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357437 04/28/11 17:40 johna Hall going to beam permit
  • 357436 04/28/11 17:27 a-onl End_of_Run_3973
  • 357435 04/28/11 17:25 a-onl Start_Run_3973,Run_type=Cosmics,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=cosmic, no beam - test epicRunstart - check sync signal,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357434 04/28/11 17:25 a-onl End_of_Run_3972
  • 357433 04/28/11 17:22 a-onl Start_Run_3972,Run_type=Cosmics,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=cosmic, no beam - test epicRunstart - check sync signal,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357432 04/28/11 17:21 a-onl End_of_Run_3971
  • 357431 04/28/11 17:16 johna MCC calls to say beam is ready
  • 357430 04/28/11 17:16 a-onl Start_Run_3971,Run_type=Cosmics,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=cosmic, no beam - test epicRunstart - check sync signal,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357429 04/28/11 17:15 a-onl End_of_Run_3970
  • 357428 04/28/11 17:10 a-onl Start_Run_3970,Run_type=Cosmics,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=cosmic, no beam - test epicRunstart - check sync signal,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357427 04/28/11 17:10 a-onl End_of_Run_3969
  • 357426 04/28/11 17:06 a-onl End_of_Run_3968
  • 357425 04/28/11 17:02 a-onl Start_Run_3968,Run_type=Cosmics,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=cosmic, no beam - test epicRunstart - check sync signal,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357424 04/28/11 17:02 a-onl End_of_Run_3967
  • 357423 04/28/11 16:58 a-onl Start_Run_3967,Run_type=Cosmics,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=cosmic, no beam - test epicRunstart - check sync signal,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357422 04/28/11 16:46 a-onl End_of_Run_3965
  • 357421 04/28/11 16:44 a-onl Start_Run_3965,Run_type=Cosmics,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=cosmic, no beam - test epicRunstart - check sync signal,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357420 04/28/11 16:34 a-onl End_of_Run_3962
  • 357419 04/28/11 16:32 a-onl Start_Run_3962,Run_type=Cosmics,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=cosmic, no beam - test epicRunstart - check sync signal,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357418 04/28/11 16:26 benesch Wien has been flipped
  • 357417 04/28/11 16:21 jpchen Day shift summary
  • 357416 04/28/11 16:21 adaq Beam Time Accounting
  • 357415 04/28/11 16:19 mhuang Target screens
  • 357414 04/28/11 15:12 Benesch Wien flip likely will occur this afternoon
  • 357413 04/28/11 14:20 a-onl Start_Run_3960,Run_type=Cosmics,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=cosmic, no beam - test epicRunstart - check sync signal,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357412 04/28/11 14:19 a-onl End_of_Run_3959
  • 357411 04/28/11 14:15 Jack (Techs) Cycled power on Q1 power supplies
  • 357410 04/28/11 14:07 a-onl Start_Run_3959,Run_type=Cosmics,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=cosmic, no beam - test epicRunstart - check sync signal,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357409 04/28/11 14:06 a-onl End_of_Run_3958
  • 357408 04/28/11 13:47 A. Camsonne Sync ADC left arm replugged
  • 357407 04/28/11 13:47 a-onl Start_Run_3958,Run_type=Cosmics,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=cosmic, no beam - test epicRunstart - check sync signal,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357406 04/28/11 13:46 jpchen Beam study, ~ 15:30
  • 357405 04/28/11 13:40 a-onl End_of_Run_3957
  • 357404 04/28/11 13:01 auto_snap Radiation and Ozone snapshot
  • 357403 04/28/11 13:01 auto_snap Miscellaneous snapshot
  • 357402 04/28/11 13:01 auto_snap Hall A Tools snapshot
  • 357401 04/28/11 13:01 auto_snap Compton Polarimeter snapshot
  • 357400 04/28/11 13:01 auto_snap BigBox Magnet snapshot
  • 357399 04/28/11 13:01 auto_snap Beam Position and Current snapshot
  • 357398 04/28/11 10:34 a-onl Start_Run_3957,Run_type=Cosmics,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=cosmic, no beam - test epicRunstart,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357397 04/28/11 10:28 A. Camsonne Added Left ARM HV to epicsRunStart script
  • 357396 04/28/11 10:27 a-onl Start_Run_3957,Run_type=Cosmics,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=cosmic, no beam - test epicRunstart,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357395 04/28/11 10:26 a-onl Start_Run_3957,Run_type=Cosmics,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=cosmic, no beam - test epicRunstart,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357394 04/28/11 08:24 a-onl Start_Run_3957,Run_type=Cosmics,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=cosmic, no beam,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357393 04/28/11 08:03 adaq Beam Time Accounting
  • 357392 04/28/11 07:59 yez Owl Shift Summary
  • 357391 04/28/11 07:49 yez Asked to put hall into control access, but need to wait for people coming to work
  • 357390 04/28/11 07:14 yez target screen - no a single alarm during the whole shift!
  • 357389 04/28/11 07:02 yez Beam is off for beam study
  • 357388 04/28/11 07:02 a-onl End_of_Run_3956
  • 357387 04/28/11 06:43 yez beam will be off at 7am
  • 357386 04/28/11 05:43 a-onl Start_Run_3956,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Dummy 20cm, 40 mu=icro amps ps = 1,2,1,2,0,2,2,400,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357385 04/28/11 05:42 a-onl End_of_Run_3955
  • 357384 04/28/11 05:01 auto_snap Radiation and Ozone snapshot
  • 357383 04/28/11 05:01 auto_snap Miscellaneous snapshot
  • 357382 04/28/11 05:01 auto_snap Hall A Tools snapshot
  • 357381 04/28/11 05:01 auto_snap Compton Polarimeter snapshot
  • 357380 04/28/11 05:01 auto_snap BigBox Magnet snapshot
  • 357379 04/28/11 05:01 auto_snap Beam Position and Current snapshot
  • 357378 04/28/11 04:52 xbai Raster Snapshot
  • 357377 04/28/11 04:42 a-onl Start_Run_3955,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Dummy 20cm, 40 mu=icro amps ps = 1,2,1,2,0,2,2,400,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357376 04/28/11 04:42 yez Beam is back
  • 357375 04/28/11 04:26 yez Beam Stop. MCC needs to cycle one of their power supply
  • 357374 04/28/11 04:17 a-onl End_of_Run_3954
  • 357373 04/28/11 04:10 a-onl Start_Run_3954,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Dummy 20cm, 40 mu=icro amps ps = 1,2,1,2,0,2,2,400,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357372 04/28/11 04:10 a-onl End_of_Run_3953
  • 357371 04/28/11 03:21 xbai Raster & spot++ snapshot
  • 357370 04/28/11 03:09 a-onl Start_Run_3953,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Dummy 20cm, 40 mu=icro amps ps = 1,2,1,2,0,2,2,400,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357369 04/28/11 03:08 a-onl End_of_Run_3952
  • 357368 04/28/11 02:23 xbai Raster & spot++ check for run 3952 with target dummy 20
  • 357367 04/28/11 02:14 xbai Target change
  • 357366 04/28/11 02:09 a-onl Start_Run_3952,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Thick carbon, 95 mu=icro amps ps = 1,4,1,4,0,3,3,400,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357365 04/28/11 02:07 a-onl End_of_Run_3951
  • 357364 04/28/11 01:21 yez target screen
  • 357363 04/28/11 01:12 xbai Raster spot++
  • 357362 04/28/11 01:02 yez This is Dummy target run
  • 357361 04/28/11 00:51 xbai Prescale & SawDTmon for Dummy
  • 357360 04/28/11 00:49 a-onl Start_Run_3951,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Thick carbon, 95 mu=icro amps ps = 1,4,1,4,0,3,3,400,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357359 04/28/11 00:44 xbai Total good events for carbon
  • 357358 04/28/11 00:44 yez target moved to Dummy 20cm
  • 357357 04/28/11 00:42 a-onl End_of_Run_3950
  • 357356 04/28/11 00:09 adaq Beam Time Accounting
  • 357355 04/28/11 00:08 johna Swing shift summary
  • Wednesday

  • 357354 04/27/11 23:27 a-onl Start_Run_3950,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Thick carbon, 95 mu=icro amps ps = 1,4,1,4,0,3,3,400,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357353 04/27/11 23:25 a-onl End_of_Run_3949
  • 357352 04/27/11 22:28 Donal Spot_L
  • 357351 04/27/11 22:25 a-onl Start_Run_3949,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=Thick carbon, 95 mu=icro amps ps = 1,4,1,4,0,3,3,400,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357350 04/27/11 22:17 Donal 3He Events
  • 357349 04/27/11 22:14 wdconinc Moving target to Carbon 5mm
  • 357348 04/27/11 22:14 a-onl End_of_Run_3948
  • 357347 04/27/11 21:51 Donal Spot_L
  • 357346 04/27/11 21:01 auto_snap Radiation and Ozone snapshot
  • 357345 04/27/11 21:01 auto_snap Miscellaneous snapshot
  • 357344 04/27/11 21:01 auto_snap Hall A Tools snapshot
  • 357343 04/27/11 21:01 auto_snap Compton Polarimeter snapshot
  • 357342 04/27/11 21:01 auto_snap BigBox Magnet snapshot
  • 357341 04/27/11 21:01 auto_snap Beam Position and Current snapshot
  • 357340 04/27/11 20:57 a-onl Start_Run_3948,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK unbuffered,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357339 04/27/11 20:56 a-onl End_of_Run_3947
  • 357338 04/27/11 19:55 a-onl Start_Run_3947,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK unbuffered,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357337 04/27/11 19:54 a-onl End_of_Run_3946
  • 357336 04/27/11 18:53 a-onl Start_Run_3946,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK unbuffered,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357335 04/27/11 18:52 a-onl End_of_Run_3945
  • 357334 04/27/11 18:49 Donal spot_L
  • 357333 04/27/11 17:49 a-onl Start_Run_3945,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK unbuffered,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357332 04/27/11 17:48 a-onl End_of_Run_3944
  • 357331 04/27/11 17:43 johna Hall C *actually* changes the half wave plate
  • 357330 04/27/11 17:08 wdconinc Target screen shots
  • 357329 04/27/11 16:34 johna Hall C changing half wave plate
  • 357328 04/27/11 16:29 a-onl Start_Run_3944,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK unbuffered,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357327 04/27/11 16:27 a-onl End_of_Run_3943
  • 357326 04/27/11 16:04 adaq Beam Time Accounting
  • 357325 04/27/11 16:04 Ole Hansen Day Shift Summary
  • 357324 04/27/11 15:27 a-onl Start_Run_3943,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK unbuffered,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357323 04/27/11 15:26 a-onl End_of_Run_3942
  • 357322 04/27/11 15:12 M Defurne screenshots
  • 357321 04/27/11 14:28 a-onl Start_Run_3942,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK unbuffered,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357320 04/27/11 14:27 a-onl End_of_Run_3941
  • 357319 04/27/11 13:44 Rondon spot++ run 3941
  • 357318 04/27/11 13:27 a-onl Start_Run_3941,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK unbuffered,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357317 04/27/11 13:26 a-onl End_of_Run_3940
  • 357316 04/27/11 13:00 auto_snap Radiation and Ozone snapshot
  • 357315 04/27/11 13:00 auto_snap Miscellaneous snapshot
  • 357314 04/27/11 13:00 auto_snap Hall A Tools snapshot
  • 357313 04/27/11 13:00 auto_snap Compton Polarimeter snapshot
  • 357312 04/27/11 13:00 auto_snap BigBox Magnet snapshot
  • 357311 04/27/11 13:00 auto_snap Beam Position and Current snapshot
  • 357310 04/27/11 12:26 a-onl Start_Run_3940,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK unbuffered ,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357309 04/27/11 12:25 a-onl End_of_Run_3939
  • 357308 04/27/11 11:43 M Defurne screenshots
  • 357307 04/27/11 11:25 a-onl Start_Run_3939,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK unbuffered HWP changed,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357306 04/27/11 11:22 Jack (Techs) reboot of iocha12 (beamline miscellaneous Slow Controls)
  • 357305 04/27/11 11:21 a-onl End_of_Run_3938
  • 357304 04/27/11 11:21 Ole Hansen Hall C changes half-wave plate
  • 357303 04/27/11 11:17 Rondon full replay run3937 has synch warning
  • 357302 04/27/11 11:15 a-onl Start_Run_3938,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK unbuffered,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357301 04/27/11 11:13 a-onl End_of_Run_3937
  • 357300 04/27/11 11:02 Jack (Techs) Q1 and BigBite power supply readbacks
  • 357299 04/27/11 10:22 Rondon Full replay run 3936 shows ROCs still not in sync
  • 357298 04/27/11 10:15 a-onl Start_Run_3937,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK unbuffered,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357297 04/27/11 10:13 a-onl End_of_Run_3936
  • 357296 04/27/11 09:29 M Defurne screenshots
  • 357295 04/27/11 09:23 Rondon ROC sync OK again, run 3936
  • 357294 04/27/11 09:12 a-onl Start_Run_3936,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK unbuffered,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357293 04/27/11 09:11 a-onl End_of_Run_3935
  • 357292 04/27/11 08:27 Ole Hansen RHRS replay ok - RQ1 polarity readback surely bogus
  • 357291 04/27/11 08:23 Ole Hansen RQ1 shows positive polarity!
  • 357290 04/27/11 08:10 a-onl Start_Run_3935,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK unbuffered,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357289 04/27/11 08:10 Ole Hansen Fast feedback was off
  • 357288 04/27/11 08:09 a-onl End_of_Run_3934
  • 357287 04/27/11 08:03 yez Owl shift Summary
  • 357286 04/27/11 08:02 adaq Beam Time Accounting
  • 357285 04/27/11 07:24 yez target screen
  • 357284 04/27/11 06:55 a-onl Start_Run_3934,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357283 04/27/11 06:54 a-onl End_of_Run_3933
  • 357282 04/27/11 05:41 a-onl Start_Run_3933,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357281 04/27/11 05:40 a-onl End_of_Run_3932
  • 357280 04/27/11 05:00 auto_snap Radiation and Ozone snapshot
  • 357279 04/27/11 05:00 auto_snap Miscellaneous snapshot
  • 357278 04/27/11 05:00 auto_snap Hall A Tools snapshot
  • 357277 04/27/11 05:00 auto_snap Compton Polarimeter snapshot
  • 357276 04/27/11 05:00 auto_snap BigBox Magnet snapshot
  • 357275 04/27/11 05:00 auto_snap Beam Position and Current snapshot
  • 357274 04/27/11 04:38 a-onl Start_Run_3932,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357273 04/27/11 04:37 a-onl End_of_Run_3931
  • 357272 04/27/11 03:52 adaq spot++ for run3931
  • 357271 04/27/11 03:49 yez target screen
  • 357270 04/27/11 03:36 yez CODA runs in un-buffer mode since run#3931
  • 357269 04/27/11 03:32 yez See few Out-Of-Sync events when checking full run data
  • 357268 04/27/11 03:22 a-onl Start_Run_3931,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357267 04/27/11 03:17 a-onl End_of_Run_3930
  • 357266 04/27/11 03:06 a-onl Start_Run_3930,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357265 04/27/11 03:05 a-onl End_of_Run_3929
  • 357264 04/27/11 02:00 a-onl Start_Run_3929,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357263 04/27/11 01:59 a-onl End_of_Run_3928
  • 357262 04/27/11 01:06 yez reboot ROC10 at the end of 3927, coda works again
  • 357261 04/27/11 00:59 a-onl Start_Run_3928,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357260 04/27/11 00:55 a-onl End_of_Run_3927
  • 357259 04/27/11 00:32 yez target screen
  • 357258 04/27/11 00:16 adaq Beam Time Accounting
  • 357257 04/27/11 00:15 johna Swing shift summary
  • Tuesday

  • 357256 04/26/11 23:51 a-onl Start_Run_3927,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357255 04/26/11 23:51 a-onl End_of_Run_3926
  • 357254 04/26/11 23:34 johna Hall C is changing state of half wave plate
  • 357253 04/26/11 22:49 a-onl Start_Run_3926,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=118,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357252 04/26/11 22:34 a-onl End_of_Run_3925
  • 357251 04/26/11 21:34 a-onl Start_Run_3925,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=90,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357250 04/26/11 21:33 a-onl End_of_Run_3924
  • 357249 04/26/11 21:00 auto_snap Radiation and Ozone snapshot
  • 357248 04/26/11 21:00 auto_snap Miscellaneous snapshot
  • 357247 04/26/11 21:00 auto_snap Hall A Tools snapshot
  • 357246 04/26/11 21:00 auto_snap Compton Polarimeter snapshot
  • 357245 04/26/11 21:00 auto_snap BigBox Magnet snapshot
  • 357244 04/26/11 21:00 auto_snap Beam Position and Current snapshot
  • 357243 04/26/11 20:31 a-onl Start_Run_3924,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=90,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357242 04/26/11 20:30 a-onl End_of_Run_3923
  • 357241 04/26/11 20:19 johna Running at 105uA while MCC trouble shoots
  • 357240 04/26/11 19:42 a-onl Start_Run_3923,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=90,leftarm_p=2.845,leftarm_theta=28,leftarm_sieve=out,rightarm_p=2.845,rightarm_theta=28,rightarm_sieve=out,comment_text=3He production 120 muA both arms OK,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357239 04/26/11 19:40 Rondon Runs 3917 to 3921 spectrometers at 2.845 GeV/c 28 deg
  • 357238 04/26/11 19:35 a-onl End_of_Run_3922
  • 357237 04/26/11 19:35 a-onl Start_Run_3922,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=90,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=3He production 120 muA Left Arm only - R arm ramping up,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357236 04/26/11 19:34 a-onl End_of_Run_3921
  • 357235 04/26/11 19:26 a-onl Start_Run_3921,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=90,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=3He production 120 muA Left Arm only - R arm ramping up,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357234 04/26/11 19:25 a-onl End_of_Run_3920
  • 357233 04/26/11 19:03 jpchen target screens
  • 357232 04/26/11 18:39 johna Hall A tools screen
  • 357231 04/26/11 18:25 a-onl Start_Run_3920,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=90,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=3He production 120 muA Left Arm only - R arm ramping up,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357230 04/26/11 18:23 a-onl End_of_Run_3919
  • 357229 04/26/11 17:32 Rondon spot++ run 3819 3He production 120 muA
  • 357228 04/26/11 17:25 johna Begin production on 3He at 120uA (left arm only)
  • 357227 04/26/11 17:23 a-onl Start_Run_3919,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=90,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=3He production 120 muA Left Arm only - R arm ramping up,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357226 04/26/11 17:21 a-onl End_of_Run_3918
  • 357225 04/26/11 17:12 a-onl Start_Run_3918,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=90,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=3He ramping to 120 muA Left Arm only - R arm ramping up,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357224 04/26/11 16:55 a-onl End_of_Run_3917
  • 357223 04/26/11 16:55 Rondon spot++ raster 3x3 run 3917
  • 357222 04/26/11 16:51 a-onl Start_Run_3917,Run_type=Production SRC,target_type=BeO,beam_e0=3.356,beam_raster=on,beam_current=5,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=BeO 5 muA Left Arm only - R arm ramping up,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357221 04/26/11 16:51 jpchen raster on 4.5x3 (MCC) =3x3 (target)
  • 357220 04/26/11 16:50 jpchen harp scan results
  • 357219 04/26/11 16:41 jpchen 5 uA, raster off, harp scan
  • 357218 04/26/11 16:39 jpchen moved target to BeO
  • 357217 04/26/11 16:13 kelleher day shift summary
  • 357216 04/26/11 16:04 adaq Beam Time Accounting
  • 357215 04/26/11 14:50 fansler LQ1
  • 357214 04/26/11 14:37 kelleher MCC: ready for beam in 10 minutes
  • 357213 04/26/11 13:44 kelleher Ramped down all HRS magnets
  • 357212 04/26/11 13:00 auto_snap Radiation and Ozone snapshot
  • 357211 04/26/11 12:59 auto_snap Miscellaneous snapshot
  • 357210 04/26/11 12:59 auto_snap Hall A Tools snapshot
  • 357209 04/26/11 12:59 auto_snap Compton Polarimeter snapshot
  • 357208 04/26/11 12:59 auto_snap BigBox Magnet snapshot
  • 357207 04/26/11 12:59 auto_snap Beam Position and Current snapshot
  • 357206 04/26/11 11:07 Fansler/Smit LQ1 16 valve
  • 357205 04/26/11 08:38 yez Putting hall into control access
  • 357204 04/26/11 08:36 yez Beam is off. Finished all data needed for 3.2kin
  • 357203 04/26/11 08:35 yez Owl Shift Summary
  • 357202 04/26/11 08:31 a-onl End_of_Run_3916
  • 357201 04/26/11 08:09 adaq Beam Time Accounting
  • 357200 04/26/11 07:56 yez spot++ forrun3916
  • 357199 04/26/11 07:55 yez reboot TS0, TS11 ...
  • 357198 04/26/11 07:52 a-onl Start_Run_3916,Run_type=Production SRC,target_type=Ca48,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=Ca48 Production RQ3 2.6704 GeV/c -rigth arm only for now, PS6=6,raster 4x4,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357197 04/26/11 07:46 yez dam*! ROC10 disconnected, reset it.
  • 357196 04/26/11 07:27 yez spot++ for run3915
  • 357195 04/26/11 07:24 a-onl Start_Run_3915,Run_type=Production SRC,target_type=Ca48,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=Ca48 Production RQ3 2.6704 GeV/c -rigth arm only for now, PS6=6,raster 4x4,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357194 04/26/11 07:14 a-onl Start_Run_3914,Run_type=Production SRC,target_type=Ca48,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=Ca48 Production RQ3 2.6704 GeV/c -rigth arm only for now, PS6=6,raster 4x4,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357193 04/26/11 07:13 a-onl End_of_Run_3913
  • 357192 04/26/11 07:12 a-onl Start_Run_3913,Run_type=Production SRC,target_type=Ca48,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=Ca48 Production RQ3 2.6704 GeV/c -rigth arm only for now, PS6=14,raster 4x4,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357191 04/26/11 07:11 mhuang Target moved to Calcium-48
  • 357190 04/26/11 07:09 a-onl End_of_Run_3912
  • 357189 04/26/11 06:43 yez spot++ for run3912
  • 357188 04/26/11 06:39 a-onl Start_Run_3912,Run_type=Production SRC,target_type=Natural Ca40,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=Ca40 Production RQ3 2.6704 GeV/c -rigth arm only for now, PS6=14,raster 4x4,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357187 04/26/11 06:39 a-onl End_of_Run_3911
  • 357186 04/26/11 06:26 yez Beam study starts after 8am meeting
  • 357185 04/26/11 06:16 yez spot++ for run3911
  • 357184 04/26/11 06:13 a-onl Start_Run_3911,Run_type=Production SRC,target_type=Natural Ca40,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=Ca40 Production RQ3 2.6704 GeV/c -rigth arm only for now, PS6=14,raster 4x4,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357183 04/26/11 06:12 a-onl End_of_Run_3910
  • 357182 04/26/11 06:11 a-onl Start_Run_3910,Run_type=Production SRC,target_type=Natural Ca40,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=Ca40 Production RQ3 2.6704 GeV/c -rigth arm only for now, PS6=14,raster 4x4,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357181 04/26/11 06:10 a-onl End_of_Run_3909
  • 357180 04/26/11 06:09 yez spot++
  • 357179 04/26/11 06:06 a-onl Start_Run_3909,Run_type=Spot++,target_type=Natural Ca40,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=Ca40 Production RQ3 2.6704 GeV/c -rigth arm only for now, PS6=14,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357178 04/26/11 06:05 mhuang Target moved to Nat. Calcium
  • 357177 04/26/11 06:03 yez Finished Carbon data at 3.2, move to Ca40 target
  • 357176 04/26/11 06:00 a-onl End_of_Run_3908
  • 357175 04/26/11 05:45 a-onl Start_Run_3908,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=Carbon 5mm Production RQ3 2.6704 GeV/c -rigth arm only for now, PS6=14,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357174 04/26/11 05:44 a-onl End_of_Run_3907
  • 357173 04/26/11 05:33 yez spot++ for run3907
  • 357172 04/26/11 05:31 a-onl Start_Run_3907,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=Carbon 5mm Production RQ3 2.6704 GeV/c -rigth arm only for now, PS6=14,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357171 04/26/11 05:28 mhuang Target moved to carbon 5mm
  • 357170 04/26/11 05:26 yez Finished one hour run on Dummy, continue to take data on Carbon 5mm
  • 357169 04/26/11 05:25 a-onl End_of_Run_3906
  • 357168 04/26/11 04:59 auto_snap Radiation and Ozone snapshot
  • 357167 04/26/11 04:59 auto_snap Miscellaneous snapshot
  • 357166 04/26/11 04:59 auto_snap Hall A Tools snapshot
  • 357165 04/26/11 04:59 auto_snap Compton Polarimeter snapshot
  • 357164 04/26/11 04:59 auto_snap BigBox Magnet snapshot
  • 357163 04/26/11 04:59 auto_snap Beam Position and Current snapshot
  • 357162 04/26/11 04:58 a-onl Start_Run_3906,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=Dummy Production RQ3 2.6704 GeV/c -rigth arm only for now, PS6=4,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357161 04/26/11 04:57 a-onl End_of_Run_3905
  • 357160 04/26/11 04:24 a-onl Start_Run_3905,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=25,rightarm_sieve=out,comment_text=Dummy Production RQ3 2.6704 GeV/c -rigth arm only for now, PS6=4,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357159 04/26/11 04:21 mhuang Target moved to dummy 20cm
  • 357158 04/26/11 04:13 yez Finished He3 data at 3.2, move to Dummy target
  • 357157 04/26/11 04:11 a-onl End_of_Run_3904
  • 357156 04/26/11 04:01 mhuang Target screens
  • 357155 04/26/11 03:53 a-onl Start_Run_3904,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=25,rightarm_sieve=out,comment_text=3He Production RQ3 2.6704 GeV/c -rigth arm only for now,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357154 04/26/11 03:51 a-onl End_of_Run_3903
  • 357153 04/26/11 03:29 yez Hall A Tools screen and Beam Curren Graph
  • 357152 04/26/11 03:28 yez spot++
  • 357151 04/26/11 02:59 a-onl Start_Run_3903,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=25,rightarm_sieve=out,comment_text=3He Production RQ3 2.6704 GeV/c -rigth arm only for now,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357150 04/26/11 02:54 yez No Beam. MCC called that there is problem at their north lineac
  • 357149 04/26/11 02:53 a-onl End_of_Run_3902
  • 357148 04/26/11 02:46 yez Beam current is actually at 110uA
  • 357147 04/26/11 02:17 a-onl Start_Run_3902,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=25,rightarm_sieve=out,comment_text=3He Production RQ3 2.6704 GeV/c -rigth arm only for now,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357146 04/26/11 01:53 yez Not True. Still has no beam. MCC has problems to send beam with above 80uA
  • 357145 04/26/11 01:46 yez Beam is back after 25mins
  • 357144 04/26/11 01:26 yez MCC sends tune beam to check some problems. Run Stop
  • 357143 04/26/11 01:20 a-onl End_of_Run_3901
  • 357142 04/26/11 00:45 yez Reboot TS12 (BBTS1) and coda runs nomally now
  • 357141 04/26/11 00:34 a-onl Start_Run_3901,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=25,rightarm_sieve=out,comment_text=3He Production RQ3 2.6704 GeV/c -rigth arm only for now,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357140 04/26/11 00:29 mhuang Target screens
  • 357139 04/26/11 00:27 yez CODA take forever to connect BBTS1
  • 357138 04/26/11 00:15 adaq Beam Time Accounting
  • 357137 04/26/11 00:15 johna Swing shift summary
  • 357136 04/26/11 00:13 Rondon run 3900 stopped CODA problem
  • 357135 04/26/11 00:05 Kalyan Target temperature spikes and occasional alarms
  • Monday

  • 357134 04/25/11 23:51 johna Running right-arm only for the night
  • 357133 04/25/11 23:45 a-onl Start_Run_3900,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=25,rightarm_sieve=out,comment_text=3He Production RQ3 2.6704 GeV/c -rigth arm only for now ,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357132 04/25/11 23:41 a-onl End_of_Run_3899
  • 357131 04/25/11 23:39 a-onl Start_Run_3899,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=1,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=25,rightarm_sieve=out,comment_text=3He 1 muA RQ3 2.6704 GeV/c -rigth arm only now Beam Recover,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357130 04/25/11 23:39 Folts lt Q1 /tech
  • 357129 04/25/11 23:03 johna Hall C changing half wave plate
  • 357128 04/25/11 22:58 Kalyan Target temperature getting stable
  • 357127 04/25/11 22:49 johna Call for controlled access
  • 357126 04/25/11 22:49 A. Camsonne Left Q1 might have a valve issue - Ed Folts is on the way to the Hall - Requested to go to Control Access
  • 357125 04/25/11 22:46 a-onl End_of_Run_3898
  • 357124 04/25/11 22:27 a-onl Start_Run_3898,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=25,rightarm_sieve=out,comment_text=3He Production RQ3 2.6704 GeV/c Left Q1 tripped during 3896 -rigth arm only now,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357123 04/25/11 22:26 a-onl End_of_Run_3897
  • 357122 04/25/11 22:23 Kalyan Target temperature excursions
  • 357121 04/25/11 22:11 johna Q1 trip, low He level
  • 357120 04/25/11 21:54 Rondon run 3896 ROC sync OK; buffered mode back
  • 357119 04/25/11 21:52 a-onl Start_Run_3897,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=25,rightarm_sieve=out,comment_text=3He Production RQ3 2.6704 GeV/c Left Q1 tripped during 3896 -rigth arm only now,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357118 04/25/11 21:46 a-onl End_of_Run_3896
  • 357117 04/25/11 21:36 johna Trying out buffered mode.
  • 357116 04/25/11 21:35 a-onl Start_Run_3896,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=25,rightarm_sieve=out,comment_text=3He Production RQ3 2.6704 GeV/c both arms, buffered,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357115 04/25/11 21:34 a-onl End_of_Run_3895
  • 357114 04/25/11 21:24 a-onl Start_Run_3895,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=25,rightarm_sieve=out,comment_text=3He Production RQ3 2.6704 GeV/c left arm only Testing Buffered configuration,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357113 04/25/11 21:18 A. Camsonne Left Dipole NMR lock at 3.055 GeV/c
  • 357112 04/25/11 21:04 Rondon spot++ run 3894 3He 120 muA
  • 357111 04/25/11 21:04 johna Still running in unbuffered mode
  • 357110 04/25/11 20:59 auto_snap Radiation and Ozone snapshot
  • 357109 04/25/11 20:59 auto_snap Miscellaneous snapshot
  • 357108 04/25/11 20:59 auto_snap Hall A Tools snapshot
  • 357107 04/25/11 20:59 auto_snap Compton Polarimeter snapshot
  • 357106 04/25/11 20:59 auto_snap BigBox Magnet snapshot
  • 357105 04/25/11 20:59 auto_snap Beam Position and Current snapshot
  • 357104 04/25/11 20:57 a-onl Start_Run_3894,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=25,rightarm_sieve=out,comment_text=3He Production RQ3 2.6704 GeV/c left arm only,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357103 04/25/11 20:54 Kalyan Target screens for helium-3 target
  • 357102 04/25/11 20:53 a-onl End_of_Run_3893
  • 357101 04/25/11 20:51 johna Both spectrometers at 21 degrees; camera snapshots attached
  • 357100 04/25/11 20:38 johna Partially set up for Kin 3.2
  • 357099 04/25/11 20:37 a-onl Start_Run_3893,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=25,rightarm_sieve=out,comment_text=3He Production RQ3 2.6704 GeV/c,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357098 04/25/11 20:35 a-onl End_of_Run_3892
  • 357097 04/25/11 20:34 a-onl Start_Run_3892,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=BeO, all PS 1 RASTER 3x3,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357096 04/25/11 20:09 Rondon CODA restarted, TSO unresponsive
  • 357095 04/25/11 20:03 Kalyan Target to 20cm Helium-3
  • 357094 04/25/11 19:53 a-onl End_of_Run_3891
  • 357093 04/25/11 19:44 a-onl Start_Run_3891,Run_type=Production SRC,target_type=BeO,beam_e0=3.356,beam_raster=on,beam_current=5,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=BeO, all PS 1 RASTER 3x3,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357092 04/25/11 19:42 Kalyan Target to BeO
  • 357091 04/25/11 19:40 a-onl End_of_Run_3890
  • 357090 04/25/11 19:29 Rondon spot++ run3890 raster 3x3
  • 357089 04/25/11 19:25 a-onl Start_Run_3890,Run_type=Production SRC,target_type=Optics: multi 1mm C foils,beam_e0=3.356,beam_raster=on,beam_current=5,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=optics all PS 1 RASTER 3x3,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357088 04/25/11 19:23 johna Set raster back to 3x3
  • 357087 04/25/11 19:22 a-onl End_of_Run_3889
  • 357086 04/25/11 19:16 a-onl Start_Run_3889,Run_type=Production SRC,target_type=Optics: multi 1mm C foils,beam_e0=3.356,beam_raster=on,beam_current=5,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=optics all PS 1,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357085 04/25/11 19:11 Kalyan Target to optics
  • 357084 04/25/11 19:09 a-onl End_of_Run_3888
  • 357083 04/25/11 19:00 Rondon spot++ run3888
  • 357082 04/25/11 18:54 Kalyan Target to calcium-40
  • 357081 04/25/11 18:49 a-onl Start_Run_3888,Run_type=Production SRC,target_type=Natural Ca40,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=calcium 40 unbuffered mode to fix ROC sync,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357080 04/25/11 18:44 a-onl End_of_Run_3887
  • 357079 04/25/11 18:40 yez Most of Out-Sync events are from T1&T3
  • 357078 04/25/11 18:39 Rondon Unbuffered mode fixes ROC sync
  • 357077 04/25/11 18:21 a-onl Start_Run_3887,Run_type=Production SRC,target_type=Ca48,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=calcium 48 unbuffered mode to test ROC sync,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357076 04/25/11 18:17 johna Switching to unbuffered mode while we diagnose sync problems
  • 357075 04/25/11 18:17 a-onl End_of_Run_3886
  • 357074 04/25/11 18:12 Rondon spot++ run 3884
  • 357073 04/25/11 18:10 a-onl Start_Run_3886,Run_type=Production SRC,target_type=Ca48,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=calcium 48,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357072 04/25/11 18:05 a-onl End_of_Run_3885
  • 357071 04/25/11 18:04 yez Syncs Problem starting from#3875
  • 357070 04/25/11 18:04 Kalyan late entry: target is calcium-48
  • 357069 04/25/11 18:02 Rondon ROC's ou of sync, run 3884
  • 357068 04/25/11 18:01 a-onl Start_Run_3885,Run_type=Production SRC,target_type=Ca48,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=calcium 48,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357067 04/25/11 17:59 a-onl End_of_Run_3884
  • 357066 04/25/11 17:50 a-onl Start_Run_3884,Run_type=Production SRC,target_type=Ca48,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=calcium 48,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357065 04/25/11 17:49 a-onl End_of_Run_3883
  • 357064 04/25/11 17:43 Rondon spot++ raster 4x4
  • 357063 04/25/11 17:39 a-onl Start_Run_3883,Run_type=Production SRC,target_type=Ca48,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=calcium 48,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357062 04/25/11 17:31 a-onl End_of_Run_3882
  • 357061 04/25/11 17:18 johna Raster changed to 4x4 (has been at 3x3)
  • 357060 04/25/11 17:16 a-onl Start_Run_3882,Run_type=Production SRC,target_type=Natural Ca40,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=calcium 40 target. Beam back to nominal, actual 4x4 raster,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357059 04/25/11 17:15 johna BPM stripchart for position scan on Ca
  • 357058 04/25/11 17:04 a-onl End_of_Run_3881
  • 357057 04/25/11 17:04 johna position scan on 40Ca
  • 357056 04/25/11 17:03 johna Raster change didn't appear to modify beam size
  • 357055 04/25/11 16:55 a-onl Start_Run_3881,Run_type=Production SRC,target_type=Natural Ca40,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=calcium 40 target. Move beam -1 in x,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357054 04/25/11 16:52 a-onl End_of_Run_3880
  • 357053 04/25/11 16:45 a-onl Start_Run_3880,Run_type=Production SRC,target_type=Natural Ca40,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=calcium 40 target. Move beam +1 in x,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357052 04/25/11 16:44 a-onl End_of_Run_3879
  • 357051 04/25/11 16:43 a-onl Start_Run_3879,Run_type=Production SRC,target_type=Natural Ca40,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=calcium 40 target. Move beam +1 in x,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357050 04/25/11 16:40 a-onl End_of_Run_3878
  • 357049 04/25/11 16:32 kelleher spot++ for 3878
  • 357048 04/25/11 16:31 kelleher correction run #3877 is +1mm in y
  • 357047 04/25/11 16:31 Kalyan yellow alarm on target
  • 357046 04/25/11 16:31 a-onl Start_Run_3878,Run_type=Production SRC,target_type=Natural Ca40,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=calcium 40 target. Move beam -1 in y,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357045 04/25/11 16:29 a-onl End_of_Run_3877
  • 357044 04/25/11 16:22 kelleher run 3877, spot++
  • 357043 04/25/11 16:20 a-onl Start_Run_3877,Run_type=Production SRC,target_type=Natural Ca40,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=calcium 40 target. Move beam up 1 in x,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357042 04/25/11 16:17 a-onl End_of_Run_3876
  • 357041 04/25/11 16:12 kelleher day shift summary
  • 357040 04/25/11 16:02 adaq Beam Time Accounting
  • 357039 04/25/11 16:02 Leckey Target Moved to Calcium 40 Target
  • 357038 04/25/11 16:00 kelleher spot++ for 3876
  • 357037 04/25/11 15:57 a-onl Start_Run_3876,Run_type=Production SRC,target_type=Natural Ca40,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=calcium 40 target.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357036 04/25/11 15:56 a-onl End_of_Run_3875
  • 357035 04/25/11 15:52 kelleher raster pictures
  • 357034 04/25/11 15:50 kelleher 40Ca run #3875 spot++
  • 357033 04/25/11 15:44 a-onl Start_Run_3875,Run_type=Production SRC,target_type=Natural Ca40,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=calcium 40 target.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357032 04/25/11 15:35 a-onl End_of_Run_3874
  • 357031 04/25/11 15:19 kelleher Hall A tools
  • 357030 04/25/11 15:18 kelleher raster 3x3
  • 357029 04/25/11 15:15 a-onl Start_Run_3874,Run_type=Production SRC,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=10 cm dummy target. Current Scan.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357028 04/25/11 14:51 A. Camsonne x >2 experiment status
  • 357027 04/25/11 14:41 a-onl End_of_Run_3873
  • 357026 04/25/11 14:39 Leckey Closed the JT 2 % on loop 1
  • 357025 04/25/11 14:22 Leckey Target Snapshot
  • 357024 04/25/11 14:21 a-onl Start_Run_3873,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm dummy target. Current Scan.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357023 04/25/11 14:20 a-onl End_of_Run_3872
  • 357022 04/25/11 14:20 Leckey Target Moved to 20 or 15 cm Dummy Target
  • 357021 04/25/11 14:19 a-onl Start_Run_3872,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm dummy target. Current Scan.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357020 04/25/11 14:13 a-onl End_of_Run_3871
  • 357019 04/25/11 14:02 a-onl Start_Run_3871,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=20,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=carbon 12 target. Current Scan.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357018 04/25/11 14:00 a-onl End_of_Run_3870
  • 357017 04/25/11 13:49 a-onl Start_Run_3870,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=carbon 12 target. Current Scan.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357016 04/25/11 12:59 auto_snap Radiation and Ozone snapshot
  • 357015 04/25/11 12:59 auto_snap Miscellaneous snapshot
  • 357014 04/25/11 12:58 auto_snap Hall A Tools snapshot
  • 357013 04/25/11 12:58 auto_snap Compton Polarimeter snapshot
  • 357012 04/25/11 12:58 auto_snap BigBox Magnet snapshot
  • 357011 04/25/11 12:58 auto_snap Beam Position and Current snapshot
  • 357010 04/25/11 12:32 a-onl End_of_Run_3869
  • 357009 04/25/11 12:24 a-onl Start_Run_3869,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=60,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=carbon 12 target. Current Scan.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357008 04/25/11 12:24 a-onl Start_Run_3866,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=carbon 12 target. Current Scan.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357007 04/25/11 12:21 a-onl End_of_Run_3868
  • 357006 04/25/11 12:20 yez Tried to add HRS HVs in Start_Of_Run script but it looks like the next three runs don't show on the halog, so I put it back
  • 357005 04/25/11 12:17 a-onl Start_Run_3868,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=60,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=carbon 12 target. Current Scan.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357004 04/25/11 12:16 a-onl Start_Run_3865,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=carbon 12 target. Current Scan.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357003 04/25/11 12:15 a-onl End_of_Run_3867
  • 357002 04/25/11 12:08 a-onl Start_Run_3864,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=carbon 12 target. Current Scan.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357001 04/25/11 12:08 a-onl Start_Run_3867,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=80,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=carbon 12 target. Current Scan.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 357000 04/25/11 12:07 a-onl End_of_Run_3866
  • 356999 04/25/11 12:05 a-onl End_of_Run_3865
  • 356998 04/25/11 11:58 a-onl End_of_Run_3864
  • 356997 04/25/11 11:51 a-onl End_of_Run_3863
  • 356996 04/25/11 11:43 a-onl Start_Run_3863,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=carbon 12 target. Current Scan.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356995 04/25/11 11:42 a-onl End_of_Run_3862
  • 356994 04/25/11 11:39 a-onl Start_Run_3862,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=carbon 12 target. Current Scan.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356993 04/25/11 11:21 a-onl End_of_Run_3861
  • 356992 04/25/11 11:12 a-onl Start_Run_3861,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=20,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=helium three target. Current Scan.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356991 04/25/11 11:10 a-onl End_of_Run_3860
  • 356990 04/25/11 10:51 a-onl Start_Run_3860,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=20,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=helium three target. Current Scan.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356989 04/25/11 10:49 a-onl End_of_Run_3859
  • 356988 04/25/11 10:34 a-onl Start_Run_3859,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=helium three target. Current Scan.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356987 04/25/11 10:33 a-onl End_of_Run_3858
  • 356986 04/25/11 10:21 a-onl Start_Run_3858,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=60,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=helium three target. Current Scan.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356985 04/25/11 10:16 a-onl End_of_Run_3857
  • 356984 04/25/11 10:12 a-onl Start_Run_3857,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=60,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=helium three target. Current Scan.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356983 04/25/11 10:11 a-onl End_of_Run_3856
  • 356982 04/25/11 10:04 LeRose Quad Field Readings
  • 356981 04/25/11 10:03 a-onl Start_Run_3856,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=80,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=helium three target. Current Scan.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356980 04/25/11 10:00 a-onl End_of_Run_3855
  • 356979 04/25/11 09:54 a-onl Start_Run_3855,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=100,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=helium three target. Current Scan.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356978 04/25/11 09:50 a-onl End_of_Run_3854
  • 356977 04/25/11 09:43 a-onl Start_Run_3854,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=helium three target. Right arm not settled - Pescaled T1 while Right arm settling by 2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356976 04/25/11 09:42 a-onl End_of_Run_3853
  • 356975 04/25/11 09:33 a-onl Start_Run_3853,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=helium three target. Right arm not settled - Pescaled T1 while Right arm settling by 2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356974 04/25/11 09:32 a-onl End_of_Run_3852
  • 356973 04/25/11 09:25 a-onl Start_Run_3852,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=helium three target. Right arm not settled - Pescaled T1 while Right arm settling by 2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356972 04/25/11 09:24 a-onl End_of_Run_3851
  • 356971 04/25/11 09:15 a-onl Start_Run_3851,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=helium three target. Right arm not settled - Pescaled T1 while Right arm settling by 2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356970 04/25/11 09:14 a-onl End_of_Run_3850
  • 356969 04/25/11 09:06 a-onl Start_Run_3850,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=helium three target. Right arm not settled - Pescaled T1 while Right arm settling by 2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356968 04/25/11 09:05 a-onl End_of_Run_3849
  • 356967 04/25/11 09:02 A. Camsonne Left HRS NMR locks nicely with probe B
  • 356966 04/25/11 08:57 a-onl Start_Run_3849,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=helium three target. Right arm not settled - Pescaled T1 while Right arm settling by 2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356965 04/25/11 08:56 a-onl End_of_Run_3848
  • 356964 04/25/11 08:53 a-onl Start_Run_3848,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=helium three target. Right arm not settled,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356963 04/25/11 08:52 a-onl End_of_Run_3847
  • 356962 04/25/11 08:47 a-onl Start_Run_3847,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=helium three target.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356961 04/25/11 08:46 a-onl End_of_Run_3846
  • 356960 04/25/11 08:45 a-onl Start_Run_3846,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=helium three target.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356959 04/25/11 08:42 a-onl End_of_Run_3845
  • 356958 04/25/11 08:38 a-onl Start_Run_3845,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.505,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25,rightarm_sieve=out,comment_text=helium three target. ,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356957 04/25/11 08:25 Leckey Target moved to 20cm Helium
  • 356956 04/25/11 08:16 a-onl End_of_Run_3844
  • 356955 04/25/11 08:02 adaq Beam Time Accounting
  • 356954 04/25/11 08:02 wdconinc Owl shift summary
  • 356953 04/25/11 07:58 a-onl Start_Run_3844,Run_type=Production SRC,target_type=Natural Ca40,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=Ca48 Production Run,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356952 04/25/11 07:56 gueye CODA crash
  • 356951 04/25/11 07:55 a-onl End_of_Run_3843
  • 356950 04/25/11 07:34 a-onl Start_Run_3843,Run_type=Production SRC,target_type=Natural Ca40,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=Ca48 Production Run,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356949 04/25/11 07:34 a-onl End_of_Run_3842
  • 356948 04/25/11 07:03 a-onl Start_Run_3842,Run_type=Production SRC,target_type=Natural Ca40,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=Ca48 Production Run,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356947 04/25/11 07:02 a-onl End_of_Run_3841
  • 356946 04/25/11 06:36 a-onl Start_Run_3841,Run_type=Production SRC,target_type=Natural Ca40,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=Ca48 Production Run,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356945 04/25/11 06:36 gueye Beam back - start run 3841
  • 356944 04/25/11 06:29 gueye MCC called - stil trying to solve the problem
  • 356943 04/25/11 06:19 a-onl End_of_Run_3840
  • 356942 04/25/11 06:19 gueye Beam is gone again - stopped run 3840
  • 356941 04/25/11 06:01 a-onl Start_Run_3840,Run_type=Production SRC,target_type=Natural Ca40,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=Ca48 Production Run,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356940 04/25/11 06:00 gueye Beam back - PS replaced
  • 356939 04/25/11 05:52 wdconinc General tools screen shot
  • 356938 04/25/11 05:52 a-onl End_of_Run_3839
  • 356937 04/25/11 05:45 wdconinc Target screen shots
  • 356936 04/25/11 05:45 a-onl Start_Run_3839,Run_type=Production SRC,target_type=Natural Ca40,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=Ca48 Production Run,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356935 04/25/11 05:44 gueye Beam back. Started run 3839
  • 356934 04/25/11 05:40 gueye Same IOC problem for run 3838
  • 356933 04/25/11 05:39 a-onl End_of_Run_3838
  • 356932 04/25/11 05:15 a-onl Start_Run_3838,Run_type=Production SRC,target_type=Natural Ca40,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=Ca48 Production Run,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356931 04/25/11 05:12 gueye Beam back!
  • 356930 04/25/11 04:58 auto_snap Radiation and Ozone snapshot
  • 356929 04/25/11 04:58 auto_snap Miscellaneous snapshot
  • 356928 04/25/11 04:58 auto_snap Hall A Tools snapshot
  • 356927 04/25/11 04:58 auto_snap Compton Polarimeter snapshot
  • 356926 04/25/11 04:58 auto_snap BigBox Magnet snapshot
  • 356925 04/25/11 04:58 auto_snap Beam Position and Current snapshot
  • 356924 04/25/11 04:44 gueye Outside experts on their way
  • 356923 04/25/11 04:40 gueye MCC called - Outside experts are on their way
  • 356922 04/25/11 03:36 gueye MCC called - CAMAC crate down
  • 356921 04/25/11 03:17 a-onl End_of_Run_3837
  • 356920 04/25/11 03:12 a-onl Start_Run_3837,Run_type=Production SRC,target_type=Natural Ca40,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=Ca48 Production Run,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356919 04/25/11 03:11 a-onl End_of_Run_3836
  • 356918 04/25/11 02:50 gueye spot++ - run 3836
  • 356917 04/25/11 02:49 gueye 1st run on Ca-40 - run 3836
  • 356916 04/25/11 02:48 gueye CODA hang when finishing run3835
  • 356915 04/25/11 02:47 a-onl Start_Run_3836,Run_type=Production SRC,target_type=Natural Ca40,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=Ca48 Production Run,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356914 04/25/11 02:39 wdconinc Moving to natural 40Ca target
  • 356913 04/25/11 02:38 a-onl End_of_Run_3835
  • 356912 04/25/11 02:11 gueye spot++ - run3835
  • 356911 04/25/11 02:09 a-onl Start_Run_3835,Run_type=Production SRC,target_type=Ca48,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=Ca48 Production Run,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356910 04/25/11 02:06 gueye ROC10 crashed - reboot CODA
  • 356909 04/25/11 01:39 gueye spot++ - run 3834
  • 356908 04/25/11 01:36 a-onl Start_Run_3834,Run_type=Production SRC,target_type=Ca48,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=Ca48 Production Run,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356907 04/25/11 01:35 a-onl End_of_Run_3833
  • 356906 04/25/11 01:09 gueye spot++ - run3833
  • 356905 04/25/11 01:06 a-onl Start_Run_3833,Run_type=Production SRC,target_type=Ca48,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=Ca48 Production Run,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356904 04/25/11 01:06 a-onl End_of_Run_3832
  • 356903 04/25/11 00:42 gueye Spot++ (run 3832) - with picture!
  • 356902 04/25/11 00:41 gueye Beam position (run 3832)
  • 356901 04/25/11 00:37 a-onl Start_Run_3832,Run_type=Production SRC,target_type=Ca48,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=Ca48 Production Run,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356900 04/25/11 00:36 a-onl End_of_Run_3831
  • 356899 04/25/11 00:30 wdconinc Run plan for owl shift
  • 356898 04/25/11 00:25 yez Set Golden run to run#3820
  • 356897 04/25/11 00:13 johna Swing shift summary
  • 356896 04/25/11 00:12 adaq Beam Time Accounting
  • 356895 04/25/11 00:10 gueye Beam position (run 3831)
  • 356894 04/25/11 00:05 a-onl Start_Run_3831,Run_type=Production SRC,target_type=Ca48,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=Ca48 Production Run,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356893 04/25/11 00:04 a-onl End_of_Run_3830
  • Sunday

  • 356892 04/24/11 23:39 yez SPOT++ for run3830
  • 356891 04/24/11 23:36 a-onl Start_Run_3830,Run_type=Production SRC,target_type=Ca48,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=Ca48 Production Run,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356890 04/24/11 23:29 a-onl End_of_Run_3829
  • 356889 04/24/11 23:28 xqiu Moved to Calcium-48 target and asked for 5uA beam.
  • 356888 04/24/11 23:27 a-onl Start_Run_3829,Run_type=Spot++,target_type=Ca48,beam_e0=3.356,beam_raster=on,beam_current=5,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=spot++, 6x4.5,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356887 04/24/11 23:23 a-onl End_of_Run_3828
  • 356886 04/24/11 22:37 a-onl Start_Run_3828,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm dummy 40 micro amps, checked ps's.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356885 04/24/11 22:37 a-onl End_of_Run_3826
  • 356884 04/24/11 22:32 a-onl End_of_Run_3827
  • 356883 04/24/11 21:56 a-onl Start_Run_3827,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm dummy 40 micro amps, checked ps's.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356882 04/24/11 21:39 Donal Spot++
  • 356881 04/24/11 21:29 yez HV of RHRS #8.4 was off. Turn it back on
  • 356880 04/24/11 21:22 xqiu Moved to 20cm Dummy target and asked for 40uA beam.
  • 356879 04/24/11 21:20 a-onl Start_Run_3826,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm dummy 40 micro amps, checked ps's.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356878 04/24/11 21:13 a-onl End_of_Run_3825
  • 356877 04/24/11 20:59 a-onl Start_Run_3825,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm He3, Beam current = 119 uA,SRC production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356876 04/24/11 20:58 auto_snap Radiation and Ozone snapshot
  • 356875 04/24/11 20:58 auto_snap Miscellaneous snapshot
  • 356874 04/24/11 20:58 auto_snap Hall A Tools snapshot
  • 356873 04/24/11 20:58 auto_snap Compton Polarimeter snapshot
  • 356872 04/24/11 20:58 auto_snap BigBox Magnet snapshot
  • 356871 04/24/11 20:58 auto_snap Beam Position and Current snapshot
  • 356870 04/24/11 20:58 a-onl End_of_Run_3824
  • 356869 04/24/11 20:43 a-onl Start_Run_3824,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm He3, Beam current = 119 uA,SRC production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356868 04/24/11 20:42 a-onl End_of_Run_3823
  • 356867 04/24/11 20:31 xqiu target status: Carbon 5mm, 95uA beam
  • 356866 04/24/11 20:27 a-onl Start_Run_3823,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm He3, Beam current = 119 uA,SRC production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356865 04/24/11 20:26 a-onl End_of_Run_3822
  • 356864 04/24/11 20:25 Donal CODA
  • 356863 04/24/11 20:25 a-onl Start_Run_3822,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm He3, Beam current = 119 uA,SRC production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356862 04/24/11 20:23 johna Fewer trips at 116uA
  • 356861 04/24/11 20:20 xqiu Ask MCC to mask the beam and moving to Carbon 5mm target.
  • 356860 04/24/11 20:14 a-onl End_of_Run_3821
  • 356859 04/24/11 19:51 a-onl Start_Run_3821,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm He3, Beam current = 119 uA,SRC production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356858 04/24/11 19:50 a-onl End_of_Run_3820
  • 356857 04/24/11 19:27 a-onl Start_Run_3820,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm He3, Beam current = 119 uA,SRC production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356856 04/24/11 19:25 a-onl End_of_Run_3819
  • 356855 04/24/11 19:21 folts tech on call
  • 356854 04/24/11 19:04 A. Camsonne Trying running at 116 uA
  • 356853 04/24/11 19:02 a-onl Start_Run_3819,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm He3, Beam current = 119 uA,SRC production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356852 04/24/11 19:01 a-onl End_of_Run_3818
  • 356851 04/24/11 18:53 johna HRS angle camera snapshots
  • 356850 04/24/11 18:42 johna HRS-R HV problems (and left arm monitoring problem)
  • 356849 04/24/11 18:37 a-onl Start_Run_3818,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm He3, Beam current = 119 uA,SRC production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356848 04/24/11 18:36 a-onl End_of_Run_3817
  • 356847 04/24/11 18:20 xqiu target status
  • 356846 04/24/11 18:18 johna Right arm triggers have stopped
  • 356845 04/24/11 18:16 a-onl Start_Run_3817,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm He3, Beam current = 119 uA,SRC production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356844 04/24/11 18:04 a-onl End_of_Run_3816
  • 356843 04/24/11 17:55 a-onl Start_Run_3816,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm He3, Beam current = 119 uA,SRC production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356842 04/24/11 17:53 a-onl End_of_Run_3815
  • 356841 04/24/11 17:36 johna Half-wave plate in
  • 356840 04/24/11 17:28 a-onl Start_Run_3815,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm He3, Beam current = 119 uA,SRC production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356839 04/24/11 17:27 a-onl End_of_Run_3814
  • 356838 04/24/11 17:00 xqiu target snapshot
  • 356837 04/24/11 16:58 a-onl Start_Run_3814,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm He3, Beam current = 119 uA,SRC production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 356836 04/24/11 16:52 a-onl End_of_Run_3813
  • 356835 04/24/11 16:24 a-onl Start_Run_3813,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=119,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm He3, Beam current = 119 uA,SRC production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356834 04/24/11 16:23 a-onl End_of_Run_3812
  • 356833 04/24/11 16:02 adaq Beam Time Accounting
  • 356832 04/24/11 16:01 Solvignon Day Shift Summary
  • 356831 04/24/11 16:00 a-onl Start_Run_3812,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=119,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm He3, Beam current = 119 uA,SRC production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356830 04/24/11 15:58 a-onl End_of_Run_3811
  • 356829 04/24/11 15:36 a-onl Start_Run_3811,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=119,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm He3, Beam current = 119 uA,SRC production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356828 04/24/11 15:34 a-onl End_of_Run_3810
  • 356827 04/24/11 15:23 pzhu target status
  • 356826 04/24/11 15:12 a-onl Start_Run_3810,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=119,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm He3, Beam current = 119 uA,SRC production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356825 04/24/11 15:08 a-onl End_of_Run_3809
  • 356824 04/24/11 14:44 a-onl Start_Run_3809,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=119,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm He3, Beam current = 119 uA,SRC production right arm junk,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356823 04/24/11 14:42 a-onl End_of_Run_3808
  • 356822 04/24/11 14:40 a-onl Start_Run_3808,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=119,leftarm_p=2.795,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm He3, Beam current = 119 uA,SRC production right arm junk,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356821 04/24/11 14:13 Solvignon Done with kin 5.2. Setting HRSs for kin 5.1.
  • 356820 04/24/11 14:12 a-onl End_of_Run_3807
  • 356819 04/24/11 13:41 pzhu target status
  • 356818 04/24/11 13:40 Solvignon Beam unstable at 120uA. Requested 115uA.
  • 356817 04/24/11 13:21 R. Michaels run 3778 is in MSS
  • 356816 04/24/11 13:15 a-onl Start_Run_3807,Run_type=Production SRC,target_type=3He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm He3, Beam current = 120 uA,SRC production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356815 04/24/11 13:13 Solvignon Done with Dummy and 12C for kin 5.2. Back to 3He production
  • 356814 04/24/11 13:02 a-onl End_of_Run_3806
  • 356813 04/24/11 12:58 auto_snap Radiation and Ozone snapshot
  • 356812 04/24/11 12:58 auto_snap Miscellaneous snapshot
  • 356811 04/24/11 12:58 auto_snap Hall A Tools snapshot
  • 356810 04/24/11 12:58 auto_snap Compton Polarimeter snapshot
  • 356809 04/24/11 12:57 auto_snap BigBox Magnet snapshot
  • 356808 04/24/11 12:57 auto_snap Beam Position and Current snapshot
  • 356807 04/24/11 12:18 yez raw data #3778 cannot move to MSS
  • 356806 04/24/11 12:16 a-onl Start_Run_3806,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm dummy, Beam current = 40 uA,SRC production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356805 04/24/11 12:15 a-onl End_of_Run_3805
  • 356804 04/24/11 12:11 A. Camsonne S2m ADC 3794
  • 356803 04/24/11 11:54 Jack Tech On Call
  • 356802 04/24/11 11:48 yez S2m started to have problems on ADC during run#3793
  • 356801 04/24/11 11:30 a-onl Start_Run_3805,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm dummy, Beam current = 40 uA,SRC production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356800 04/24/11 11:29 a-onl End_of_Run_3804
  • 356799 04/24/11 11:17 Solvignon Hall A general tools during checklist
  • 356798 04/24/11 11:15 Solvignon Spot++ for checklist
  • 356797 04/24/11 10:32 A. Camsonne Scalers 40 uA Al dummy
  • 356796 04/24/11 10:29 a-onl Start_Run_3804,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm dummy, Beam current = 40 uA,SRC production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356795 04/24/11 10:28 A. Camsonne S2 ADC values for Left arm
  • 356794 04/24/11 10:28 a-onl End_of_Run_3803
  • 356793 04/24/11 10:23 chench run3803 left arm is junk since LQ1 is tripped.
  • 356792 04/24/11 10:20 Solvignon LQ1 ramping up
  • 356791 04/24/11 10:07 a-onl Start_Run_3803,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=20 cm dummy, Beam current = 40 uA,Left arm check ,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356790 04/24/11 10:04 a-onl End_of_Run_3802
  • 356789 04/24/11 10:04 Solvignon LQ1 tripped
  • 356788 04/24/11 10:00 Solvignon LQ1 cryo low but still holding
  • 356787 04/24/11 09:21 a-onl Start_Run_3802,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, 20 cm dummy, Beam current = 40 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356786 04/24/11 09:21 Solvignon Beam HWP OUT
  • 356785 04/24/11 09:19 a-onl End_of_Run_3801
  • 356784 04/24/11 09:11 a-onl Start_Run_3801,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, 20 cm dummy, Beam current = 40 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356783 04/24/11 09:10 a-onl End_of_Run_3800
  • 356782 04/24/11 08:35 chench Screenshot of beam monitor and DAQ hands up display
  • 356781 04/24/11 08:25 a-onl Start_Run_3800,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, 20 cm dummy, Beam current = 40 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356780 04/24/11 08:16 chench online GUI from run3799
  • 356779 04/24/11 08:11 a-onl End_of_Run_3799
  • 356778 04/24/11 08:04 adaq Beam Time Accounting
  • 356777 04/24/11 08:03 W. Deconinck Owl shift summary
  • 356776 04/24/11 08:02 a-onl Start_Run_3799,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, Carbon 5mm Target, Beam current = 95 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356775 04/24/11 07:59 a-onl End_of_Run_3798
  • 356774 04/24/11 07:49 R. Michaels re: S2m performance
  • 356773 04/24/11 07:43 xbai He3 run plot of S2m when it started
  • 356772 04/24/11 07:35 R. Michaels added He3 to pulldown menu
  • 356771 04/24/11 07:32 A. Camsonne S2m changes
  • 356770 04/24/11 07:26 xbai L-S2m issue still exists
  • 356769 04/24/11 07:22 A. Camsonne Power cycled ROC3 and reloaded trigger
  • 356768 04/24/11 07:20 a-onl Start_Run_3798,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, Carbon 5mm Target, Beam current = 95 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356767 04/24/11 07:08 wdconinc Alex looking into L-S2m issue
  • 356766 04/24/11 07:06 a-onl End_of_Run_3797
  • 356765 04/24/11 07:03 wdconinc Target screen shots
  • 356764 04/24/11 06:34 a-onl Start_Run_3797,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, Carbon 5mm Target, Beam current = 95 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356763 04/24/11 06:32 a-onl End_of_Run_3796
  • 356762 04/24/11 06:04 xbai Carbon target spot++ & Raster status snapshot
  • 356761 04/24/11 06:02 wdconinc Hall A general tools screen shot
  • 356760 04/24/11 06:02 xbai Changed target from 3He to Carbon 5mm
  • 356759 04/24/11 05:59 wdconinc Beam is Back!
  • 356758 04/24/11 05:58 a-onl Start_Run_3796,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, Carbon 5mm Target, Beam current = 95 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356757 04/24/11 05:53 wdconinc Vac valve closed in linac, waiting for open sesame
  • 356756 04/24/11 05:52 a-onl End_of_Run_3795
  • 356755 04/24/11 05:40 a-onl Start_Run_3795,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, Carbon 5mm Target, Beam current = 95 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356754 04/24/11 05:35 xbai start coda
  • 356753 04/24/11 05:30 wdconinc Moving target to Carbon 5mm
  • 356752 04/24/11 05:00 a-onl Start_Run_3794,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, 3He 20cm Target, Beam current = 120 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356751 04/24/11 04:59 a-onl End_of_Run_3793
  • 356750 04/24/11 04:57 auto_snap Radiation and Ozone snapshot
  • 356749 04/24/11 04:57 auto_snap Miscellaneous snapshot
  • 356748 04/24/11 04:57 auto_snap Hall A Tools snapshot
  • 356747 04/24/11 04:57 auto_snap Compton Polarimeter snapshot
  • 356746 04/24/11 04:57 auto_snap BigBox Magnet snapshot
  • 356745 04/24/11 04:57 auto_snap Beam Position and Current snapshot
  • 356744 04/24/11 04:11 xbai CODA Restart
  • 356743 04/24/11 04:10 a-onl Start_Run_3793,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, 3He 20cm Target, Beam current = 120 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356742 04/24/11 04:00 a-onl End_of_Run_3792
  • 356741 04/24/11 03:07 a-onl Start_Run_3792,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, 3He 20cm Target, Beam current = 120 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356740 04/24/11 03:06 a-onl End_of_Run_3791
  • 356739 04/24/11 02:16 a-onl Start_Run_3791,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, 3He 20cm Target, Beam current = 120 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356738 04/24/11 02:15 a-onl End_of_Run_3790
  • 356737 04/24/11 01:26 a-onl Start_Run_3790,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, 3He 20cm Target, Beam current = 120 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356736 04/24/11 01:23 a-onl End_of_Run_3789
  • 356735 04/24/11 01:16 yez Use the name '20cm He4 ' for 20cm He3 target
  • 356734 04/24/11 00:59 wdconinc Hall C announces: IHWP change
  • 356733 04/24/11 00:34 xbai/wdconin No 3He_20cm entry in runstart target_type dropdown list
  • 356732 04/24/11 00:31 a-onl Start_Run_3789,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, 3He 20cm Target, Beam current = 120 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356731 04/24/11 00:30 a-onl End_of_Run_3788
  • 356730 04/24/11 00:12 wdconinc Plans for the owl shift: production on 3He and Carbon
  • 356729 04/24/11 00:03 adaq Beam Time Accounting
  • 356728 04/24/11 00:00 johna Swing shfit summary
  • Saturday

  • 356727 04/23/11 23:44 miham Spot++ check for run #3788
  • 356726 04/23/11 23:41 miham Coda GUI screenshot for run #3788
  • 356725 04/23/11 23:40 a-onl Start_Run_3788,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, 3He 20cm Target, Beam current = 120 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356724 04/23/11 23:39 a-onl End_of_Run_3787
  • 356723 04/23/11 23:08 xqiu target snapshot
  • 356722 04/23/11 22:54 A. Camsonne Turned beamline HV crate off and on
  • 356721 04/23/11 22:52 miham Spot++ check for run #3787
  • 356720 04/23/11 22:50 miham Coda GUI screenshot for run #3787
  • 356719 04/23/11 22:46 a-onl Start_Run_3787,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, 3He 20cm Target, Beam current = 120 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356718 04/23/11 22:33 johna Beam down; MCC needs to bypass an RF cavity
  • 356717 04/23/11 22:31 a-onl End_of_Run_3786
  • 356716 04/23/11 21:56 miham Spot++ check for run #3786
  • 356715 04/23/11 21:55 miham Coda GUI screenshot for run #3786
  • 356714 04/23/11 21:54 a-onl Start_Run_3786,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, 3He 20cm Target, Beam current = 120 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356713 04/23/11 21:53 a-onl End_of_Run_3785
  • 356712 04/23/11 21:06 miham Spot++ check for run #3785
  • 356711 04/23/11 21:05 miham Coda GUI screenshot for run #3785
  • 356710 04/23/11 21:03 a-onl Start_Run_3785,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, 3He 20cm Target, Beam current = 120 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356709 04/23/11 21:03 a-onl End_of_Run_3784
  • 356708 04/23/11 20:57 auto_snap Radiation and Ozone snapshot
  • 356707 04/23/11 20:57 auto_snap Miscellaneous snapshot
  • 356706 04/23/11 20:57 auto_snap Hall A Tools snapshot
  • 356705 04/23/11 20:57 auto_snap Compton Polarimeter snapshot
  • 356704 04/23/11 20:57 auto_snap BigBox Magnet snapshot
  • 356703 04/23/11 20:57 auto_snap Beam Position and Current snapshot
  • 356702 04/23/11 20:22 xqiu target snapshot
  • 356701 04/23/11 20:14 miham Spot++ check for run #3784
  • 356700 04/23/11 20:13 miham Coda GUI screenshot for run #3784
  • 356699 04/23/11 20:11 a-onl Start_Run_3784,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, 3He 20cm Target, Beam current = 120 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356698 04/23/11 20:10 a-onl End_of_Run_3783
  • 356697 04/23/11 19:24 Jack (Techs) Sorry. Old news. Different probe placements.
  • 356696 04/23/11 19:15 miham Spot++ check for run #3783
  • 356695 04/23/11 19:14 miham Coda GUI screen shot for run #3783
  • 356694 04/23/11 19:13 a-onl Start_Run_3783,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, 3He 20cm Target, Beam current = 120 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356693 04/23/11 19:12 a-onl End_of_Run_3782
  • 356692 04/23/11 19:10 Jack (Techs) RQ1 and LQ1 field readings are different
  • 356691 04/23/11 18:21 miham Spot++ check for run #3782
  • 356690 04/23/11 18:19 miham Coda GUI screen shot for run #3782
  • 356689 04/23/11 18:18 a-onl Start_Run_3782,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, 3He 20cm Target, Beam current = 120 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356688 04/23/11 18:17 a-onl End_of_Run_3781
  • 356687 04/23/11 17:47 xqiu target status
  • 356686 04/23/11 17:28 miham Spot++ check for run #3781
  • 356685 04/23/11 17:28 xqiu target snapshot
  • 356684 04/23/11 17:27 miham Coda GUI screen shot for run #3781
  • 356683 04/23/11 17:27 a-onl Start_Run_3781,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, 3He 20cm Target, Beam current = 120 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356682 04/23/11 17:26 xqiu target snapshot
  • 356681 04/23/11 17:26 a-onl End_of_Run_3780
  • 356680 04/23/11 16:56 johna Still have beamline lecroy HV alarm (and turned speakers back on so we will start hearing alarms)
  • 356679 04/23/11 16:42 miham Spot++ check for run #3780
  • 356678 04/23/11 16:38 miham Coda GUI screen shot for run #3780
  • 356677 04/23/11 16:36 johna BCM stripchart for run 3779
  • 356676 04/23/11 16:35 a-onl Start_Run_3780,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, 3He 20cm Target, Beam current = 120 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356675 04/23/11 16:34 a-onl End_of_Run_3779
  • 356674 04/23/11 16:15 benesch 0R04 troubleshooting
  • 356673 04/23/11 16:05 adaq Beam Time Accounting
  • 356672 04/23/11 16:03 Solvignon Day Shift Summary
  • 356671 04/23/11 15:59 a-onl Start_Run_3779,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, 3He 20cm Target, Beam current = 120 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356670 04/23/11 15:58 a-onl End_of_Run_3778
  • 356669 04/23/11 15:48 Solvignon Stable beam at 120uA
  • 356668 04/23/11 15:38 pzhu target status
  • 356667 04/23/11 15:33 A. Camsonne Target vacuum yellow alarm
  • 356666 04/23/11 15:32 C. Gu Spot_L for Run#3778
  • 356665 04/23/11 15:29 a-onl Start_Run_3778,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Production SRC, 4He 20cm Target, Beam current = 120 uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356664 04/23/11 15:25 A. Camsonne High Power heater of He3 loop Hi Limit changed from 380W to 390 W
  • 356663 04/23/11 15:12 A. Camsonne Bogdan's detector rates at 5 uA
  • 356662 04/23/11 15:11 C. Gu spot_L result
  • 356661 04/23/11 15:10 a-onl End_of_Run_3777
  • 356660 04/23/11 15:08 a-onl Start_Run_3777,Run_type=Spot++,target_type=BeO,beam_e0=3.356,beam_raster=on,beam_current=5,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=Spot++ run,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356659 04/23/11 12:57 auto_snap Radiation and Ozone snapshot
  • 356658 04/23/11 12:57 auto_snap Miscellaneous snapshot
  • 356657 04/23/11 12:57 auto_snap Hall A Tools snapshot
  • 356656 04/23/11 12:57 auto_snap Compton Polarimeter snapshot
  • 356655 04/23/11 12:57 auto_snap BigBox Magnet snapshot
  • 356654 04/23/11 12:56 auto_snap Beam Position and Current snapshot
  • 356653 04/23/11 12:24 Solvignon Beam status update
  • 356652 04/23/11 11:51 Folts Lt Q1 power supply fault
  • 356651 04/23/11 11:00 Solvignon Left Q1 power supply problem: ask for controlled access
  • 356650 04/23/11 10:42 pzhu moved the target to empty
  • 356649 04/23/11 10:41 Solvignon Left Q1 cryo increasing slowly
  • 356648 04/23/11 10:29 Solvignon Left Q1 trip: low cryo
  • 356647 04/23/11 10:01 a-onl End_of_Run_3776
  • 356646 04/23/11 09:48 C. Gu Stop Run#3775 because of 15 min beam trip
  • 356645 04/23/11 09:40 a-onl Start_Run_3776,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=30,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=production run, He3 target 20cm, beam current is 30uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356644 04/23/11 09:31 a-onl End_of_Run_3775
  • 356643 04/23/11 09:21 C. Gu Spot_L for Run#3775
  • 356642 04/23/11 09:07 Solvignon Beam getting worse and current getting lower
  • 356641 04/23/11 08:58 a-onl Start_Run_3775,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=production run, He3 target 20cm, beam current is 110uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356640 04/23/11 08:57 a-onl End_of_Run_3774
  • 356639 04/23/11 08:02 adaq Beam Time Accounting
  • 356638 04/23/11 08:01 yez Owl Shift Summary
  • 356637 04/23/11 07:57 yez beam is not stable
  • 356636 04/23/11 07:29 a-onl End_of_Run_3772
  • 356635 04/23/11 07:25 a-onl Start_Run_3774,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=production run, He3 target 20cm, beam current is 110uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356634 04/23/11 07:17 a-onl End_of_Run_3773
  • 356633 04/23/11 06:40 a-onl Start_Run_3773,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=production run, He3 target 20cm, beam current is 110uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356632 04/23/11 06:35 gueye Run 3772 stopped - MCC running QE measurement
  • 356631 04/23/11 06:08 a-onl Start_Run_3772,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=production run, He3 target 20cm, beam current is 110uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356630 04/23/11 06:07 a-onl End_of_Run_3771
  • 356629 04/23/11 05:22 yez Start to run at 110uA since MCC has problems to delivery 120uA to us
  • 356628 04/23/11 05:22 a-onl Start_Run_3771,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=production run, He3 target 20cm, beam current is 110uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356627 04/23/11 05:12 gueye Run 3770 stopped: vacuum problem.
  • 356626 04/23/11 05:10 a-onl End_of_Run_3770
  • 356625 04/23/11 04:56 auto_snap Radiation and Ozone snapshot
  • 356624 04/23/11 04:56 auto_snap Miscellaneous snapshot
  • 356623 04/23/11 04:56 auto_snap Hall A Tools snapshot
  • 356622 04/23/11 04:56 auto_snap Compton Polarimeter snapshot
  • 356621 04/23/11 04:56 auto_snap BigBox Magnet snapshot
  • 356620 04/23/11 04:56 auto_snap Beam Position and Current snapshot
  • 356619 04/23/11 04:24 a-onl Start_Run_3770,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=production run, He3 target 20cm, beam current is 120uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356618 04/23/11 04:23 a-onl End_of_Run_3769
  • 356617 04/23/11 03:42 xbai Snapshot of Target Status
  • 356616 04/23/11 03:41 gueye Beam spotsize from spot_L
  • 356615 04/23/11 03:40 a-onl Start_Run_3769,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=production run, He3 target 20cm, beam current is 120uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356614 04/23/11 03:39 a-onl End_of_Run_3768
  • 356613 04/23/11 03:30 yez It looks like the red alarm is related to Compton
  • 356612 04/23/11 03:07 yez LeCory HV of Beamline becomes red but things look OK
  • 356611 04/23/11 03:04 gueye run 3767 shorten to check for beam position
  • 356610 04/23/11 02:59 a-onl Start_Run_3768,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=production run, He3 target 20cm, beam current is 120uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356609 04/23/11 02:58 a-onl End_of_Run_3767
  • 356608 04/23/11 02:50 a-onl Start_Run_3767,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=production run, He3 target 20cm, beam current is 120uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356607 04/23/11 02:49 a-onl End_of_Run_3766
  • 356606 04/23/11 02:49 yez Gas sensors alarm but went off immediately, HV alarm
  • 356605 04/23/11 02:17 yez Have to put more cryo coolant on He3 target
  • 356604 04/23/11 02:12 yez rates and scaler for He3 at 120uA
  • 356603 04/23/11 02:06 xbai Target move from Dummy 20cm to He3 20cm, Open JT_Valve
  • 356602 04/23/11 02:05 a-onl Start_Run_3766,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=production run, He3 target 20cm, beam current is 120uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356601 04/23/11 01:53 gueye target change
  • 356600 04/23/11 01:52 a-onl End_of_Run_3765
  • 356599 04/23/11 01:02 xbai Target Screen
  • 356598 04/23/11 00:53 yez rates and scalers
  • 356597 04/23/11 00:53 yez Reboot CODA becase TS0 lost communication
  • 356596 04/23/11 00:52 a-onl Start_Run_3765,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run, Dummy target, beam current is 40uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356595 04/23/11 00:45 a-onl End_of_Run_3763
  • 356594 04/23/11 00:06 adaq Beam Time Accounting
  • 356593 04/23/11 00:06 johna Swing shift summary
  • Friday

  • 356592 04/22/11 23:50 johna BCM for run 3762
  • 356591 04/22/11 23:49 miham Spot++ for run #3763
  • 356590 04/22/11 23:47 miham Coda GUI and Scalers Screen shots for run #3763
  • 356589 04/22/11 23:46 a-onl Start_Run_3763,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run, Dummy target, beam current is 40uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356588 04/22/11 23:45 a-onl End_of_Run_3762
  • 356587 04/22/11 22:49 miham Spot++ check for run #3762
  • 356586 04/22/11 22:47 miham Coda GUI and Scalers screen shot for run #3762
  • 356585 04/22/11 22:46 a-onl Start_Run_3762,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run, Dummy target, beam current is 40uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356584 04/22/11 22:43 a-onl End_of_Run_3761
  • 356583 04/22/11 22:04 miham Spot++ check for run #3761
  • 356582 04/22/11 21:45 miham Coda GUI and Scalers screen shots for run #3761(dummy target)
  • 356581 04/22/11 21:42 a-onl Start_Run_3761,Run_type=Production SRC,target_type=20cm dummy,beam_e0=3.356,beam_raster=on,beam_current=40,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run, Dummy target, beam current is 40uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356580 04/22/11 21:40 efuchey Target change
  • 356579 04/22/11 21:35 a-onl End_of_Run_3760
  • 356578 04/22/11 21:21 miham Spot++ check for run #3760
  • 356577 04/22/11 21:20 efuchey Actual target change time
  • 356576 04/22/11 21:19 miham Coda GUI and Scalers screen shot for run #3760
  • 356575 04/22/11 21:17 a-onl Start_Run_3760,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run, Carbon target, beam current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356574 04/22/11 21:16 efuchey Target screenshots
  • 356573 04/22/11 21:15 a-onl End_of_Run_3759
  • 356572 04/22/11 21:11 efuchey Target changed
  • 356571 04/22/11 20:56 auto_snap Radiation and Ozone snapshot
  • 356570 04/22/11 20:56 auto_snap Miscellaneous snapshot
  • 356569 04/22/11 20:56 auto_snap Hall A Tools snapshot
  • 356568 04/22/11 20:56 auto_snap Compton Polarimeter snapshot
  • 356567 04/22/11 20:56 auto_snap BigBox Magnet snapshot
  • 356566 04/22/11 20:56 auto_snap Beam Position and Current snapshot
  • 356565 04/22/11 20:52 miham Spot++ check for run #3759 (1st carbon run)
  • 356564 04/22/11 20:49 miham Coda GUI and Scalers screen shots for run #3759
  • 356563 04/22/11 20:46 a-onl Start_Run_3759,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run, Carbon target, beam current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356562 04/22/11 20:27 a-onl End_of_Run_3758
  • 356561 04/22/11 19:29 miham Spot++ check for run #3758
  • 356560 04/22/11 19:28 miham Coda GUI and Scalers screen shots for run #3758
  • 356559 04/22/11 19:25 a-onl Start_Run_3758,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,110A, Helium 3 target,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356558 04/22/11 19:13 miham Roc10 Error at the end of run #3757
  • 356557 04/22/11 19:10 a-onl End_of_Run_3757
  • 356556 04/22/11 18:30 miham Spot++ check for run #3757
  • 356555 04/22/11 18:27 miham Coda GUI and Scalers screen shots for run #3757
  • 356554 04/22/11 18:25 a-onl Start_Run_3757,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,110A, Helium 3 target,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356553 04/22/11 18:25 johna Hall C putting in the half wave plate
  • 356552 04/22/11 18:22 a-onl End_of_Run_3756
  • 356551 04/22/11 17:32 miham Spot++ check for run #3756
  • 356550 04/22/11 17:28 miham Coda GUI and Scaler screen shots for run #3756
  • 356549 04/22/11 17:26 jpchen changed fan speed from 36 Hz to 48 Hz on 3He loop
  • 356548 04/22/11 17:25 a-onl Start_Run_3756,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,110A, Helium 3 target,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356547 04/22/11 17:21 a-onl End_of_Run_3755
  • 356546 04/22/11 16:51 miham Spot++ check for run #3755
  • 356545 04/22/11 16:46 miham Coda GUI and Scaler screen shot for run #3755
  • 356544 04/22/11 16:42 a-onl Start_Run_3755,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,110A, Helium 3 target,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356543 04/22/11 16:40 a-onl End_of_Run_3754
  • 356542 04/22/11 16:23 adaq Beam Time Accounting
  • 356541 04/22/11 16:10 saw Day shift summary
  • 356540 04/22/11 15:42 a-onl Start_Run_3754,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,110A, Helium 3 target,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356539 04/22/11 15:40 a-onl End_of_Run_3753
  • 356538 04/22/11 14:52 saw spot++
  • 356537 04/22/11 14:46 a-onl Start_Run_3753,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,110A, Helium 3 target,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356536 04/22/11 14:34 saw Target work in hall done, request beam back
  • 356535 04/22/11 14:01 a-onl End_of_Run_3752
  • 356534 04/22/11 14:00 saw Request for escorted access
  • 356533 04/22/11 13:57 pzhu target status
  • 356532 04/22/11 13:52 saw Meekins studying target, I=50uA
  • 356531 04/22/11 13:50 pzhu He3Pressure
  • 356530 04/22/11 13:40 pzhu target status
  • 356529 04/22/11 13:25 a-onl Start_Run_3752,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,110A, Helium 3 target,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356528 04/22/11 13:24 a-onl End_of_Run_3751
  • 356527 04/22/11 13:24 saw Hall C changing IHWP from IN to OUT.
  • 356526 04/22/11 13:12 saw Beam current to 120uA
  • 356525 04/22/11 13:11 saw Beam current to 120uA
  • 356524 04/22/11 12:56 auto_snap Radiation and Ozone snapshot
  • 356523 04/22/11 12:56 auto_snap Miscellaneous snapshot
  • 356522 04/22/11 12:56 auto_snap Hall A Tools snapshot
  • 356521 04/22/11 12:55 auto_snap Compton Polarimeter snapshot
  • 356520 04/22/11 12:55 auto_snap BigBox Magnet snapshot
  • 356519 04/22/11 12:55 auto_snap Beam Position and Current snapshot
  • 356518 04/22/11 12:40 a-onl Start_Run_3751,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,110A, Helium 3 target,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356517 04/22/11 12:34 a-onl End_of_Run_3750
  • 356516 04/22/11 12:16 a-onl Start_Run_3750,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,110A, Helium 3 target,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356515 04/22/11 12:15 pzhu target status
  • 356514 04/22/11 12:14 a-onl End_of_Run_3749
  • 356513 04/22/11 12:07 saw spot++
  • 356512 04/22/11 11:58 a-onl Start_Run_3749,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,110A, Helium 3 target,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356511 04/22/11 11:41 meekins Hall A target
  • 356510 04/22/11 11:22 saw Meekins done in the hall
  • 356509 04/22/11 10:10 saw Power cycled hacweb1
  • 356508 04/22/11 09:58 saw hacweb1 crash?
  • 356507 04/22/11 09:56 saw Beam down for target work (density increase)
  • 356506 04/22/11 09:51 a-onl End_of_Run_3748
  • 356505 04/22/11 09:24 a-onl Start_Run_3748,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,110A, Helium 3 target,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356504 04/22/11 09:15 saw RF problems
  • 356503 04/22/11 09:11 a-onl End_of_Run_3747
  • 356502 04/22/11 08:53 a-onl Start_Run_3747,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,110A, Helium 3 target,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356501 04/22/11 08:46 a-onl End_of_Run_3746
  • 356500 04/22/11 08:11 adaq Beam Time Accounting
  • 356499 04/22/11 08:10 yez Owl Shift Summary
  • 356498 04/22/11 08:04 a-onl Start_Run_3746,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,110A,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356497 04/22/11 08:03 a-onl End_of_Run_3745
  • 356496 04/22/11 07:40 Jack (Techs) Right Q1 work done during the target changeover
  • 356495 04/22/11 07:23 yez Beam runs at 100uA during the last 15mins.
  • 356494 04/22/11 07:09 Folts right spectrometer movement restrictions
  • 356493 04/22/11 07:08 a-onl Start_Run_3745,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,110A,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356492 04/22/11 07:06 a-onl End_of_Run_3744
  • 356491 04/22/11 06:40 a-onl Start_Run_3744,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,110A,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356490 04/22/11 06:32 a-onl End_of_Run_3743
  • 356489 04/22/11 06:23 yez rates and scaler
  • 356488 04/22/11 06:23 yez target screen
  • 356487 04/22/11 06:02 a-onl Start_Run_3743,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,110A,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356486 04/22/11 06:02 a-onl End_of_Run_3742
  • 356485 04/22/11 05:52 yez beam is not stable and run at lower current
  • 356484 04/22/11 05:24 yez beam position was unlocked at the first 15mins of run#3742
  • 356483 04/22/11 05:17 yez spot++
  • 356482 04/22/11 05:05 a-onl Start_Run_3742,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,110A,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356481 04/22/11 05:02 yez L-HRS Dipole NMR locked
  • 356480 04/22/11 04:55 auto_snap Radiation and Ozone snapshot
  • 356479 04/22/11 04:55 auto_snap Miscellaneous snapshot
  • 356478 04/22/11 04:55 auto_snap Hall A Tools snapshot
  • 356477 04/22/11 04:55 auto_snap Compton Polarimeter snapshot
  • 356476 04/22/11 04:55 auto_snap BigBox Magnet snapshot
  • 356475 04/22/11 04:55 auto_snap Beam Position and Current snapshot
  • 356474 04/22/11 04:13 a-onl End_of_Run_3741
  • 356473 04/22/11 03:44 a-onl Start_Run_3741,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,110A,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356472 04/22/11 03:43 a-onl End_of_Run_3740
  • 356471 04/22/11 03:38 yez target screen
  • 356470 04/22/11 03:33 yez rates and sclers
  • 356469 04/22/11 02:54 yez beam was off for 15 mins but not it is back to 110uA
  • 356468 04/22/11 02:53 a-onl Start_Run_3740,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,110A,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356467 04/22/11 02:40 a-onl End_of_Run_3739
  • 356466 04/22/11 02:35 yez Beam was not stable at 110uA so we ask to send 100uA untill they fix the issue
  • 356465 04/22/11 02:31 a-onl Start_Run_3739,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=100,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,100A,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356464 04/22/11 02:30 a-onl End_of_Run_3738
  • 356463 04/22/11 02:17 a-onl Start_Run_3738,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,110uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356462 04/22/11 02:16 a-onl End_of_Run_3737
  • 356461 04/22/11 01:51 a-onl Start_Run_3737,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,110uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356460 04/22/11 01:50 a-onl End_of_Run_3736
  • 356459 04/22/11 01:49 yez Triggers and Target Y plots with PID cuts for run#3735
  • 356458 04/22/11 01:05 a-onl Start_Run_3736,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=110,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,110uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356457 04/22/11 01:03 a-onl End_of_Run_3735
  • 356456 04/22/11 00:29 yez target screen for 110uA
  • 356455 04/22/11 00:26 yez Run#3734 is for rate checking
  • 356454 04/22/11 00:25 yez Beam current for run 3735 is 110uA
  • 356453 04/22/11 00:23 adaq Beam Time Accounting
  • 356452 04/22/11 00:13 adaq Target Screens
  • 356451 04/22/11 00:12 Yi Qiang Beam current set at 110 uA
  • 356450 04/22/11 00:09 a-onl Start_Run_3735,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,120uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356449 04/22/11 00:07 Yi Qiang Swing Shift Summary
  • 356448 04/22/11 00:05 a-onl End_of_Run_3734
  • Thursday

  • 356447 04/21/11 23:59 a-onl Start_Run_3734,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=120,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=pruduction run,120uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356446 04/21/11 23:57 a-onl End_of_Run_3733
  • 356445 04/21/11 23:45 pzhu rate
  • 356444 04/21/11 23:42 pzhu raster,rate,spot++
  • 356443 04/21/11 23:40 a-onl Start_Run_3733,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=1,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=1uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356442 04/21/11 23:39 mhuang Alarm handler config file
  • 356441 04/21/11 23:29 a-onl End_of_Run_3732
  • 356440 04/21/11 23:27 a-onl Start_Run_3732,Run_type=Spot++,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=5,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=spot++ 5uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356439 04/21/11 23:26 a-onl End_of_Run_3731
  • 356438 04/21/11 23:25 Yi Qiang Harp scan shows normal beam size
  • 356437 04/21/11 23:24 pzhu spot++
  • 356436 04/21/11 23:22 a-onl Start_Run_3731,Run_type=Spot++,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=5,leftarm_p=2.995,leftarm_theta=25,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25,rightarm_sieve=out,comment_text=spot++ 5uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356435 04/21/11 23:22 Yi Qiang Beam is back to the Hall
  • 356434 04/21/11 23:21 mhuang Alarm Handler and Target Lifter
  • 356433 04/21/11 21:33 a-onl End_of_Run_3730
  • 356432 04/21/11 20:55 auto_snap Radiation and Ozone snapshot
  • 356431 04/21/11 20:55 auto_snap Miscellaneous snapshot
  • 356430 04/21/11 20:55 auto_snap Hall A Tools snapshot
  • 356429 04/21/11 20:55 auto_snap Compton Polarimeter snapshot
  • 356428 04/21/11 20:55 auto_snap BigBox Magnet snapshot
  • 356427 04/21/11 20:55 auto_snap Beam Position and Current snapshot
  • 356426 04/21/11 20:54 mhuang Target Screens
  • 356425 04/21/11 20:12 johna HRS-R Q3 setting
  • 356424 04/21/11 19:56 pzhu rate,spot++,raster
  • 356423 04/21/11 19:54 a-onl Start_Run_3730,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=1,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=production 1uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356422 04/21/11 19:36 mhuang Target moved to He3 20cm
  • 356421 04/21/11 19:31 pzhu raster
  • 356420 04/21/11 19:29 Yi Qiang Both HRS are ready for Kin 5.2
  • 356419 04/21/11 19:27 a-onl End_of_Run_3729
  • 356418 04/21/11 19:26 pzhu spot++ 3x3
  • 356417 04/21/11 19:26 pzhu spot++ 3x3
  • 356416 04/21/11 19:24 a-onl Start_Run_3729,Run_type=Spot++,target_type=BeO,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=Spot check,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356415 04/21/11 19:24 pzhu spot++
  • 356414 04/21/11 19:20 a-onl End_of_Run_3728
  • 356413 04/21/11 19:20 a-onl Start_Run_3728,Run_type=Spot++,target_type=BeO,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=Spot check,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356412 04/21/11 19:17 mhuang Target Screens
  • 356411 04/21/11 19:16 mhuang Target moved to BeO
  • 356410 04/21/11 19:12 a-onl End_of_Run_3727
  • 356409 04/21/11 18:16 Yi Qiang Power Permit, Target to Empty
  • 356408 04/21/11 17:14 Yi Qiang Both HRS moved to 25 degrees
  • 356407 04/21/11 16:53 a-onl Start_Run_3727,Run_type=Spot++,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=Right arm - Spot check,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356406 04/21/11 16:50 A. Camsonne Right arm BPM should be fine now
  • 356405 04/21/11 16:35 a-onl End_of_Run_3726
  • 356404 04/21/11 16:31 a-onl Start_Run_3726,Run_type=Spot++,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=Right arm - Spot check,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356403 04/21/11 16:30 M Jones Turned off HV crate hvqui_bb for Bigbite E plane
  • 356402 04/21/11 16:30 a-onl End_of_Run_3725
  • 356401 04/21/11 16:23 a-onl Start_Run_3725,Run_type=Spot++,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=Right arm - Spot check,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356400 04/21/11 16:22 a-onl End_of_Run_3724
  • 356399 04/21/11 16:20 a-onl Start_Run_3724,Run_type=Spot++,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=Right arm - Spot check,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356398 04/21/11 16:19 a-onl Start_Run_3723,Run_type=Spot++,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=Right arm - Spot check,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356397 04/21/11 16:19 a-onl End_of_Run_3722
  • 356395 04/21/11 16:15 johna Clarification on trigger types
  • 356394 04/21/11 15:59 hsmith Hall A check list
  • 356393 04/21/11 15:42 a-onl Start_Run_3722,Run_type=Spot++,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=Right arm - Spot check,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356392 04/21/11 15:42 a-onl End_of_Run_3721
  • 356391 04/21/11 15:24 a-onl Start_Run_3721,Run_type=Spot++,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=Right arm - Spot check,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356390 04/21/11 15:23 a-onl End_of_Run_3720
  • 356389 04/21/11 15:21 a-onl Start_Run_3720,Run_type=Spot++,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=Right arm - Spot check,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356388 04/21/11 15:21 a-onl End_of_Run_3719
  • 356387 04/21/11 15:14 a-onl Start_Run_3719,Run_type=Spot++,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=Right arm - Spot check,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356386 04/21/11 15:13 a-onl End_of_Run_3718
  • 356385 04/21/11 15:10 a-onl Start_Run_3718,Run_type=Spot++,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=Right arm - Spot check,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356384 04/21/11 15:09 a-onl End_of_Run_3717
  • 356383 04/21/11 15:07 a-onl Start_Run_3717,Run_type=Spot++,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=Right arm - Spot check,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356382 04/21/11 15:06 a-onl End_of_Run_3716
  • 356381 04/21/11 15:04 a-onl Start_Run_3716,Run_type=Spot++,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=Right arm - Spot check,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356380 04/21/11 15:03 a-onl End_of_Run_3715
  • 356379 04/21/11 15:01 a-onl Start_Run_3715,Run_type=Spot++,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=Right arm - Spot check,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356378 04/21/11 14:55 Fansler RQ2 N2
  • 356377 04/21/11 14:55 a-onl End_of_Run_3714
  • 356376 04/21/11 14:42 a-onl Start_Run_3714,Run_type=Spot++,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=Right arm - Spot check,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356375 04/21/11 14:33 yez T3,T3,T7 connected to R-HRS Scaler
  • 356374 04/21/11 14:30 V. Sulkosky E-plane HV On Lower Half
  • 356373 04/21/11 14:27 V. Sulkosky Trigger Types for x>2
  • 356372 04/21/11 14:22 V. Sulkosky Trigger Inputs to R-HRS Scalers
  • 356371 04/21/11 12:55 auto_snap Radiation and Ozone snapshot
  • 356370 04/21/11 12:55 auto_snap Miscellaneous snapshot
  • 356369 04/21/11 12:55 auto_snap Hall A Tools snapshot
  • 356368 04/21/11 12:55 auto_snap Compton Polarimeter snapshot
  • 356367 04/21/11 12:55 auto_snap BigBox Magnet snapshot
  • 356366 04/21/11 12:54 auto_snap Beam Position and Current snapshot
  • 356365 04/21/11 12:13 narbe (Momentum)Settings for HRS Magnets
  • 356364 04/21/11 11:43 Higinbotham Arms Set To 21 degrees and Quads Cycled
  • 356363 04/21/11 11:36 Higinbotham Spectrometer Angles Checked
  • 356362 04/21/11 04:54 auto_snap Radiation and Ozone snapshot
  • 356361 04/21/11 04:54 auto_snap Miscellaneous snapshot
  • 356360 04/21/11 04:54 auto_snap Hall A Tools snapshot
  • 356359 04/21/11 04:54 auto_snap Compton Polarimeter snapshot
  • 356358 04/21/11 04:54 auto_snap BigBox Magnet snapshot
  • 356357 04/21/11 04:54 auto_snap Beam Position and Current snapshot
  • Wednesday

  • 356356 04/20/11 23:25 a-onl End_of_Run_3713
  • 356355 04/20/11 22:49 a-onl Start_Run_3713,Run_type=Production SRC,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=Right arm ,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356354 04/20/11 22:13 R. Michaels Updated CODA version tried, rolled back for now.
  • 356353 04/20/11 22:10 a-onl End_of_Run_3712
  • 356352 04/20/11 22:08 a-onl Start_Run_3712,Run_type=Production SRC,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=DAQ test,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356351 04/20/11 22:03 a-onl End_of_Run_3711
  • 356350 04/20/11 21:58 a-onl Start_Run_3711,Run_type=Production SRC,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=DAQ test,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356349 04/20/11 21:53 a-onl End_of_Run_3710
  • 356348 04/20/11 21:51 a-onl Start_Run_3710,Run_type=Production SRC,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=DAQ test,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356347 04/20/11 21:46 a-onl Start_Run_3709,Run_type=Production SRC,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=DAQ test,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356346 04/20/11 20:54 auto_snap Radiation and Ozone snapshot
  • 356345 04/20/11 20:54 auto_snap Miscellaneous snapshot
  • 356344 04/20/11 20:54 auto_snap Hall A Tools snapshot
  • 356343 04/20/11 20:54 auto_snap Compton Polarimeter snapshot
  • 356342 04/20/11 20:54 auto_snap BigBox Magnet snapshot
  • 356341 04/20/11 20:54 auto_snap Beam Position and Current snapshot
  • 356340 04/20/11 15:27 fansler LQ3 N2
  • 356339 04/20/11 12:54 auto_snap Radiation and Ozone snapshot
  • 356338 04/20/11 12:54 auto_snap Miscellaneous snapshot
  • 356337 04/20/11 12:54 auto_snap Hall A Tools snapshot
  • 356336 04/20/11 12:54 auto_snap Compton Polarimeter snapshot
  • 356335 04/20/11 12:54 auto_snap BigBox Magnet snapshot
  • 356334 04/20/11 12:53 auto_snap Beam Position and Current snapshot
  • 356333 04/20/11 04:53 auto_snap Radiation and Ozone snapshot
  • 356332 04/20/11 04:53 auto_snap Miscellaneous snapshot
  • 356331 04/20/11 04:53 auto_snap Hall A Tools snapshot
  • 356330 04/20/11 04:53 auto_snap Compton Polarimeter snapshot
  • 356329 04/20/11 04:53 auto_snap BigBox Magnet snapshot
  • 356328 04/20/11 04:53 auto_snap Beam Position and Current snapshot
  • Tuesday

  • 356327 04/19/11 20:53 auto_snap Radiation and Ozone snapshot
  • 356326 04/19/11 20:53 auto_snap Miscellaneous snapshot
  • 356325 04/19/11 20:53 auto_snap Hall A Tools snapshot
  • 356324 04/19/11 20:53 auto_snap Compton Polarimeter snapshot
  • 356323 04/19/11 20:53 auto_snap BigBox Magnet snapshot
  • 356322 04/19/11 20:53 auto_snap Beam Position and Current snapshot
  • 356321 04/19/11 12:53 auto_snap Radiation and Ozone snapshot
  • 356320 04/19/11 12:53 auto_snap Miscellaneous snapshot
  • 356319 04/19/11 12:53 auto_snap Hall A Tools snapshot
  • 356318 04/19/11 12:53 auto_snap Compton Polarimeter snapshot
  • 356317 04/19/11 12:53 auto_snap BigBox Magnet snapshot
  • 356316 04/19/11 12:53 auto_snap Beam Position and Current snapshot
  • 356315 04/19/11 11:54 V. Sulkosky /adaql10/work1 Continued clean up (Better format)
  • 356314 04/19/11 11:53 V. Sulkosky /adaql10/work1 Continued clean up
  • 356313 04/19/11 10:38 Fansler Lead Flow/ Vacuum
  • 356312 04/19/11 08:30 R. Michaels re: run 3647 (it was JUNK)
  • 356311 04/19/11 08:02 adaq Beam Time Accounting
  • 356310 04/19/11 08:00 efuchey 4/19/2011 Owl shift summary
  • 356309 04/19/11 07:43 efuchey switching to restricted access
  • 356308 04/19/11 07:31 efuchey Beam taken away by MCC
  • 356307 04/19/11 06:47 a-onl End_of_Run_3708
  • 356306 04/19/11 06:07 a-onl Start_Run_3708,Run_type=Production SRC,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=kinematic 4.2 on 4cm dummy LHRS,R-HRS,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356305 04/19/11 06:06 a-onl End_of_Run_3707
  • 356304 04/19/11 05:07 efuchey R-HRS momentum change
  • 356303 04/19/11 05:06 a-onl Start_Run_3707,Run_type=Production SRC,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=kinematic 4.2 on 4cm dummy LHRS,R-HRS,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356302 04/19/11 05:05 a-onl End_of_Run_3706
  • 356301 04/19/11 05:02 a-onl Start_Run_3706,Run_type=Production SRC,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=kinematic 4.2 on 4cm dummy LHRS,R-HRS,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356300 04/19/11 05:02 a-onl End_of_Run_3705
  • 356299 04/19/11 04:52 auto_snap Radiation and Ozone snapshot
  • 356298 04/19/11 04:52 auto_snap Miscellaneous snapshot
  • 356297 04/19/11 04:52 auto_snap Hall A Tools snapshot
  • 356296 04/19/11 04:52 auto_snap Compton Polarimeter snapshot
  • 356295 04/19/11 04:52 auto_snap BigBox Magnet snapshot
  • 356294 04/19/11 04:52 auto_snap Beam Position and Current snapshot
  • 356293 04/19/11 04:29 yez Summary of R-HRS Optics Data taking during this owl shift
  • 356292 04/19/11 04:18 efuchey HV screenshot
  • 356291 04/19/11 04:12 efuchey Hall A General tools
  • 356290 04/19/11 03:37 a-onl Start_Run_3705,Run_type=Production SRC,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.522,rightarm_theta=23,rightarm_sieve=out,comment_text=kinematic 4.2 on 4cm dummy LHRS,R-HRS is not ready,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356289 04/19/11 03:35 a-onl End_of_Run_3704
  • 356288 04/19/11 03:14 efuchey R-HRS momentum change
  • 356287 04/19/11 03:14 a-onl Start_Run_3704,Run_type=Optics,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.600,rightarm_theta=23,rightarm_sieve=out,comment_text=kinematic 4.2 on 4cm dummy LHRS,R-HRS deltaP scan -3%,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356286 04/19/11 03:12 a-onl End_of_Run_3703
  • 356285 04/19/11 03:02 a-onl Start_Run_3703,Run_type=Production SRC,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=10,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.600,rightarm_theta=23,rightarm_sieve=out,comment_text=kinematic 4.2 4cm dummy LHRS,R-HRS is not ready for delta scan,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356284 04/19/11 03:01 a-onl End_of_Run_3702
  • 356283 04/19/11 02:42 a-onl Start_Run_3702,Run_type=Production SRC,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.600,rightarm_theta=23,rightarm_sieve=out,comment_text=kinematic 4.2 4cm dummy LHRS,R-HRS is not ready for delta scan ,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356282 04/19/11 02:39 kelleher Target to Dummy
  • 356281 04/19/11 02:37 a-onl End_of_Run_3701
  • 356280 04/19/11 02:16 yez Raster was off during run#3699
  • 356279 04/19/11 02:15 a-onl Start_Run_3701,Run_type=Optics,target_type=Optics: multi 1mm C foils,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.600,rightarm_theta=23,rightarm_sieve=out,comment_text=optics RHRS, Vertex run with Multi-C; kinematic 4.2 4cm dummy LHRS,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356278 04/19/11 02:12 a-onl End_of_Run_3700
  • 356277 04/19/11 01:50 a-onl Start_Run_3700,Run_type=Optics,target_type=Optics: multi 1mm C foils,beam_e0=3.356,beam_raster=off,beam_current=95,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.600,rightarm_theta=23,rightarm_sieve=out,comment_text=optics RHRS, Vertex run with Multi-C; kinematic 4.2 4cm dummy LHRS,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356276 04/19/11 01:48 kelleher Target Optics
  • 356275 04/19/11 01:46 a-onl End_of_Run_3699
  • 356274 04/19/11 01:28 yez run#3698 is for momentum scan with dP=+0%
  • 356273 04/19/11 01:26 yez run#3695 is for momentum scan with dP=+3%
  • 356272 04/19/11 01:26 a-onl Start_Run_3699,Run_type=Optics,target_type=BeO,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.600,rightarm_theta=23,rightarm_sieve=out,comment_text=optics RHRS, Vertex run with BeO; kinematic 4.2 4cm dummy LHRS,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356271 04/19/11 01:22 kelleher Target to BeO
  • 356270 04/19/11 01:21 a-onl End_of_Run_3698
  • 356269 04/19/11 01:02 a-onl Start_Run_3698,Run_type=Optics,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.600,rightarm_theta=23,rightarm_sieve=out,comment_text=optics RHRS; kinematic 4.2 4cm dummy LHRS,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356268 04/19/11 01:01 a-onl End_of_Run_3697
  • 356267 04/19/11 00:59 efuchey R-HRS momentum change
  • 356266 04/19/11 00:58 a-onl Start_Run_3697,Run_type=Production SRC,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.678,rightarm_theta=23,rightarm_sieve=out,comment_text=optics RHRS; kinematic 4.2 4cm dummy LHRS, RHRS is changing momentum,junk,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356265 04/19/11 00:36 a-onl End_of_Run_3696
  • 356264 04/19/11 00:19 a-onl Start_Run_3696,Run_type=Production SRC,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.678,rightarm_theta=23,rightarm_sieve=out,comment_text=optics RHRS; kinematic 4.2 4cm dummy LHRS, RHRS is changing momentum,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356263 04/19/11 00:16 a-onl End_of_Run_3695
  • 356262 04/19/11 00:08 V. Sulkosky Swing Shift Summary April 18, 2011
  • 356261 04/19/11 00:07 adaq Beam Time Accounting
  • Monday

  • 356260 04/18/11 23:58 Hisham Target is warmed up.
  • 356259 04/18/11 23:56 a-onl Start_Run_3695,Run_type=Optics,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.678,rightarm_theta=23,rightarm_sieve=out,comment_text=optics RHRS; kinematic 4.2 4cm dummy LHRS,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356258 04/18/11 23:54 a-onl End_of_Run_3694
  • 356257 04/18/11 23:47 Hisham Target status
  • 356256 04/18/11 23:47 V. Sulkosky Hall A Tools New R-HRS Setting
  • 356255 04/18/11 23:45 V. Sulkosky Right HRS Dipole NMR Locked
  • 356254 04/18/11 23:25 a-onl Start_Run_3694,Run_type=Production SRC,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=2.678,rightarm_theta=23,rightarm_sieve=out,comment_text=production kinematic 4.2 4cm dummy,R-HRS is not ready,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356253 04/18/11 22:50 V. Sulkosky Controlled Access for Target Warm Up
  • 356252 04/18/11 22:25 R. Michaels story of runs 3655, 3566, 3674, 3675 files
  • 356251 04/18/11 22:03 a-onl End_of_Run_3693
  • 356250 04/18/11 21:55 a-onl Start_Run_3693,Run_type=Production SRC,target_type=10cm dummy,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=3.035,rightarm_theta=23,rightarm_sieve=out,comment_text=production kinematic 4.2 4cm dummy,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356249 04/18/11 21:48 a-onl End_of_Run_3692
  • 356248 04/18/11 21:30 a-onl Start_Run_3692,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=3.035,rightarm_theta=23,rightarm_sieve=out,comment_text=production kinematic 4.2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356247 04/18/11 21:28 Hisham Trget moved to Carbon 5mm
  • 356246 04/18/11 21:20 a-onl End_of_Run_3691
  • 356245 04/18/11 20:52 auto_snap Radiation and Ozone snapshot
  • 356244 04/18/11 20:52 auto_snap Miscellaneous snapshot
  • 356243 04/18/11 20:52 auto_snap Hall A Tools snapshot
  • 356242 04/18/11 20:52 auto_snap Compton Polarimeter snapshot
  • 356241 04/18/11 20:52 auto_snap BigBox Magnet snapshot
  • 356240 04/18/11 20:52 auto_snap Beam Position and Current snapshot
  • 356239 04/18/11 20:45 V. Sulkosky Spot++ for 3691
  • 356238 04/18/11 20:43 V. Sulkosky Spot++ 3689
  • 356237 04/18/11 20:35 V. Sulkosky Hall A Tools Kin 4.2
  • 356236 04/18/11 20:32 a-onl Start_Run_3691,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=3.035,rightarm_theta=23,rightarm_sieve=out,comment_text=production kinematic 4.2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356235 04/18/11 20:31 a-onl End_of_Run_3690
  • 356234 04/18/11 20:04 R. Michaels re: moving filenames
  • 356233 04/18/11 19:44 a-onl Start_Run_3690,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=3.035,rightarm_theta=23,rightarm_sieve=out,comment_text=production kinematic 4.2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356232 04/18/11 19:43 a-onl End_of_Run_3689
  • 356231 04/18/11 19:43 a-onl Start_Run_3689,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=95,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=3.035,rightarm_theta=23,rightarm_sieve=out,comment_text=production kinematic 4.2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356230 04/18/11 19:38 a-onl End_of_Run_3688
  • 356229 04/18/11 19:34 a-onl Start_Run_3688,Run_type=Spot++,target_type=4He 20cm cell,beam_e0=3.356,beam_raster=on,beam_current=1,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=3.035,rightarm_theta=23,rightarm_sieve=out,comment_text=SPOT++,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356228 04/18/11 19:31 Hisham Target moved to He4
  • 356227 04/18/11 19:28 V. Sulkosky Raster Check 3687
  • 356226 04/18/11 19:28 yez CODA crashed during run#3647 hence we lose entire data of this run.
  • 356225 04/18/11 19:24 a-onl End_of_Run_3687
  • 356224 04/18/11 19:20 a-onl Start_Run_3687,Run_type=Spot++,target_type=BeO,beam_e0=3.356,beam_raster=on,beam_current=5,leftarm_p=3.035,leftarm_theta=23,leftarm_sieve=out,rightarm_p=3.035,rightarm_theta=23,rightarm_sieve=out,comment_text=SPOT++,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356223 04/18/11 19:04 adaq Beam Time Accounting
  • 356222 04/18/11 19:03 V. Sulkosky Controlled Access
  • 356221 04/18/11 18:44 yez Rename file names of raw data for #3656 and #3675
  • 356220 04/18/11 18:23 V. Sulkosky BCM Log File Location
  • 356219 04/18/11 18:19 V. Sulkosky BCM Calibration Info
  • 356218 04/18/11 18:14 a-onl End_of_Run_3686
  • 356217 04/18/11 17:37 V. Sulkosky Hall A Tools Kin 4.2
  • 356216 04/18/11 17:36 V. Sulkosky Left NMR Locked Kin 4.2
  • 356215 04/18/11 17:27 a-onl Start_Run_3686,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=Production on LD2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356214 04/18/11 17:04 a-onl End_of_Run_3685
  • 356213 04/18/11 16:53 V. Sulkosky BCM Calibration
  • 356212 04/18/11 16:52 a-onl Start_Run_3685,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=Production on LD2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356211 04/18/11 16:50 a-onl End_of_Run_3684
  • 356210 04/18/11 16:49 a-onl Start_Run_3684,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=Production on LD2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356209 04/18/11 16:45 Hisham Target status
  • 356208 04/18/11 16:25 johna Setting HRS momenta for kinematics 4.2
  • 356207 04/18/11 16:07 gaskell Day shift summary: 04/18/2011
  • 356206 04/18/11 16:02 brads LHRS back angle camera has been bumped, not useful
  • 356205 04/18/11 15:59 brads HRS camera angles also stored in CODA file
  • 356204 04/18/11 15:53 gaskell Kin 3.1 complete - moving on to BCM calibration procedure
  • 356203 04/18/11 15:49 a-onl End_of_Run_3683
  • 356202 04/18/11 15:20 a-onl Start_Run_3683,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=Production on LD2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356201 04/18/11 15:19 a-onl End_of_Run_3682
  • 356200 04/18/11 15:09 a-onl Start_Run_3682,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=Production on LD2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356199 04/18/11 15:04 a-onl End_of_Run_3681
  • 356198 04/18/11 15:00 gaskell Hall A Tools
  • 356197 04/18/11 14:51 a-onl Start_Run_3681,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=Production on LD2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356196 04/18/11 14:51 V. Sulkosky Re: Run Number glitched again for run 3675, rename raw data file of 3674 -- better format
  • 356195 04/18/11 14:50 a-onl End_of_Run_3680
  • 356194 04/18/11 14:39 a-onl Start_Run_3680,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=Production on LD2,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356193 04/18/11 14:37 johna Non-blurred verson of the HRS-L angle camera
  • 356192 04/18/11 14:37 a-onl End_of_Run_3679
  • 356191 04/18/11 14:31 kelleher target to LD2
  • 356190 04/18/11 14:08 johna Right HRS angle camera snapshots, kin3
  • 356189 04/18/11 14:05 johna Right HRS angle camera snapshots, kin3
  • 356188 04/18/11 14:04 johna Left HRS angle camera snapshots, kin 3
  • 356187 04/18/11 13:28 a-onl Start_Run_3679,Run_type=Optics,target_type=Optics: multi 1mm C foils,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=Optics,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356186 04/18/11 13:27 yez Run Number glitched again for run 3675, rename raw data file of 3674 -- better format
  • 356185 04/18/11 13:27 yez Run Number glitched again for run 3675, rename raw data file of 3674
  • 356184 04/18/11 13:26 a-onl Start_Run_3678,Run_type=Other,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=JUNK,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356183 04/18/11 13:26 a-onl End_of_Run_3678
  • 356182 04/18/11 13:20 gaskell done with 4He - changing to optics target
  • 356181 04/18/11 13:19 kelleher target to optics
  • 356180 04/18/11 13:19 kelleher Target to helium
  • 356179 04/18/11 13:17 a-onl End_of_Run_3677
  • 356178 04/18/11 13:13 a-onl Start_Run_3677,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356177 04/18/11 13:12 a-onl End_of_Run_3676
  • 356176 04/18/11 13:02 a-onl Start_Run_3676,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356175 04/18/11 13:01 a-onl End_of_Run_3675
  • 356174 04/18/11 12:52 auto_snap Radiation and Ozone snapshot
  • 356173 04/18/11 12:52 auto_snap Miscellaneous snapshot
  • 356172 04/18/11 12:52 auto_snap Hall A Tools snapshot
  • 356171 04/18/11 12:52 auto_snap Compton Polarimeter snapshot
  • 356170 04/18/11 12:52 auto_snap BigBox Magnet snapshot
  • 356169 04/18/11 12:52 auto_snap Beam Position and Current snapshot
  • 356168 04/18/11 12:42 a-onl Start_Run_3675,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356167 04/18/11 12:41 a-onl End_of_Run_3674
  • 356166 04/18/11 12:26 a-onl Start_Run_3674,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356165 04/18/11 12:25 a-onl End_of_Run_3673
  • 356164 04/18/11 12:24 gaskell spot++ for BeO run
  • 356163 04/18/11 12:21 a-onl Start_Run_3673,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356162 04/18/11 12:09 a-onl End_of_Run_3672
  • 356161 04/18/11 11:55 a-onl Start_Run_3672,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356160 04/18/11 11:54 elli Prescales changed
  • 356159 04/18/11 11:54 a-onl End_of_Run_3671
  • 356158 04/18/11 11:51 a-onl Start_Run_3671,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356157 04/18/11 11:48 kelleher Target to BeO
  • 356156 04/18/11 11:46 a-onl End_of_Run_3670
  • 356155 04/18/11 11:35 a-onl Start_Run_3670,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356154 04/18/11 11:34 a-onl End_of_Run_3669
  • 356153 04/18/11 11:27 a-onl Start_Run_3669,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356152 04/18/11 11:26 a-onl End_of_Run_3668
  • 356151 04/18/11 11:17 gaskell Beam back - energy ~ 3 MeV higher
  • 356150 04/18/11 11:16 a-onl Start_Run_3668,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356149 04/18/11 11:02 gaskell MCC will adjust dipole string
  • 356148 04/18/11 10:58 a-onl End_of_Run_3667
  • 356147 04/18/11 10:56 ellie Prescales Changed
  • 356146 04/18/11 10:53 a-onl Start_Run_3667,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356145 04/18/11 10:52 a-onl End_of_Run_3666
  • 356144 04/18/11 10:50 a-onl Start_Run_3666,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356143 04/18/11 10:48 a-onl End_of_Run_3665
  • 356142 04/18/11 10:46 gaskell LHRS NMR unlocked
  • 356141 04/18/11 10:45 a-onl Start_Run_3665,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356140 04/18/11 10:43 a-onl End_of_Run_3664
  • 356139 04/18/11 10:42 a-onl Start_Run_3664,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.905,leftarm_theta=21,leftarm_sieve=out,rightarm_p=2.905,rightarm_theta=21,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356138 04/18/11 10:40 LeRose Left NMR
  • 356137 04/18/11 10:10 gaskell LHRS dipole NMR unhappy
  • 356136 04/18/11 09:42 gaskell Finished kin 3.2 --> going to kin 3.1
  • 356135 04/18/11 09:30 a-onl End_of_Run_3663
  • 356134 04/18/11 09:25 a-onl Start_Run_3663,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356133 04/18/11 09:24 a-onl End_of_Run_3662
  • 356132 04/18/11 09:16 a-onl Start_Run_3662,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356131 04/18/11 09:01 kelleher Target to Carbon
  • 356130 04/18/11 08:53 a-onl End_of_Run_3661
  • 356129 04/18/11 08:21 a-onl Start_Run_3661,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356128 04/18/11 08:20 a-onl End_of_Run_3660
  • 356127 04/18/11 08:05 adaq Beam Time Accounting
  • 356126 04/18/11 08:05 wdconinc Owl shift summary
  • 356125 04/18/11 07:40 a-onl Start_Run_3660,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356124 04/18/11 07:37 a-onl Start_Run_3659,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356123 04/18/11 07:36 a-onl End_of_Run_3658
  • 356122 04/18/11 07:21 a-onl Start_Run_3658,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356121 04/18/11 07:20 a-onl End_of_Run_3657
  • 356120 04/18/11 07:12 wdconinc Run 3657: R-D1 still oscillating around set point
  • 356119 04/18/11 07:09 a-onl Start_Run_3657,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=3.055,leftarm_theta=21,leftarm_sieve=out,rightarm_p=3.055,rightarm_theta=21,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356118 04/18/11 06:53 Higinbotham Spectrometer Moved To 21 Degrees
  • 356117 04/18/11 06:25 wdconinc Controlled access for spectrometer move
  • 356116 04/18/11 06:25 Higinbotham Setting Moment for Kin 3.2
  • 356115 04/18/11 06:13 a-onl End_of_Run_3656
  • 356114 04/18/11 06:11 bittnerj/wdc Run number glitch: backed up original 3655
  • 356113 04/18/11 05:56 a-onl Start_Run_3656,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356112 04/18/11 05:55 a-onl End_of_Run_3655
  • 356111 04/18/11 05:29 a-onl Start_Run_3655,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356110 04/18/11 05:28 a-onl End_of_Run_3654
  • 356109 04/18/11 05:09 wdconinc General tools
  • 356108 04/18/11 04:51 auto_snap Radiation and Ozone snapshot
  • 356107 04/18/11 04:51 auto_snap Miscellaneous snapshot
  • 356106 04/18/11 04:51 auto_snap Hall A Tools snapshot
  • 356105 04/18/11 04:51 auto_snap Compton Polarimeter snapshot
  • 356104 04/18/11 04:51 auto_snap BigBox Magnet snapshot
  • 356103 04/18/11 04:51 auto_snap Beam Position and Current snapshot
  • 356102 04/18/11 04:42 bittnerj empty vertex solved
  • 356101 04/18/11 04:36 a-onl Start_Run_3654,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356100 04/18/11 04:35 a-onl End_of_Run_3653
  • 356099 04/18/11 04:14 bittnerj/wdc Problem with L.tr.vz for some runs only
  • 356098 04/18/11 03:48 bittnerj run 3646 target vertex reconstruction problem
  • 356097 04/18/11 03:34 a-onl Start_Run_3653,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356096 04/18/11 03:33 a-onl End_of_Run_3652
  • 356095 04/18/11 03:15 wdconinc Target screen shots
  • 356094 04/18/11 02:54 wdconinc Target: Parked LD2, active LHe3
  • 356093 04/18/11 02:51 a-onl Start_Run_3652,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356092 04/18/11 02:50 wdconinc Beam: 95uA on Helium target
  • 356091 04/18/11 02:39 wdconinc Moving target to Helium 20cm (loop 1)
  • 356090 04/18/11 02:38 a-onl End_of_Run_3651
  • 356089 04/18/11 01:58 a-onl Start_Run_3651,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356088 04/18/11 01:57 a-onl End_of_Run_3650
  • 356087 04/18/11 01:36 wdconinc Correction: Run *3649* ended with EB2 crash
  • 356086 04/18/11 01:32 a-onl Start_Run_3650,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356085 04/18/11 01:31 wdconinc Run 3648 ended with EB2 crash
  • 356084 04/18/11 01:24 bittnerj eb2 crashed
  • 356083 04/18/11 01:21 wdconinc Starting closing JT2, followed by opening JT1
  • 356082 04/18/11 01:17 a-onl Start_Run_3649,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production, current is 95uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356081 04/18/11 01:12 wdconinc Beam: 95uA on Carbon
  • 356080 04/18/11 01:11 bittnerj roc5 message
  • 356079 04/18/11 01:08 wdconinc Moving target to Carbon (5mm
  • 356078 04/18/11 01:07 a-onl End_of_Run_3648
  • 356077 04/18/11 00:41 wdconinc Owl shift start & plans
  • 356076 04/18/11 00:30 S. Sirca ROCs rebooted
  • 356075 04/18/11 00:28 a-onl Start_Run_3648,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=45,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production, current is 45uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356074 04/18/11 00:07 yez Doesn't this run should be Optics run?
  • 356073 04/18/11 00:04 a-onl Start_Run_3647,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=45,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production, current is 45uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356072 04/18/11 00:04 S. Sirca Swing shift SUMMARY
  • 356071 04/18/11 00:03 a-onl End_of_Run_3646
  • 356070 04/18/11 00:02 adaq Beam Time Accounting
  • Sunday

  • 356069 04/17/11 23:55 pzhu target status
  • 356068 04/17/11 23:09 miham Spot++ check for run #3646
  • 356067 04/17/11 23:08 miham Coda GUI screen shot for run #3646
  • 356066 04/17/11 23:07 a-onl Start_Run_3646,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=45,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production, current is 45uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356065 04/17/11 23:05 a-onl End_of_Run_3645
  • 356064 04/17/11 22:52 miham Begin-of-Run Missing for run #3645
  • 356063 04/17/11 22:21 pzhu target status 45uA
  • 356062 04/17/11 22:07 miham Spot++ check for run #3645
  • 356061 04/17/11 22:06 miham Coda GUI screen shot for run #3645
  • 356060 04/17/11 22:03 a-onl End_of_Run_3644
  • 356059 04/17/11 21:35 yez Golden Run for Kin5.1 is set to be run3643 with 40K events
  • 356058 04/17/11 21:03 miham Spot++ check for run #3644
  • 356057 04/17/11 21:02 S. Sirca HRSR D1 seems to have settled in
  • 356056 04/17/11 21:02 miham Coda GUI screen shot for run #3644
  • 356055 04/17/11 21:00 a-onl Start_Run_3644,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=45,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production, current is 45uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356054 04/17/11 20:59 a-onl End_of_Run_3643
  • 356053 04/17/11 20:51 auto_snap Radiation and Ozone snapshot
  • 356052 04/17/11 20:51 auto_snap Miscellaneous snapshot
  • 356051 04/17/11 20:51 auto_snap Hall A Tools snapshot
  • 356050 04/17/11 20:51 auto_snap Compton Polarimeter snapshot
  • 356049 04/17/11 20:51 auto_snap BigBox Magnet snapshot
  • 356048 04/17/11 20:51 auto_snap Beam Position and Current snapshot
  • 356047 04/17/11 20:45 yez It is a known feature
  • 356046 04/17/11 20:40 miham Spot++ check for run #3643
  • 356045 04/17/11 20:39 miham Coda GUI screen shot for run #3643
  • 356044 04/17/11 20:37 a-onl Start_Run_3643,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=45,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production, current is 45uA, HRS-R Dipole still ramping up.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356043 04/17/11 20:35 a-onl End_of_Run_3642
  • 356042 04/17/11 20:35 miham Coda GUI screen shot for run #3642
  • 356041 04/17/11 20:34 pzhu target status
  • 356040 04/17/11 20:30 a-onl Start_Run_3642,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=45,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production, current is 45uA, HRS-R Dipole still ramping up.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356039 04/17/11 20:08 S. Sirca HRSL D1 NMR loses lock frequently
  • 356038 04/17/11 19:56 S. Sirca HRSL OK for kinematics 5.1
  • 356037 04/17/11 19:35 a-onl End_of_Run_3641
  • 356036 04/17/11 19:33 pzhu target status
  • 356035 04/17/11 19:15 miham Spot++ check for run #3641
  • 356034 04/17/11 19:13 miham Coda GUI screen shot for run #3641
  • 356033 04/17/11 19:12 a-onl Start_Run_3641,Run_type=Production SRC,target_type=Optics: multi 1mm C foils,beam_e0=3.353,beam_raster=on,beam_current=5,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Optics, current is 5uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356032 04/17/11 19:02 a-onl End_of_Run_3640
  • 356031 04/17/11 18:59 narbe Update on xgt2
  • 356030 04/17/11 18:56 miham Spot++ check for run #3640
  • 356029 04/17/11 18:52 miham Coda GUI screen shot for run #3640
  • 356028 04/17/11 18:49 a-onl Start_Run_3640,Run_type=Production SRC,target_type=BeO,beam_e0=3.353,beam_raster=on,beam_current=5,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Calibration, current is 5uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356027 04/17/11 18:37 S. Sirca Why does the Loop 2 heater fluctuate so much?
  • 356026 04/17/11 18:37 a-onl End_of_Run_3639
  • 356025 04/17/11 18:31 miham Spot++ check for run #3639
  • 356024 04/17/11 18:29 miham Coda GUI screen shot for run #3639
  • 356023 04/17/11 18:27 a-onl Start_Run_3639,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=45,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling, current is 45uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356022 04/17/11 18:21 pzhu target status in 30uA
  • 356021 04/17/11 18:21 a-onl End_of_Run_3638
  • 356020 04/17/11 18:20 S. Sirca Max current on LD2 shall be 40uA
  • 356019 04/17/11 18:17 A. Camsonne Reduced JT on Loop 1 for more power on LD2 target
  • 356018 04/17/11 18:13 miham Spot++ check for run #3638
  • 356017 04/17/11 18:12 miham Coda GUI screen shot for run #3638
  • 356016 04/17/11 18:10 a-onl Start_Run_3638,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=30,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling, current is 30uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356015 04/17/11 18:07 a-onl End_of_Run_3637
  • 356014 04/17/11 17:40 miham Coda GUI screen shot for run #3637
  • 356013 04/17/11 17:39 miham Spot++ check for run #3637
  • 356012 04/17/11 17:37 a-onl Start_Run_3637,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=15,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling, current is 15uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356011 04/17/11 17:35 a-onl End_of_Run_3636
  • 356010 04/17/11 17:31 a-onl Start_Run_3636,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=15,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling, current is 15uA, we are twicking target during this run.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356009 04/17/11 17:13 a-onl End_of_Run_3635
  • 356008 04/17/11 17:12 a-onl Start_Run_3635,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=60,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling, current is 60uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356007 04/17/11 17:01 a-onl End_of_Run_3634
  • 356006 04/17/11 17:00 miham High gain in one of the channels in S2 plane
  • 356005 04/17/11 16:56 A. Camsonne Large offset for unser
  • 356004 04/17/11 16:55 miham Coda GUI screen shot for run #3634
  • 356003 04/17/11 16:54 a-onl Start_Run_3634,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=15,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling, current is 15uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 356002 04/17/11 16:52 a-onl End_of_Run_3633
  • 356001 04/17/11 16:43 miham Spot++ check for run #3633
  • 356000 04/17/11 16:42 miham Coda GUI screen shot for run #3633
  • 355999 04/17/11 16:41 a-onl Start_Run_3633,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=15,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling, current is 15uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355998 04/17/11 16:39 a-onl End_of_Run_3632
  • 355997 04/17/11 16:34 miham Coda GUI screen shot for run #3632
  • 355996 04/17/11 16:29 miham Spot++ check for run #3632
  • 355995 04/17/11 16:28 a-onl Start_Run_3632,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=30,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling, current is 30uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355994 04/17/11 16:28 yez target screen -> Give more coolant on LD2 and reduce coolant on He4
  • 355993 04/17/11 16:27 a-onl End_of_Run_3631
  • 355992 04/17/11 16:26 a-onl Start_Run_3631,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=30,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling, current is 30uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355991 04/17/11 16:22 a-onl End_of_Run_3630
  • 355990 04/17/11 16:22 miham Spot++ check for run #3630
  • 355989 04/17/11 16:19 S. Sirca Calibration of beam losses needed
  • 355988 04/17/11 16:18 miham Coda screen shot for run #3630 (45uA)
  • 355987 04/17/11 16:08 a-onl Start_Run_3630,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=45,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling, current is 45uA,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355986 04/17/11 16:04 John Watson Day Shift Summary: Sunday 4/17/11
  • 355985 04/17/11 16:02 a-onl End_of_Run_3629
  • 355984 04/17/11 16:02 adaq Beam Time Accounting
  • 355983 04/17/11 15:52 a-onl Start_Run_3629,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=60,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355982 04/17/11 15:50 johna Trigger efficiency tests
  • 355981 04/17/11 15:48 a-onl End_of_Run_3628
  • 355980 04/17/11 15:37 a-onl Start_Run_3628,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=80,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355979 04/17/11 15:34 a-onl End_of_Run_3627
  • 355978 04/17/11 15:24 a-onl Start_Run_3627,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355977 04/17/11 15:22 a-onl End_of_Run_3626
  • 355976 04/17/11 15:12 a-onl Start_Run_3626,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355975 04/17/11 15:11 a-onl End_of_Run_3625
  • 355974 04/17/11 15:00 a-onl Start_Run_3625,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355973 04/17/11 14:58 a-onl End_of_Run_3624
  • 355972 04/17/11 14:47 a-onl Start_Run_3624,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=100,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355971 04/17/11 14:47 a-onl End_of_Run_3623
  • 355970 04/17/11 14:43 a-onl Start_Run_3623,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=100,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355969 04/17/11 14:30 yez target move to c12 & screen snapshot
  • 355968 04/17/11 14:22 a-onl End_of_Run_3622
  • 355967 04/17/11 14:06 a-onl Start_Run_3622,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=15,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355966 04/17/11 14:02 a-onl End_of_Run_3621
  • 355965 04/17/11 13:51 a-onl Start_Run_3621,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=30,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355964 04/17/11 13:46 a-onl End_of_Run_3620
  • 355963 04/17/11 13:46 a-onl Start_Run_3620,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=30,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355962 04/17/11 13:44 a-onl End_of_Run_3619
  • 355961 04/17/11 13:36 yez New Efficiency triggers comparing with old setting
  • 355960 04/17/11 13:35 a-onl Start_Run_3619,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=45,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355959 04/17/11 13:33 a-onl End_of_Run_3618
  • 355958 04/17/11 13:22 a-onl Start_Run_3618,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=60,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355957 04/17/11 13:19 a-onl End_of_Run_3617
  • 355956 04/17/11 13:04 a-onl End_of_Run_3616
  • 355955 04/17/11 12:53 a-onl Start_Run_3616,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355954 04/17/11 12:53 A. Camsonne Electron trigger efficiency
  • 355953 04/17/11 12:51 auto_snap Radiation and Ozone snapshot
  • 355952 04/17/11 12:51 auto_snap Miscellaneous snapshot
  • 355951 04/17/11 12:51 auto_snap Hall A Tools snapshot
  • 355950 04/17/11 12:51 auto_snap Compton Polarimeter snapshot
  • 355949 04/17/11 12:51 auto_snap BigBox Magnet snapshot
  • 355948 04/17/11 12:51 auto_snap Beam Position and Current snapshot
  • 355947 04/17/11 12:50 a-onl End_of_Run_3615
  • 355946 04/17/11 12:50 a-onl Start_Run_3615,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355945 04/17/11 12:49 a-onl End_of_Run_3614
  • 355944 04/17/11 12:38 a-onl Start_Run_3614,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355943 04/17/11 12:38 a-onl End_of_Run_3613
  • 355942 04/17/11 12:27 yez Download trigger of both arm with expert mode
  • 355941 04/17/11 12:26 a-onl Start_Run_3613,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355940 04/17/11 12:24 atrig Trigger-Download
  • 355939 04/17/11 12:21 atrig Trigger-Download
  • 355938 04/17/11 12:21 atrig Trigger-Download
  • 355937 04/17/11 12:20 A. Camsonne Asked to redownload the trigger
  • 355936 04/17/11 12:19 atrig Trigger-Download
  • 355935 04/17/11 12:18 a-onl End_of_Run_3612
  • 355934 04/17/11 12:17 atrig Trigger-Download
  • 355933 04/17/11 12:15 yez electron/pion ratio becomes much better with new Triggers and after Alex fixed the Cer Linear_Fan
  • 355932 04/17/11 12:07 a-onl Start_Run_3612,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355931 04/17/11 12:05 a-onl End_of_Run_3611
  • 355930 04/17/11 12:02 a-onl Start_Run_3611,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355929 04/17/11 12:01 a-onl End_of_Run_3610
  • 355928 04/17/11 11:53 yez deadtime monitor and scalers
  • 355927 04/17/11 11:50 a-onl Start_Run_3610,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355926 04/17/11 11:45 yez T4 scaler is not giving value
  • 355925 04/17/11 11:42 a-onl End_of_Run_3609
  • 355924 04/17/11 11:38 a-onl Start_Run_3609,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=100,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355923 04/17/11 11:37 a-onl End_of_Run_3608
  • 355922 04/17/11 11:37 Lindgren Deadtime monitor 100 ua
  • 355921 04/17/11 11:36 yez scaler for 100uA
  • 355920 04/17/11 11:35 Lindgren dead time monitor for 110 ua
  • 355919 04/17/11 11:34 a-onl Start_Run_3608,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=100,leftarm_p=2.505,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.505,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Rate Test target boiling ,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355918 04/17/11 11:31 a-onl End_of_Run_3607
  • 355917 04/17/11 11:29 yez scalers for 100 uA
  • 355916 04/17/11 11:23 a-onl Start_Run_3607,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=110,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=cosmics during recovery after the storm. Right arm not ready yet.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355915 04/17/11 11:15 yez target screen - make He4 ready for beam
  • 355914 04/17/11 11:07 yez target move to He4 LOOP1
  • 355913 04/17/11 11:02 Lindgren Spot_L
  • 355912 04/17/11 11:00 a-onl End_of_Run_3606
  • 355911 04/17/11 10:53 a-onl Start_Run_3606,Run_type=Spot++,target_type=BeO,beam_e0=3.353,beam_raster=on,beam_current=5,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=cosmics during recovery after the storm. Right arm not ready yet.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355910 04/17/11 10:51 a-onl End_of_Run_3605
  • 355909 04/17/11 10:45 a-onl Start_Run_3605,Run_type=Spot++,target_type=BeO,beam_e0=3.353,beam_raster=on,beam_current=5,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=cosmics during recovery after the storm. Right arm not ready yet.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355908 04/17/11 10:44 a-onl End_of_Run_3604
  • 355907 04/17/11 10:41 a-onl Start_Run_3604,Run_type=Spot++,target_type=BeO,beam_e0=3.353,beam_raster=on,beam_current=5,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=cosmics during recovery after the storm. Right arm not ready yet.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355906 04/17/11 10:40 a-onl End_of_Run_3603
  • 355905 04/17/11 10:33 a-onl Start_Run_3603,Run_type=Spot++,target_type=BeO,beam_e0=3.353,beam_raster=on,beam_current=5,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=cosmics during recovery after the storm. Right arm not ready yet.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355904 04/17/11 10:31 a-onl End_of_Run_3602
  • 355903 04/17/11 10:12 Folts vacuum problem
  • 355901 04/17/11 09:56 Folts call in
  • 355900 04/17/11 09:01 John Watson Hall A Status
  • 355899 04/17/11 08:53 yez Target Screen
  • 355898 04/17/11 08:44 yez Left Dipole comes back
  • 355897 04/17/11 08:37 yez LHRS Dipole trip
  • 355896 04/17/11 08:37 John Watson Status: 08:30 Sunday 4/17/11
  • 355895 04/17/11 08:21 Fatiha Tech: Next tech on call
  • 355894 04/17/11 08:20 a-onl Start_Run_3602,Run_type=Cosmics,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=0,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=cosmics during recovery after the storm. Right arm not ready yet.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355893 04/17/11 08:18 a-onl End_of_Run_3601
  • 355892 04/17/11 08:13 Lindgren Cosmics
  • 355891 04/17/11 08:12 Lindgren Hall A in controlled Access
  • 355890 04/17/11 08:06 Fatiha Shift Sumary
  • 355889 04/17/11 08:06 adaq Beam Time Accounting
  • 355888 04/17/11 08:06 S. Sirca shift summary and BTA
  • 355887 04/17/11 08:00 sirca Swing shift SUMMARY
  • 355886 04/17/11 07:27 adaq Beam Time Accounting
  • 355885 04/17/11 07:20 Fatiha Snapshot: Beam line entrance to the hall
  • 355884 04/17/11 07:19 wdconinc Controlled access for vacuum group
  • 355883 04/17/11 07:02 Fatiha Paged Tech on call: Todd Ewing
  • 355882 04/17/11 06:44
  • 355881 04/17/11 05:57 wdconinc Vacuum expert in BSY, VTC1C16 is the problematic region
  • 355880 04/17/11 05:17 R. Michaels please make shift summaries
  • 355879 04/17/11 05:15 wdconinc Still no beam due to Hall A beamline vacuum issues...
  • 355878 04/17/11 04:50 auto_snap Radiation and Ozone snapshot
  • 355877 04/17/11 04:50 auto_snap Miscellaneous snapshot
  • 355876 04/17/11 04:50 auto_snap Hall A Tools snapshot
  • 355875 04/17/11 04:50 auto_snap Compton Polarimeter snapshot
  • 355874 04/17/11 04:50 auto_snap BigBox Magnet snapshot
  • 355873 04/17/11 04:50 auto_snap Beam Position and Current snapshot
  • 355872 04/17/11 04:22 wdconinc Target screen shot
  • 355871 04/17/11 03:52 Fatiha We can not get beam right now
  • 355870 04/17/11 03:46 wdconinc Scattering chamber vacuum spike
  • 355869 04/17/11 03:43 wdconinc Moved target to BeO
  • 355868 04/17/11 03:32 Fatiha Hall A tool, magnets set for boiling study kinematics
  • 355867 04/17/11 02:56 A. Camsonne Cosmics and trigger latch look ok
  • 355866 04/17/11 02:30 A. Camsonne Added *rol->dabufp++ = 0xfabc0003; to bb_sfi3.crl
  • 355865 04/17/11 02:18 a-onl Start_Run_3601,Run_type=Cosmics,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=0,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=cosmics during recovery after the storm. Right arm not ready yet.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355864 04/17/11 02:16 a-onl End_of_Run_3600
  • 355863 04/17/11 01:45 a-onl Start_Run_3600,Run_type=Cosmics,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=0,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=cosmics during recovery after the storm. Right arm not ready yet.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355862 04/17/11 01:44 a-onl End_of_Run_3599
  • 355861 04/17/11 01:43 A. Camsonne Using the crl code reading only trigger latch
  • 355860 04/17/11 01:26 a-onl Start_Run_3599,Run_type=Cosmics,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=0,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=cosmics during recovery after the storm. Right arm not ready yet.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355859 04/17/11 01:20 a-onl End_of_Run_3598
  • 355858 04/17/11 01:19 Jack hatsv3 and Left Arm rack tripped off
  • 355857 04/17/11 01:14 A. Camsonne All detectors look ok now
  • 355856 04/17/11 01:14 A. Camsonne All detectors look ok now
  • 355855 04/17/11 01:11 a-onl Start_Run_3598,Run_type=Cosmics,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=0,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=cosmics during recovery after the storm. Right arm not ready yet. ,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355854 04/17/11 01:07 a-onl End_of_Run_3597
  • 355853 04/17/11 00:58 A. Camsonne Whole rack for left arm logic was tripped
  • 355852 04/17/11 00:57 a-onl Start_Run_3597,Run_type=Cosmics,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=0,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=cosmics during recovery after the storm. Right arm not ready yet. After rebooting ROC1 and BBTS1,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355851 04/17/11 00:56 a-onl End_of_Run_3596
  • 355850 04/17/11 00:47 a-onl Start_Run_3596,Run_type=Cosmics,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=0,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=cosmics during recovery after the storm. Right arm not ready yet. After rebooting ROC1 and BBTS1,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • Saturday

  • 355849 04/16/11 23:59 miham Missing signals from HRS-L arm!
  • 355848 04/16/11 23:44 miham Run #3594 missing end-of-run script
  • 355847 04/16/11 23:40 a-onl Start_Run_3595,Run_type=Cosmics,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=0,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=cosmics during recovery after the storm. Right arm not ready yet.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355846 04/16/11 23:08 jpchen Hall A cryotarget recovering completed
  • 355845 04/16/11 23:00 S. Sirca Regulator for HRSR Q1 should be ON
  • 355844 04/16/11 22:37 jpchen CHL trip caused ESR problem and Hall A Helium target vented, recovering now
  • 355843 04/16/11 22:35 a-onl Start_Run_3594,Run_type=Cosmics,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=0,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=cosmics during recovery after the storm. Right arm not ready yet.,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355842 04/16/11 21:59 S. Sirca TODO: reboot hatsv3
  • 355841 04/16/11 21:57 S. Sirca HRS magnets back up except Right Dipole
  • 355840 04/16/11 21:23 Ewing techs
  • 355839 04/16/11 20:59 A. Camsonne Wrong polarity for middle group of linear fan
  • 355838 04/16/11 20:50 auto_snap Radiation and Ozone snapshot
  • 355837 04/16/11 20:50 auto_snap Miscellaneous snapshot
  • 355836 04/16/11 20:50 auto_snap Hall A Tools snapshot
  • 355835 04/16/11 20:50 auto_snap Compton Polarimeter snapshot
  • 355834 04/16/11 20:50 auto_snap BigBox Magnet snapshot
  • 355833 04/16/11 20:50 auto_snap Beam Position and Current snapshot
  • 355832 04/16/11 20:40 ewing techs
  • 355831 04/16/11 20:33 miham Run #3593 is JUNK!
  • 355830 04/16/11 19:57 A. Camsonne Vacuum going bad in the beam line
  • 355829 04/16/11 19:53 A. Camsonne NIM crate for Left arm gas cerenkov voltages reading 0
  • 355828 04/16/11 19:13 A. Shahinyan Target
  • 355827 04/16/11 19:04 A. Camsonne Left HRS cerenkov trigger inefficiency
  • 355826 04/16/11 19:02 S. Sirca HRS-L and -R magnets ramped down
  • 355825 04/16/11 18:53 a-onl End_of_Run_3593
  • 355824 04/16/11 18:45 miham Coda Screen shot for run #3593. HRS-L run ONLY!
  • 355823 04/16/11 18:39 a-onl Start_Run_3593,Run_type=Production SRC,target_type=LD 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.795,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.795,rightarm_theta=25.0,rightarm_sieve=out,comment_text=production for HRSL, HRSR dipole not ready, still ramping down,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355822 04/16/11 18:39 A. Shahinyan Target
  • 355821 04/16/11 18:04 a-onl End_of_Run_3592
  • 355820 04/16/11 17:58 V. Sulkosky /adaql10/work1 Continued clean up
  • 355819 04/16/11 17:33 miham Coda Screen shot for run #3592
  • 355818 04/16/11 17:32 miham Spot++ check for run #3592
  • 355817 04/16/11 17:29 a-onl Start_Run_3592,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355816 04/16/11 17:29 a-onl End_of_Run_3591
  • 355815 04/16/11 17:14 johna Left HRS angle camera snapshots
  • 355814 04/16/11 17:11 johna Right HRS angle camera snapshots
  • 355813 04/16/11 17:01 miham Spot++ check for run #3591
  • 355812 04/16/11 16:59 miham Coda GUI screen shot for run #3591
  • 355811 04/16/11 16:58 A. Shahinyan Target
  • 355810 04/16/11 16:57 a-onl Start_Run_3591,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355809 04/16/11 16:55 a-onl End_of_Run_3590
  • 355808 04/16/11 16:18 miham Spot++ check for run #3590
  • 355807 04/16/11 16:17 miham Spot++ check for run #3590
  • 355806 04/16/11 16:12 A. Shahinyan Target
  • 355805 04/16/11 16:10 John Watson Shift Summary Saturday Day (4/16/11)
  • 355804 04/16/11 16:04 adaq Beam Time Accounting
  • 355803 04/16/11 16:00 A. Camsonne Overlap of gas cerenkov with T1 and T3 looks good
  • 355802 04/16/11 15:59 John Watson Spot++ Run 3589 (thick C, 95 uA)
  • 355801 04/16/11 15:56 yez dead time screen
  • 355800 04/16/11 15:56 A. Camsonne Overlap of gas cerenkov with T1 and T3 looks good
  • 355799 04/16/11 15:53 a-onl Start_Run_3590,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355798 04/16/11 15:51 a-onl End_of_Run_3589
  • 355797 04/16/11 15:47 a-onl Start_Run_3589,Run_type=Spot++,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=1,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Test DAQ timing check,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355796 04/16/11 15:46 a-onl End_of_Run_3588
  • 355795 04/16/11 15:45 a-onl Start_Run_3588,Run_type=Other,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=1,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Test DAQ timing check,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355794 04/16/11 15:22 yez target screen
  • 355793 04/16/11 14:56 A. Camsonne Scope traces for T1,T3 and left and right cerenkov
  • 355792 04/16/11 14:54 a-onl End_of_Run_3587
  • 355791 04/16/11 14:49 a-onl Start_Run_3587,Run_type=Other,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=60,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Test DAQ timing check,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355790 04/16/11 14:41 V. Sulkosky New left and right arm Triggers
  • 355789 04/16/11 14:33 V. Sulkosky L-HRS Patch Panel Cable changes
  • 355788 04/16/11 14:29 V. Sulkosky R-HRS Patch Panel Cable changes
  • 355787 04/16/11 14:12 a-onl Start_Run_3586,Run_type=Other,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=60,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Test DAQ timing check,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355786 04/16/11 12:50 auto_snap Radiation and Ozone snapshot
  • 355785 04/16/11 12:50 auto_snap Miscellaneous snapshot
  • 355784 04/16/11 12:50 auto_snap Hall A Tools snapshot
  • 355783 04/16/11 12:50 auto_snap Compton Polarimeter snapshot
  • 355782 04/16/11 12:50 auto_snap BigBox Magnet snapshot
  • 355781 04/16/11 12:50 auto_snap Beam Position and Current snapshot
  • 355780 04/16/11 12:48 a-onl End_of_Run_3585
  • 355779 04/16/11 12:21 yez MCC is still calibrating the ion chamber but they told me that we can continue to take data
  • 355778 04/16/11 12:16 a-onl Start_Run_3585,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=60,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355777 04/16/11 12:15 a-onl End_of_Run_3584
  • 355776 04/16/11 11:53 yez A late halog: MCC needs to calibrate Ion chamber and it takes one hour(started from 25min ago)
  • 355775 04/16/11 11:26 a-onl Start_Run_3584,Run_type=Calibration,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=60,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Ion Chamber Calibration.MCC still tuning beam,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355774 04/16/11 11:11 a-onl End_of_Run_3583
  • 355773 04/16/11 10:54 Lindgren Run 3582 sync scaler diff
  • 355772 04/16/11 10:52 Lindgren Run 3582 Sync plots
  • 355771 04/16/11 10:47 a-onl Start_Run_3583,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=60,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production.MCC still tuning beam,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355770 04/16/11 10:23 John Watson Hall A tools and current monitor--Sat. a.m.Hall A tools and current monitor--Sat. a.m.
  • 355769 04/16/11 10:17 yez target move to LH4 20cm
  • 355768 04/16/11 10:16 a-onl End_of_Run_3582
  • 355767 04/16/11 10:10 a-onl Start_Run_3582,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=45,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355766 04/16/11 10:07 a-onl End_of_Run_3581
  • 355765 04/16/11 09:54 yez target screen
  • 355764 04/16/11 09:50 yez During run 3581, only the first few minutes beam is available. Most of data are cosmic
  • 355763 04/16/11 09:22 yez Start to take Carbon data. MCC currently only can send 5uA beam to us
  • 355762 04/16/11 09:21 a-onl Start_Run_3581,Run_type=Production SRC,target_type=Thick Carbon,beam_e0=3.353,beam_raster=on,beam_current=5,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355761 04/16/11 09:19 yez target move to Carbon 5mm
  • 355760 04/16/11 09:19 a-onl End_of_Run_3580
  • 355759 04/16/11 09:05 yez Beam off. MCC has some issues to deliver beam to hall A
  • 355758 04/16/11 09:04 yez Beam had problem during most of this run time.
  • 355757 04/16/11 09:01 a-onl Start_Run_3580,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355756 04/16/11 08:58 a-onl End_of_Run_3579
  • 355755 04/16/11 08:34 a-onl Start_Run_3579,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355754 04/16/11 08:32 a-onl End_of_Run_3578
  • 355753 04/16/11 08:07 adaq Beam Time Accounting
  • 355752 04/16/11 08:05 Fatiha Shift Summary
  • 355751 04/16/11 07:26 xqiu spot++ check for run #3578
  • 355750 04/16/11 07:25 xqiu sync plots #2 for run #3578
  • 355749 04/16/11 07:24 xqiu sync plots #1 fro run #3578
  • 355748 04/16/11 07:18 a-onl Start_Run_3578,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355747 04/16/11 07:17 a-onl End_of_Run_3577
  • 355746 04/16/11 07:11 xqiu Physics check for the run #3576
  • 355745 04/16/11 06:24 xqiu spot++ check for run #3577
  • 355744 04/16/11 06:21 xqiu sync plots #2 for run #3577
  • 355743 04/16/11 06:20 xqiu sync plots #1 for run #3577
  • 355742 04/16/11 06:11 a-onl Start_Run_3577,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355741 04/16/11 06:10 a-onl End_of_Run_3576
  • 355740 04/16/11 06:05 xqiu physics check for run #3575
  • 355739 04/16/11 05:21 xqiu spot++ check for run #3576
  • 355738 04/16/11 05:20 xqiu sync plots #2 for run #3576
  • 355737 04/16/11 05:19 xqiu sync plots #1 for run #3576
  • 355736 04/16/11 05:14 xqiu Physics check for run #3574
  • 355735 04/16/11 05:07 xqiu ROC10 crashed at the end of the run #3575.
  • 355734 04/16/11 05:06 a-onl Start_Run_3576,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355733 04/16/11 04:50 auto_snap Radiation and Ozone snapshot
  • 355732 04/16/11 04:49 auto_snap Miscellaneous snapshot
  • 355731 04/16/11 04:49 auto_snap Hall A Tools snapshot
  • 355730 04/16/11 04:49 auto_snap Compton Polarimeter snapshot
  • 355729 04/16/11 04:49 auto_snap BigBox Magnet snapshot
  • 355728 04/16/11 04:49 auto_snap Beam Position and Current snapshot
  • 355727 04/16/11 04:29 Fatiha Beam current for the last 4h
  • 355726 04/16/11 04:28 Fatiha Hall A Tools
  • 355725 04/16/11 04:22 xqiu spot++ check for run #3575
  • 355724 04/16/11 04:21 xqiu sync plots #2 for run #3575
  • 355723 04/16/11 04:20 xqiu sync plots #1 for the run #3575
  • 355722 04/16/11 04:10 xqiu Physics check for run #3573
  • 355721 04/16/11 04:09 a-onl Start_Run_3575,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355720 04/16/11 04:08 a-onl End_of_Run_3574
  • 355719 04/16/11 04:02 wdconinc Target screen shots
  • 355718 04/16/11 03:49 xqiu Deadtime monitor and DAQ heads up display shot
  • 355717 04/16/11 03:47 xqiu Hall A scaler data
  • 355716 04/16/11 03:17 xqiu spot++ check for run #3574
  • 355715 04/16/11 03:16 xqiu sync plots #2 for run #3574
  • 355714 04/16/11 03:15 xqiu sync plots #1 for run #3574
  • 355713 04/16/11 03:07 a-onl Start_Run_3574,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355712 04/16/11 03:06 a-onl End_of_Run_3573
  • 355711 04/16/11 03:02 xqiu physics check for run #3572
  • 355710 04/16/11 02:51 xqiu spot++ check for run #3573
  • 355709 04/16/11 02:48 xqiu sync plot #2 in the run #3573
  • 355708 04/16/11 02:47 xqiu sync plots in the run #3573
  • 355707 04/16/11 02:41 xqiu The first 40000 events in the run #3573, the L.tr.vz is there.
  • 355706 04/16/11 02:23 xqiu The L.tr.vz in the run #3570
  • 355705 04/16/11 02:19 xqiu Missing L.tr.vz PLOT
  • 355704 04/16/11 02:05 a-onl Start_Run_3573,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355703 04/16/11 01:58 a-onl End_of_Run_3572
  • 355702 04/16/11 01:46 xqiu sync plots #2 for run #3572
  • 355701 04/16/11 01:45 xqiu sync plot #1 for run #3572
  • 355700 04/16/11 01:27 xqiu spot++ check for run #3572
  • 355699 04/16/11 01:12 xqiu sync plots for run #3571
  • 355698 04/16/11 01:10 xqiu sync check for run 3571
  • 355697 04/16/11 00:58 a-onl Start_Run_3572,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355696 04/16/11 00:56 a-onl End_of_Run_3571
  • 355695 04/16/11 00:02 adaq Beam Time Accounting
  • 355694 04/16/11 00:00 miham Spot++ check for run #3571
  • Friday

  • 355693 04/15/11 23:57 miham Coda GUI screen shot for run #3571
  • 355692 04/15/11 23:57 S. Sirca Swing shift SUMMARY
  • 355691 04/15/11 23:55 a-onl Start_Run_3571,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355690 04/15/11 23:54 a-onl End_of_Run_3570
  • 355689 04/15/11 23:08 miham Spot++ check for run #3570
  • 355688 04/15/11 23:06 miham Sync plots #2 for run #3570
  • 355687 04/15/11 23:06 miham Sync plots #1 for run #3570
  • 355686 04/15/11 23:01 miham Coda GUI screen shot for run #3570
  • 355685 04/15/11 22:58 A. Camsonne Trying buffered configuration B2ArmNTS11
  • 355684 04/15/11 22:55 a-onl Start_Run_3570,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355683 04/15/11 22:48 A. Camsonne Trying rebooting ROCs
  • 355682 04/15/11 22:39 miham CODA problem with communication
  • 355681 04/15/11 22:36 miham End-of-run file is missing for run #3569
  • 355680 04/15/11 21:37 miham Spot++ check for run #3569
  • 355679 04/15/11 21:35 miham Coda GUI screen shot for run #3569
  • 355678 04/15/11 21:34 a-onl Start_Run_3569,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355677 04/15/11 21:32 a-onl End_of_Run_3568
  • 355676 04/15/11 20:49 auto_snap Radiation and Ozone snapshot
  • 355675 04/15/11 20:49 auto_snap Miscellaneous snapshot
  • 355674 04/15/11 20:49 auto_snap Hall A Tools snapshot
  • 355673 04/15/11 20:49 auto_snap Compton Polarimeter snapshot
  • 355672 04/15/11 20:49 auto_snap BigBox Magnet snapshot
  • 355671 04/15/11 20:49 auto_snap Beam Position and Current snapshot
  • 355670 04/15/11 20:31 miham Screen shot of Coda GUI for run #3568. Run Numbers do not update!!!
  • 355669 04/15/11 20:27 miham Spot++ check for run #3568
  • 355668 04/15/11 20:22 a-onl Start_Run_3568,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=on,beam_current=95,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355667 04/15/11 20:20 a-onl End_of_Run_3567
  • 355666 04/15/11 20:11 V. Sulkosky /adaql10/work1 partial clean up
  • 355665 04/15/11 19:23 miham Spot++ check for run #3567
  • 355664 04/15/11 19:21 miham Coda GUI screen shot for run #3567
  • 355663 04/15/11 19:20 a-onl Start_Run_3567,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=off,beam_current=95,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355662 04/15/11 19:18 a-onl End_of_Run_3566
  • 355661 04/15/11 19:17 S. Sirca Comment on BB
  • 355660 04/15/11 18:23 miham Spot++ check for run #3566
  • 355659 04/15/11 18:21 miham Screen shot of the Coda GUI for run #3566
  • 355658 04/15/11 18:19 a-onl Start_Run_3566,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=off,beam_current=87,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355657 04/15/11 18:16 a-onl End_of_Run_3565
  • 355656 04/15/11 17:25 miham Screen shot of HRS scalers for run #3565
  • 355655 04/15/11 17:24 miham Coda GUI screen shot for run #3565
  • 355654 04/15/11 17:23 miham Spot++ check for run #3565
  • 355653 04/15/11 17:19 yez L-HRS: Put HV of PRL back to what it was, before the production data start taking
  • 355652 04/15/11 17:17 a-onl Start_Run_3565,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=off,beam_current=80,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Hopefully production,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355651 04/15/11 17:06 miham Run #3563 is a 4He run, not BeO
  • 355650 04/15/11 17:03 a-onl End_of_Run_3564
  • 355649 04/15/11 17:02 S. Sirca spot check with run 3563
  • 355648 04/15/11 17:02 a-onl Start_Run_3564,Run_type=Spot++,target_type=4He 20cm cell,beam_e0=3.353,beam_raster=off,beam_current=5,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Spot++ during beam recovery procedure,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355647 04/15/11 17:00 a-onl End_of_Run_3563
  • 355646 04/15/11 16:58 a-onl Start_Run_3563,Run_type=Spot++,target_type=BeO,beam_e0=3.353,beam_raster=off,beam_current=5,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Spot++ during beam recovery procedure,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355645 04/15/11 16:56 C. Gu Target move to 20cm He4
  • 355644 04/15/11 16:50 S. Sirca spot check with run 3562
  • 355643 04/15/11 16:49 a-onl End_of_Run_3562
  • 355642 04/15/11 16:44 S. Sirca spot check with run 3561
  • 355641 04/15/11 16:44 a-onl Start_Run_3562,Run_type=Spot++,target_type=BeO,beam_e0=3.353,beam_raster=off,beam_current=5,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Spot++ during beam recovery procedure,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355640 04/15/11 16:42 a-onl End_of_Run_3561
  • 355639 04/15/11 16:33 a-onl Start_Run_3561,Run_type=Spot++,target_type=BeO,beam_e0=3.353,beam_raster=off,beam_current=5,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2.995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=Spot++ during beam recovery procedure,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355638 04/15/11 16:26 C. Gu Target move to BeO
  • 355637 04/15/11 16:26 a-onl End_of_Run_3560
  • 355636 04/15/11 16:19 C. Gu Target Screenshot
  • 355635 04/15/11 16:15 adaq Beam Time Accounting
  • 355634 04/15/11 16:10 S. Sirca HRS-R Q1 ramping up
  • 355633 04/15/11 16:08 fansler RQ1
  • 355632 04/15/11 15:27 a-onl Start_Run_3560,Run_type=Spot++,target_type=BeO,beam_e0=3.355,beam_raster=off,beam_current=5,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=checking detectors while doing harp scan,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355631 04/15/11 15:24 a-onl End_of_Run_3559
  • 355630 04/15/11 15:18 Higinbotham Feedback Problem
  • 355629 04/15/11 15:07 John Watson RHRS Q1 trips
  • 355628 04/15/11 14:51 a-onl Start_Run_3559,Run_type=Other,target_type=BeO,beam_e0=3.355,beam_raster=off,beam_current=5,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=checking detectors while doing harp scan,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355627 04/15/11 14:51 yez target screen
  • 355626 04/15/11 14:49 a-onl End_of_Run_3558
  • 355625 04/15/11 14:49 A. Camsonne Right Arm NMR compensating coil tweaked
  • 355624 04/15/11 14:41 a-onl Start_Run_3558,Run_type=Other,target_type=BeO,beam_e0=3.355,beam_raster=off,beam_current=5,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=checking detectors while doing harp scan,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355623 04/15/11 14:39 a-onl End_of_Run_3557
  • 355622 04/15/11 14:35 a-onl Start_Run_3557,Run_type=Other,target_type=BeO,beam_e0=3.355,beam_raster=off,beam_current=5,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=checking detectors while doing harp scan,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355621 04/15/11 14:33 a-onl End_of_Run_3556
  • 355620 04/15/11 14:17 a-onl Start_Run_3556,Run_type=Other,target_type=BeO,beam_e0=3.355,beam_raster=off,beam_current=5,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=checking detectors while doing harp scan,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355619 04/15/11 14:16 a-onl End_of_Run_3555
  • 355618 04/15/11 14:11 a-onl Start_Run_3555,Run_type=Other,target_type=BeO,beam_e0=3.355,beam_raster=off,beam_current=5,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=checking detectors while doing harp scan,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355617 04/15/11 14:10 a-onl End_of_Run_3554
  • 355616 04/15/11 13:56 a-onl Start_Run_3554,Run_type=Other,target_type=BeO,beam_e0=3.355,beam_raster=off,beam_current=5,leftarm_p=2.995,leftarm_theta=25.0,leftarm_sieve=out,rightarm_p=2995,rightarm_theta=25.0,rightarm_sieve=out,comment_text=checking detectors while doing harp scan,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355615 04/15/11 13:30 a-onl End_of_Run_3553
  • 355614 04/15/11 13:18 a-onl Start_Run_3553,Run_type=Spot++,target_type=Optics: multi 1mm C foils,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=Switch to X>2 experiment,SPOT++,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355613 04/15/11 13:09 yez target move to BeO
  • 355612 04/15/11 12:49 auto_snap Radiation and Ozone snapshot
  • 355611 04/15/11 12:49 auto_snap Miscellaneous snapshot
  • 355610 04/15/11 12:49 auto_snap Hall A Tools snapshot
  • 355609 04/15/11 12:49 auto_snap Compton Polarimeter snapshot
  • 355608 04/15/11 12:49 auto_snap BigBox Magnet snapshot
  • 355607 04/15/11 12:49 auto_snap Beam Position and Current snapshot
  • 355606 04/15/11 12:22 yez HVs of HAND and Bigbit are off
  • 355605 04/15/11 11:56 Jack (Techs) Right Dipole DAC replaced
  • 355604 04/15/11 11:52 hsmith Right dipole polarity change
  • 355603 04/15/11 11:12 Hanretty Left and Right HRS moved
  • 355602 04/15/11 10:16 a-onl Start_Run_3552,Run_type=Cosmics,target_type=Optics: multi 1mm C foils,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=Switch to X>2 experiment,Cosmic Ray run during pass change and R-HRS polarity change,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355601 04/15/11 10:14 yez R-HRS: HV of Shower 4
  • 355600 04/15/11 10:13 yez R-HRS: HV of Shower 3
  • 355599 04/15/11 10:12 yez R-HRS: HV of Shower 2
  • 355598 04/15/11 10:11 yez R-HRS: HV of Preshower 3 & Shower 1
  • 355597 04/15/11 10:08 yez R-HRS: HV of Preshower 2
  • 355596 04/15/11 10:07 yez R-HRS: HV of S2m&S0&PS1
  • 355595 04/15/11 10:03 yez R-HRS: HV of S2m
  • 355594 04/15/11 10:00 yez R-HRS: HV of S1&Cer
  • 355593 04/15/11 09:55 yez Increase HV of PRL2.#18,#19
  • 355592 04/15/11 09:41 a-onl End_of_Run_3551
  • 355591 04/15/11 09:34 kelleher Owl Shift+ Summary
  • 355590 04/15/11 09:31 a-onl Start_Run_3551,Run_type=Optics,target_type=Optics: multi 1mm C foils,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=4 uA BB optics, magnet on, raster off,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355589 04/15/11 09:25 kelleher BigBite power ON
  • 355588 04/15/11 09:22 a-onl End_of_Run_3550
  • 355587 04/15/11 09:13 a-onl Start_Run_3550,Run_type=Optics,target_type=Optics: multi 1mm C foils,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=4 uA BB optics, magnet off, raster off,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355586 04/15/11 09:11 a-onl End_of_Run_3549
  • 355585 04/15/11 09:11 a-onl Start_Run_3549,Run_type=Optics,target_type=Optics: multi 1mm C foils,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=4 uA BB optics, magnet off, raster off,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355584 04/15/11 09:05 a-onl End_of_Run_3548
  • 355583 04/15/11 08:57 kelleher spot ++
  • 355582 04/15/11 08:55 a-onl Start_Run_3548,Run_type=Optics,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=4 uA BB optics, magnet off, raster on,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355581 04/15/11 08:51 a-onl End_of_Run_3547
  • 355580 04/15/11 08:32 a-onl Start_Run_3547,Run_type=Optics,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=4 uA BB optics, magnet off, raster off,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355579 04/15/11 08:31 a-onl End_of_Run_3546
  • 355578 04/15/11 08:22 a-onl Start_Run_3546,Run_type=Optics,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=4 uA BB optics, magnet on, raster off,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355577 04/15/11 08:19 a-onl End_of_Run_3545
  • 355576 04/15/11 08:09 adaq Beam Time Accounting
  • 355575 04/15/11 08:06 a-onl Start_Run_3545,Run_type=Optics,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=4 uA BB optics, magnet on, raster on,bbite_current=518,bbite_theta=100,hand_angle=92,bbite_sieve=sieve,hand_lead=on
  • 355574 04/15/11 08:04 a-onl End_of_Run_3544
  • 355573 04/15/11 08:02 Leckey Target moved to BeO
  • 355572 04/15/11 07:58 a-onl Start_Run_3544,Run_type=Check out,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=DAQ test,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355571 04/15/11 07:55 a-onl End_of_Run_3543
  • 355570 04/15/11 07:44 orchen HRS Coincedence time for 850MeV/c kinematics
  • 355569 04/15/11 07:26 a-onl Start_Run_3543,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=40,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=40 uA production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355568 04/15/11 07:26 a-onl End_of_Run_3542
  • 355567 04/15/11 07:20 a-onl Start_Run_3542,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=40,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=40 uA production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355566 04/15/11 07:19 a-onl End_of_Run_3541
  • 355565 04/15/11 07:18 a-onl Start_Run_3541,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=40,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=40 uA production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355564 04/15/11 07:16 a-onl End_of_Run_3540
  • 355563 04/15/11 07:15 Leckey Target Regulation Issue Solved
  • 355562 04/15/11 06:53 Leckey Loop temp dropped again to 18K
  • 355561 04/15/11 06:51 a-onl Start_Run_3540,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=10,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=10 uA production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355560 04/15/11 06:49 a-onl End_of_Run_3539
  • 355559 04/15/11 06:06 Leckey Target: He 1K cooler??
  • 355558 04/15/11 06:04 a-onl Start_Run_3539,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=10,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=10 uA production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355557 04/15/11 06:03 a-onl End_of_Run_3538
  • 355556 04/15/11 05:29 kelleher P_miss for run #3536
  • 355555 04/15/11 05:18 a-onl Start_Run_3538,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=10,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=10 uA production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355554 04/15/11 05:17 a-onl End_of_Run_3537
  • 355553 04/15/11 04:55 kelleher Hall A Tools
  • 355552 04/15/11 04:49 auto_snap Radiation and Ozone snapshot
  • 355551 04/15/11 04:48 auto_snap Miscellaneous snapshot
  • 355550 04/15/11 04:48 auto_snap Hall A Tools snapshot
  • 355549 04/15/11 04:48 auto_snap Compton Polarimeter snapshot
  • 355548 04/15/11 04:48 auto_snap BigBox Magnet snapshot
  • 355547 04/15/11 04:48 auto_snap Beam Position and Current snapshot
  • 355546 04/15/11 04:29 a-onl Start_Run_3537,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=10,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=10 uA production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355545 04/15/11 04:29 a-onl End_of_Run_3536
  • 355544 04/15/11 03:44 wdconinc Rebooted TS0 and TS11
  • 355543 04/15/11 03:43 a-onl Start_Run_3536,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=10,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=10 uA production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355542 04/15/11 03:40 a-onl End_of_Run_3535
  • 355541 04/15/11 03:30 Leckey Target Snapshot
  • 355540 04/15/11 02:55 a-onl Start_Run_3535,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=10,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=10 uA production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355539 04/15/11 02:54 a-onl End_of_Run_3534
  • 355538 04/15/11 02:09 a-onl Start_Run_3534,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=10,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=10 uA production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355537 04/15/11 02:08 a-onl End_of_Run_3533
  • 355536 04/15/11 02:03 kelleher Half Waveplate: OUT
  • 355535 04/15/11 01:23 a-onl Start_Run_3533,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=10,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=10 uA production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355534 04/15/11 01:19 a-onl End_of_Run_3532
  • 355533 04/15/11 01:00 V. Sulkosky RHRS Angle and Momentum Change
  • 355532 04/15/11 00:56 a-onl Start_Run_3532,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=10,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=10 uA production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355531 04/15/11 00:55 a-onl End_of_Run_3531
  • 355530 04/15/11 00:51 a-onl Start_Run_3531,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=2,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=2 uA production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355529 04/15/11 00:47 wdconinc spot++
  • 355528 04/15/11 00:46 a-onl End_of_Run_3530
  • 355527 04/15/11 00:46 kelleher RHRS dipole regulation off
  • 355526 04/15/11 00:44 a-onl Start_Run_3530,Run_type=Spot++,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.099,rightarm_theta=20.0,rightarm_sieve=out,comment_text=1 uA spot check,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355525 04/15/11 00:37 A. Camsonne 20 db attenuator added to HAPPEX upstream cavity
  • 355524 04/15/11 00:02 adaq Beam Time Accounting
  • Thursday

  • 355523 04/14/11 23:56 S. Sirca Swing shift SUMMARY
  • 355522 04/14/11 23:35 miham Analysis plots for run #3529
  • 355521 04/14/11 23:30 a-onl End_of_Run_3529
  • 355520 04/14/11 23:04 miham Analysis plots for run #3528
  • 355519 04/14/11 23:00 miham Screen shot of the HRS scalers for run #3529
  • 355518 04/14/11 22:59 miham Coda GUI screen shot for run #3529
  • 355517 04/14/11 22:55 a-onl Start_Run_3529,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=60,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.449,rightarm_theta=38.5,rightarm_sieve=out,comment_text=60 uA prod.,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355516 04/14/11 22:53 a-onl End_of_Run_3528
  • 355515 04/14/11 22:04 S. Sirca spot_L check for run 3528
  • 355514 04/14/11 22:00 miham Analysis Plots for run #3527
  • 355513 04/14/11 21:59 yez target screen
  • 355512 04/14/11 21:56 miham Screen shot of the HRS scalers for run #3528
  • 355511 04/14/11 21:55 miham Coda GUI screen shot for run #3528
  • 355510 04/14/11 21:53 a-onl Start_Run_3528,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=60,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.449,rightarm_theta=38.5,rightarm_sieve=out,comment_text=60 uA prod.,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355509 04/14/11 21:52 a-onl End_of_Run_3527
  • 355508 04/14/11 21:10 miham Screen shot of the HRS scalers for run #3527
  • 355507 04/14/11 21:09 miham Coda GUI screen shot for run #3527
  • 355506 04/14/11 21:07 miham Analysis plots for run #3526
  • 355505 04/14/11 20:50 a-onl Start_Run_3527,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=60,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.449,rightarm_theta=38.5,rightarm_sieve=out,comment_text=60 uA prod.,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355504 04/14/11 20:49 a-onl End_of_Run_3526
  • 355503 04/14/11 20:48 auto_snap Radiation and Ozone snapshot
  • 355502 04/14/11 20:48 auto_snap Miscellaneous snapshot
  • 355501 04/14/11 20:48 auto_snap Hall A Tools snapshot
  • 355500 04/14/11 20:48 auto_snap Compton Polarimeter snapshot
  • 355499 04/14/11 20:48 auto_snap BigBox Magnet snapshot
  • 355498 04/14/11 20:48 auto_snap Beam Position and Current snapshot
  • 355497 04/14/11 20:05 miham Analysis plots for run #3525
  • 355496 04/14/11 19:52 miham Screen shot of the HRS scalers for run #3526
  • 355495 04/14/11 19:51 miham Coda GUI screen shot for run #3526
  • 355494 04/14/11 19:50 V. Sulkosky CT for ~ 350 MeV/c upto Run 3524
  • 355493 04/14/11 19:49 a-onl Start_Run_3526,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=60,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.449,rightarm_theta=38.5,rightarm_sieve=out,comment_text=60 uA prod.,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355492 04/14/11 19:48 a-onl End_of_Run_3525
  • 355491 04/14/11 19:33 yez target screen
  • 355490 04/14/11 19:12 miham Analysis plots for run #3524
  • 355489 04/14/11 19:09 miham Screen shot of the HRS scalers for run #3525
  • 355488 04/14/11 19:08 miham Coda GUI screen shot for run #3525
  • 355487 04/14/11 18:48 a-onl Start_Run_3525,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=60,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.449,rightarm_theta=38.5,rightarm_sieve=out,comment_text=60 uA prod.,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355486 04/14/11 18:47 a-onl End_of_Run_3524
  • 355485 04/14/11 18:40 adq two coinc time peaks
  • 355484 04/14/11 17:47 miham Screen shot of HRS scalers for run #3524
  • 355483 04/14/11 17:46 miham Coda GUI screen shot for run #3524
  • 355482 04/14/11 17:44 a-onl Start_Run_3524,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=60,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.449,rightarm_theta=38.5,rightarm_sieve=out,comment_text=60 uA prod.,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355481 04/14/11 17:43 miham Analysis plots for run #3523
  • 355480 04/14/11 17:38 a-onl End_of_Run_3523
  • 355479 04/14/11 17:35 yez target: open JT valve twice during run#3523
  • 355478 04/14/11 17:08 miham Analysis plots for run #3522
  • 355477 04/14/11 17:07 miham Coda GUI screen shot for run #3523
  • 355476 04/14/11 17:05 miham Screen shot of HRS scalers for run #3523
  • 355475 04/14/11 16:56 a-onl Start_Run_3523,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=60,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.449,rightarm_theta=38.5,rightarm_sieve=out,comment_text=60 uA prod.,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355474 04/14/11 16:54 a-onl End_of_Run_3522
  • 355473 04/14/11 16:28 yez target screen
  • 355472 04/14/11 16:12 adaq Beam Time Accounting
  • 355471 04/14/11 15:54 a-onl Start_Run_3522,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=60,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.449,rightarm_theta=38.5,rightarm_sieve=out,comment_text=60 uA prod.,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355470 04/14/11 15:53 John Watson p_miss events, Thursday Day Shift
  • 355469 04/14/11 15:24 posik run 3521 p miss plots
  • 355468 04/14/11 15:17 a-onl End_of_Run_3521
  • 355467 04/14/11 15:04 J.Watson Spot++
  • 355466 04/14/11 14:56 posik run 3520 crashed
  • 355465 04/14/11 14:51 posik run 3519 TS12 crash
  • 355464 04/14/11 14:50 a-onl Start_Run_3521,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=60,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.449,rightarm_theta=38.5,rightarm_sieve=out,comment_text=60 uA prod.,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355463 04/14/11 14:41 a-onl Start_Run_3520,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=60,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.449,rightarm_theta=38.5,rightarm_sieve=out,comment_text=60 uA prod.,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355462 04/14/11 14:34 a-onl End_of_Run_3519
  • 355461 04/14/11 14:28 Solvignon Boiling study results on 20cm 4he
  • 355460 04/14/11 13:58 a-onl Start_Run_3519,Run_type=Cosmics,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.449,rightarm_theta=38.5,rightarm_sieve=out,comment_text=cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355459 04/14/11 13:54 posik should be 3516 at 70uA
  • 355458 04/14/11 13:50 posik run 3517 missing p plots at 80 uA
  • 355457 04/14/11 13:46 a-onl End_of_Run_3518
  • 355456 04/14/11 13:44 posik run 383 80uA missing p plots
  • 355455 04/14/11 13:40 a-onl Start_Run_3518,Run_type=Production SRC,target_type=4He 10cm cell,beam_e0=4.46,beam_raster=on,beam_current=90,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.449,rightarm_theta=38.5,rightarm_sieve=out,comment_text=90 uA 10cm cell,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355454 04/14/11 13:38 posik run 3515 missing p
  • 355453 04/14/11 13:33 a-onl End_of_Run_3517
  • 355452 04/14/11 13:21 a-onl Start_Run_3517,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=80,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.449,rightarm_theta=38.5,rightarm_sieve=out,comment_text=80 uA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355451 04/14/11 13:18 a-onl End_of_Run_3516
  • 355450 04/14/11 13:05 a-onl Start_Run_3516,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=60,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.449,rightarm_theta=38.5,rightarm_sieve=out,comment_text=70 uA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355449 04/14/11 13:01 a-onl End_of_Run_3515
  • 355448 04/14/11 12:52 a-onl Start_Run_3515,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=60,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.449,rightarm_theta=38.5,rightarm_sieve=out,comment_text=60 uA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355447 04/14/11 12:51 a-onl End_of_Run_3514
  • 355446 04/14/11 12:48 auto_snap Radiation and Ozone snapshot
  • 355445 04/14/11 12:48 auto_snap Miscellaneous snapshot
  • 355444 04/14/11 12:48 auto_snap Hall A Tools snapshot
  • 355443 04/14/11 12:48 auto_snap Compton Polarimeter snapshot
  • 355442 04/14/11 12:48 auto_snap BigBox Magnet snapshot
  • 355441 04/14/11 12:48 auto_snap Beam Position and Current snapshot
  • 355440 04/14/11 12:35 a-onl Start_Run_3514,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=60,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.449,rightarm_theta=38.5,rightarm_sieve=out,comment_text=60 uA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355439 04/14/11 12:33 Fatiha Data looks very good!
  • 355438 04/14/11 12:32 a-onl End_of_Run_3513
  • 355437 04/14/11 12:21 V. Sulkosky Added Acceptance Cuts to Pmiss
  • 355436 04/14/11 11:59 A. Camsonne Spot++ set to default to spot_L
  • 355435 04/14/11 10:47 posik run 3506 missing p plots
  • 355434 04/14/11 10:45 posik run 3505 missing p plots
  • 355433 04/14/11 10:38 A. Camsonne BCM HAPPEX cavity test
  • 355432 04/14/11 10:34 a-onl Start_Run_3513,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=50,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=test UB2ArmNTS11,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355431 04/14/11 10:33 a-onl End_of_Run_3512
  • 355430 04/14/11 10:32 A. Camsonne Right arm BPMA still not working
  • 355429 04/14/11 10:27 A. Camsonne Configuration for Two arms readout only UB2ArmNTS11
  • 355428 04/14/11 10:25 a-onl Start_Run_3512,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=50,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=test UB2ArmNTS11,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355427 04/14/11 10:23 a-onl End_of_Run_3511
  • 355426 04/14/11 10:22 a-onl Start_Run_3511,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=50,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=test B2ArmNTS11,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355425 04/14/11 09:44 a-onl Start_Run_3509,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=50,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=Test TwoArmBuffered config / cosmics / Right arm BPM ADC check,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355424 04/14/11 09:16 a-onl Start_Run_3508,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=50,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=Test TwoArmBuffered config / cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355423 04/14/11 09:11 a-onl Start_Run_3507,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=50,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=Test TwoArmBuffered config / cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355422 04/14/11 08:40 R. Michaels some raw data files in raw.dup but not raw
  • 355421 04/14/11 08:33 a-onl End_of_Run_3506
  • 355420 04/14/11 08:10 adaq Beam Time Accounting
  • 355419 04/14/11 08:10 nmuangma owl shift summary
  • 355418 04/14/11 08:05 a-onl Start_Run_3506,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=50,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=Production 50 uA - Deadtime 8% ps1=10 and ps3=7,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355417 04/14/11 08:05 a-onl End_of_Run_3505
  • 355416 04/14/11 07:20 a-onl Start_Run_3505,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=50,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=Production 50 uA - Deadtime 8% ps1=10 and ps3=7,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355415 04/14/11 07:19 a-onl End_of_Run_3504
  • 355414 04/14/11 06:33 a-onl Start_Run_3504,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=50,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=Production 50 uA - Deadtime 8% ps1=10 and ps3=7,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355413 04/14/11 06:32 a-onl End_of_Run_3503
  • 355412 04/14/11 06:11 nmuangma QE scan
  • 355411 04/14/11 05:45 a-onl Start_Run_3503,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=50,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=Production 50 uA - Deadtime 8% ps1=10 and ps3=7,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355410 04/14/11 05:44 a-onl End_of_Run_3502
  • 355409 04/14/11 05:06 nmuangma spot++ 2x2mm
  • 355408 04/14/11 05:03 nmuangma The BigBite is now off
  • 355407 04/14/11 04:58 a-onl Start_Run_3502,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=50,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=Production 50 uA - Deadtime 8% ps1=10 and ps3=7,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355406 04/14/11 04:57 a-onl End_of_Run_3501
  • 355405 04/14/11 04:48 auto_snap Radiation and Ozone snapshot
  • 355404 04/14/11 04:47 auto_snap Miscellaneous snapshot
  • 355403 04/14/11 04:47 auto_snap Hall A Tools snapshot
  • 355402 04/14/11 04:47 auto_snap Compton Polarimeter snapshot
  • 355401 04/14/11 04:47 auto_snap BigBox Magnet snapshot
  • 355400 04/14/11 04:47 auto_snap Beam Position and Current snapshot
  • 355399 04/14/11 04:12 a-onl Start_Run_3501,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=50,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=Production 50 uA - Deadtime 8% ps1=10 and ps3=7,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355398 04/14/11 04:11 a-onl End_of_Run_3500
  • 355397 04/14/11 03:25 a-onl Start_Run_3500,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=50,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=Production 50 uA - Deadtime 8% ps1=10 and ps3=7,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355396 04/14/11 03:24 a-onl End_of_Run_3499
  • 355395 04/14/11 02:41 a-onl Start_Run_3499,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=50,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=Production 50 uA - Deadtime 8% ps1=10 and ps3=7,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355394 04/14/11 02:40 a-onl End_of_Run_3498
  • 355393 04/14/11 02:11 Fatiha Beam current screen
  • 355392 04/14/11 02:04 A. Camsonne Production started at 50 uA
  • 355391 04/14/11 02:03 A. Camsonne Scaler rates 50 uA 20 cm He4 target
  • 355390 04/14/11 02:01 nmuangma BigBite MWDC snap shot
  • 355389 04/14/11 01:58 Fatiha Current Scan list
  • 355388 04/14/11 01:54 a-onl Start_Run_3498,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=50,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=Production 50 uA - Deadtime 8% ps1=10 and ps3=7,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355387 04/14/11 01:53 a-onl End_of_Run_3497
  • 355386 04/14/11 01:50 a-onl Start_Run_3497,Run_type=Check out,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=60,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=Trying 60 uA Current Scan,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355385 04/14/11 01:42 a-onl End_of_Run_3496
  • 355384 04/14/11 01:42 nmuangma BigBite Magnet ON
  • 355383 04/14/11 01:38 a-onl Start_Run_3496,Run_type=Check out,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=5uA Current Scan,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355382 04/14/11 01:32 Leckey Target moved to 20cm Helium
  • 355381 04/14/11 01:27 nmuangma BigBite Optic Position Scan.
  • 355380 04/14/11 01:25 a-onl End_of_Run_3495
  • 355379 04/14/11 01:17 a-onl Start_Run_3495,Run_type=Optics,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=BB optic position magnet on,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355378 04/14/11 01:16 a-onl End_of_Run_3494
  • 355377 04/14/11 01:07 a-onl Start_Run_3494,Run_type=Optics,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=BB optic position,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355376 04/14/11 01:02 nmuangma BigBite Magnet OFF : optic absolute positon scan.
  • 355375 04/14/11 01:01 Leckey Target moved to BeO
  • 355374 04/14/11 00:58 a-onl End_of_Run_3493
  • 355373 04/14/11 00:54 a-onl Start_Run_3493,Run_type=Optics,target_type=Optics: multi 1mm C foils,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=BB optic,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355372 04/14/11 00:53 a-onl End_of_Run_3492
  • 355371 04/14/11 00:53 a-onl Start_Run_3492,Run_type=Check out,target_type=Optics: multi 1mm C foils,beam_e0=4.46,beam_raster=on,beam_current=50,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=BB optic,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355370 04/14/11 00:52 a-onl End_of_Run_3491
  • 355369 04/14/11 00:50 a-onl Start_Run_3491,Run_type=Check out,target_type=Optics: multi 1mm C foils,beam_e0=4.46,beam_raster=on,beam_current=50,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=BB optic,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355368 04/14/11 00:49 Leckey Target moved to Carbon optics
  • 355367 04/14/11 00:38 Leckey Target moved to slanted Carbon
  • 355366 04/14/11 00:38 nmuangma Hall C requests Hall A to turn off the beam
  • 355365 04/14/11 00:33 A. Camsonne Asked to move to C target - Will keep beam off for Hall C to check if Hall A high current running affect width of QWeak
  • 355364 04/14/11 00:29 a-onl End_of_Run_3490
  • 355363 04/14/11 00:28 Fatiha Prescale set for production @ 50 uA
  • 355362 04/14/11 00:17 Miha/Fatiha/ Coincidence timing and Pmiss spectrum @ 40 uA
  • 355361 04/14/11 00:07 a-onl Start_Run_3490,Run_type=Check out,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=50,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=50uA part of current scan,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355360 04/14/11 00:05 S. Sirca Swing shift SUMMARY
  • 355359 04/14/11 00:04 adaq Beam Time Accounting
  • 355358 04/14/11 00:04 a-onl End_of_Run_3489
  • Wednesday

  • 355357 04/13/11 23:58 miham Screen shot of BB and HRS scalers for run #3489 (40uA)
  • 355356 04/13/11 23:57 miham Screen shot of Coda GUI for run #3489 (40uA)
  • 355355 04/13/11 23:55 a-onl Start_Run_3489,Run_type=Check out,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=40,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=40uA part of current scan,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355354 04/13/11 23:53 a-onl End_of_Run_3488
  • 355353 04/13/11 23:51 a-onl Start_Run_3488,Run_type=Check out,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=30uA part of current scan,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355352 04/13/11 23:36 a-onl End_of_Run_3487
  • 355351 04/13/11 23:33 yez target screen
  • 355350 04/13/11 23:32 miham Screen shot of BB and HRS scalers for run #3487 (30uA)
  • 355349 04/13/11 23:31 miham Screen shot of the Coda GUI for run #3487 (30uA)
  • 355348 04/13/11 23:29 a-onl Start_Run_3487,Run_type=Check out,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=30,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=30uA part of current scan,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355347 04/13/11 23:22 a-onl End_of_Run_3486
  • 355346 04/13/11 23:19 miham Screen shot of the Coda GUI for run #3486 (20uA)
  • 355345 04/13/11 23:19 yez L-HRS: Cerenkov Overfload issue resolved, as well as PRL2#14 ADC
  • 355344 04/13/11 23:18 S. Sirca scalers for run 3486
  • 355343 04/13/11 23:17 a-onl Start_Run_3486,Run_type=Check out,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=20,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=20uA part of current scan,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355342 04/13/11 23:16 a-onl End_of_Run_3485
  • 355341 04/13/11 23:14 miham Screen shot of the Coda GUI for run #3485 (20uA)
  • 355340 04/13/11 23:12 S. Sirca scalers for run 3485
  • 355339 04/13/11 23:11 a-onl Start_Run_3485,Run_type=Check out,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=20,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=20uA part of current scan,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355338 04/13/11 23:05 miham Screen shot of the Coda GUI and DT monitor for run #3484 (5uA )
  • 355337 04/13/11 23:05 a-onl End_of_Run_3484
  • 355336 04/13/11 23:03 S. Sirca Scalers run 3484
  • 355335 04/13/11 22:58 a-onl Start_Run_3484,Run_type=Check out,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=Run to check rates and set prescales,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355334 04/13/11 22:56 yez Pedestal Suppression is on after #3483
  • 355333 04/13/11 22:54 S. Sirca spot_L check with run 3483
  • 355332 04/13/11 22:54 a-onl End_of_Run_3483
  • 355331 04/13/11 22:53 yez Turn off pedestal suppression during run#3483
  • 355330 04/13/11 22:52 a-onl Start_Run_3483,Run_type=Spot++,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=Beam 1uA spot++ with prod. target,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355329 04/13/11 22:44 A. Camsonne Pedestal subtraction string added to database for ROC3
  • 355328 04/13/11 22:43 yez target move to 20cm He4
  • 355327 04/13/11 22:42 S. Sirca spot_R check
  • 355326 04/13/11 22:41 a-onl End_of_Run_3482
  • 355325 04/13/11 22:41 S. Sirca spot_L (!) check
  • 355324 04/13/11 22:38 yez Alexandre fixed this issue
  • 355323 04/13/11 22:36 a-onl Start_Run_3482,Run_type=Spot++,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=Beam 5uA spot++ with pulser,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355322 04/13/11 22:35 a-onl End_of_Run_3481
  • 355321 04/13/11 22:35 a-onl Start_Run_3481,Run_type=Spot++,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=Beam 5uA spot++ with pulser,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355320 04/13/11 22:34 a-onl End_of_Run_3480
  • 355319 04/13/11 22:34 S. Sirca How to speed up BPM adjustments
  • 355318 04/13/11 22:33 a-onl Start_Run_3480,Run_type=Spot++,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=Beam 5uA spot++,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355317 04/13/11 22:31 a-onl End_of_Run_3479
  • 355316 04/13/11 22:23 a-onl Start_Run_3479,Run_type=Other,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.510,rightarm_theta=47,rightarm_sieve=out,comment_text=Beam 5uA while tuning,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355315 04/13/11 22:05 yez target move to BeO
  • 355314 04/13/11 22:05 adaq Beam Time Accounting
  • 355313 04/13/11 22:04 S. Sirca Harp B not operational
  • 355312 04/13/11 22:03 a-onl End_of_Run_3478
  • 355311 04/13/11 21:58 yez L-HRS: Pedestal Suppression is not working on ROC3
  • 355310 04/13/11 21:05 yez L.Cer.ADC plot in 13bit mode
  • 355309 04/13/11 20:47 auto_snap Radiation and Ozone snapshot
  • 355308 04/13/11 20:47 auto_snap Miscellaneous snapshot
  • 355307 04/13/11 20:47 auto_snap Hall A Tools snapshot
  • 355306 04/13/11 20:47 auto_snap Compton Polarimeter snapshot
  • 355305 04/13/11 20:47 auto_snap BigBox Magnet snapshot
  • 355304 04/13/11 20:47 auto_snap Beam Position and Current snapshot
  • 355303 04/13/11 20:03 Hanretty Ready, willing and able. However....
  • 355302 04/13/11 19:57 a-onl Start_Run_3478,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Cosmics with pedestal adjusted for ADC in ROC4 slot 21,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355301 04/13/11 19:57 a-onl End_of_Run_3477
  • 355300 04/13/11 19:56 yez Modify pedestal values in ped4.dat
  • 355299 04/13/11 19:35 yez R-HRS: Shower#76 is fixed
  • 355298 04/13/11 19:25 A. Camsonne ADC 1881 added in ROC4 slot 21 - new database directory created
  • 355297 04/13/11 18:52 yez target screen
  • 355296 04/13/11 17:51 a-onl Start_Run_3477,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Cosmics with pedestal adjusted for ADC in ROC4 slot 21,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355295 04/13/11 17:49 a-onl End_of_Run_3476
  • 355294 04/13/11 17:49 A. Camsonne Pedestals for ROC4 ADC and Left arm gas cerenkov
  • 355293 04/13/11 17:35 Folts checklist/tech
  • 355292 04/13/11 17:34 a-onl Start_Run_3476,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355291 04/13/11 17:34 a-onl End_of_Run_3475
  • 355290 04/13/11 17:33 a-onl Start_Run_3475,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355289 04/13/11 17:32 A. Camsonne Pulser run for Cerenkov Left ADC ROC4 Slot 21 Run 3474
  • 355288 04/13/11 17:32 a-onl End_of_Run_3474
  • 355287 04/13/11 17:32 a-onl Start_Run_3474,Run_type=Pedestal,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Pulser run for Left HRS Cerenkov pedestal,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355286 04/13/11 17:23 V. Sulkosky T6 is now BB singles
  • 355285 04/13/11 17:13 a-onl End_of_Run_3473
  • 355284 04/13/11 17:09 a-onl Start_Run_3473,Run_type=Other,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Test DAQ - ADC added in ROC4 slot 21,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355283 04/13/11 17:07 a-onl End_of_Run_3472
  • 355282 04/13/11 17:07 yez target screen after parking LD
  • 355281 04/13/11 17:02 a-onl Start_Run_3472,Run_type=Other,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Test DAQ - ADC added in ROC4 slot 21,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355280 04/13/11 16:55 a-onl End_of_Run_3471
  • 355279 04/13/11 16:51 a-onl Start_Run_3471,Run_type=Other,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Test DAQ - ADC added in ROC4 slot 21,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355278 04/13/11 16:47 a-onl End_of_Run_3470
  • 355277 04/13/11 16:46 a-onl Start_Run_3470,Run_type=Other,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Test DAQ - ADC added in ROC4 slot 21,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355276 04/13/11 16:45 a-onl End_of_Run_3469
  • 355275 04/13/11 16:45 yez target screen
  • 355274 04/13/11 16:39 a-onl Start_Run_3469,Run_type=Other,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Test DAQ - ADC added in ROC4 slot 21,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355273 04/13/11 16:38 a-onl End_of_Run_3468
  • 355272 04/13/11 16:23 a-onl Start_Run_3468,Run_type=Other,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Test DAQ - ADC added in ROC4 slot 21,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355271 04/13/11 16:22 a-onl End_of_Run_3467
  • 355270 04/13/11 16:16 a-onl Start_Run_3467,Run_type=Other,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Test DAQ - ADC added in ROC4 slot 21,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355269 04/13/11 16:13 A. Camsonne Blaster test BigBite crates
  • 355268 04/13/11 16:13 A. Camsonne Blaster test BigBite crates
  • 355267 04/13/11 15:53 a-onl Start_Run_3466,Run_type=Other,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Test DAQ - ADC added in ROC4 slot 21,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355266 04/13/11 15:41 a-onl Start_Run_3465,Run_type=Other,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Test DAQ - ADC added in ROC4 slot 21,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355265 04/13/11 15:40 V. Sulkosky BigBite Moved Out
  • 355264 04/13/11 15:39 V. Sulkosky R-HRS @ 47 degs
  • 355263 04/13/11 15:36 a-onl End_of_Run_3464
  • 355262 04/13/11 15:33 a-onl Start_Run_3464,Run_type=Other,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Test DAQ - ADC added in ROC4 slot 21,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355261 04/13/11 15:24 a-onl End_of_Run_3463
  • 355260 04/13/11 15:21 a-onl Start_Run_3463,Run_type=Other,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Test DAQ - ADC added in ROC4 slot 21,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355259 04/13/11 15:20 a-onl End_of_Run_3462
  • 355258 04/13/11 15:17 a-onl Start_Run_3462,Run_type=Other,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Test DAQ - ADC added in ROC4 slot 21,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355257 04/13/11 15:14 a-onl End_of_Run_3461
  • 355256 04/13/11 15:14 a-onl Start_Run_3461,Run_type=Other,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Test DAQ - ADC added in ROC4 slot 21,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355255 04/13/11 15:11 a-onl End_of_Run_3460
  • 355254 04/13/11 15:09 a-onl Start_Run_3460,Run_type=Other,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Test DAQ - ADC added in ROC4 slot 21,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355253 04/13/11 15:05 yez R-HRS: swap Sh.76 & Sh.80 at the patch panel upstair
  • 355252 04/13/11 15:00 A. Camsonne Bigbite network switch replaced
  • 355251 04/13/11 14:36 yez New DB ~/20110413/ is for L-Cerenkov ADC change, only for future experiment.
  • 355250 04/13/11 14:34 a-onl End_of_Run_3456
  • 355249 04/13/11 14:32 orchen 750MeV/c summery
  • 355248 04/13/11 14:32 a-onl End_of_Run_3456
  • 355247 04/13/11 14:31 adq ADC cardS changes
  • 355246 04/13/11 14:23 yez L-HRS: Reduce HV of PRL2#14 from -908V to 850V
  • 355245 04/13/11 14:22 a-onl Start_Run_3456,Run_type=Other,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Test DAQ - ADC added in ROC4 slot 21,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355244 04/13/11 14:04 yez L-HRS: Move the ADC card in RCO3.Slot21 to ROC4.Slot21, and move Cer ADC cables
  • 355243 04/13/11 13:55 A. Camsonne HRS crl and pedestal files backed up in directory sav_13apr2011
  • 355242 04/13/11 13:29 a-onl End_of_Run_3455
  • 355241 04/13/11 13:02 orchen BB sieve in
  • 355240 04/13/11 12:47 auto_snap Radiation and Ozone snapshot
  • 355239 04/13/11 12:47 auto_snap Miscellaneous snapshot
  • 355238 04/13/11 12:47 auto_snap Hall A Tools snapshot
  • 355237 04/13/11 12:47 auto_snap Compton Polarimeter snapshot
  • 355236 04/13/11 12:47 auto_snap BigBox Magnet snapshot
  • 355235 04/13/11 12:47 auto_snap Beam Position and Current snapshot
  • 355234 04/13/11 12:02 posik target moved to empty
  • 355233 04/13/11 11:08 adaq SRC run stoped sorry it was not ended
  • 355232 04/13/11 11:05 posik target screen
  • 355231 04/13/11 11:05 Annand Beam off: run 3455 cosmics
  • 355230 04/13/11 11:05 Hanretty Beam Studies
  • 355229 04/13/11 11:04 a-onl Start_Run_3455,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355228 04/13/11 11:04 posik target strip charts
  • 355227 04/13/11 11:02 J. Watson Spot++ for Run 3454
  • 355226 04/13/11 11:01 a-onl End_of_Run_3454
  • 355225 04/13/11 09:55 a-onl Start_Run_3454,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355224 04/13/11 09:54 a-onl End_of_Run_3453
  • 355223 04/13/11 09:24 V. Sulkosky Cycled L11.3
  • 355222 04/13/11 09:20 R. Michaels never mind
  • 355221 04/13/11 08:38 a-onl Start_Run_3453,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355220 04/13/11 08:37 a-onl End_of_Run_3452
  • 355219 04/13/11 08:07 adaq Beam Time Accounting
  • 355218 04/13/11 07:41 nmuangma owl shift summary
  • 355217 04/13/11 07:22 a-onl Start_Run_3452,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355216 04/13/11 07:22 a-onl End_of_Run_3451
  • 355215 04/13/11 06:06 a-onl Start_Run_3451,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355214 04/13/11 06:04 a-onl End_of_Run_3450
  • 355213 04/13/11 06:03 nmuangma QE scan (MCC)
  • 355212 04/13/11 04:47 a-onl Start_Run_3450,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355211 04/13/11 04:47 auto_snap Radiation and Ozone snapshot
  • 355210 04/13/11 04:47 auto_snap Miscellaneous snapshot
  • 355209 04/13/11 04:47 auto_snap Hall A Tools snapshot
  • 355208 04/13/11 04:46 a-onl End_of_Run_3449
  • 355207 04/13/11 04:46 auto_snap Compton Polarimeter snapshot
  • 355206 04/13/11 04:46 auto_snap BigBox Magnet snapshot
  • 355205 04/13/11 04:46 auto_snap Beam Position and Current snapshot
  • 355204 04/13/11 03:36 Leckey Bigbite and Scaler Snapshot
  • 355203 04/13/11 03:35 Leckey Neutron Detector Snapshot
  • 355202 04/13/11 03:33 Leckey Target Snapshot
  • 355201 04/13/11 03:29 a-onl Start_Run_3449,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355200 04/13/11 03:29 a-onl End_of_Run_3448
  • 355199 04/13/11 02:09 a-onl Start_Run_3448,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355198 04/13/11 02:08 a-onl End_of_Run_3447
  • 355197 04/13/11 01:52 nmuangma half wave plate OUT (requested by hall C)
  • 355196 04/13/11 01:39 orchen BB and HAND thresholds
  • 355195 04/13/11 01:01 orchen BB time resolution and number of events for 750MeV/c kinematics
  • 355194 04/13/11 00:51 a-onl Start_Run_3447,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355193 04/13/11 00:50 a-onl End_of_Run_3446
  • 355192 04/13/11 00:07 adaq Beam Time Accounting
  • 355191 04/13/11 00:06 S. Sirca Swing shift SUMMARY
  • Tuesday

  • 355190 04/12/11 23:42 miham Screen shot of HAND scalers for run #3446
  • 355189 04/12/11 23:40 miham Screen shots of BB and HRS scalers for run #3446
  • 355188 04/12/11 23:37 miham Screen shot of Coda GUI during run #3446
  • 355187 04/12/11 23:36 a-onl Start_Run_3446,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355186 04/12/11 23:35 a-onl End_of_Run_3445
  • 355185 04/12/11 23:16 miham Screen shot of HAND scalers for run #3445
  • 355184 04/12/11 23:15 miham Screen shot of BB and HRS scalers for run #3445
  • 355183 04/12/11 23:14 miham Screen shot of Coda GUI for run #3445
  • 355182 04/12/11 22:57 orchen pion rejector is fine - root plotting issue
  • 355181 04/12/11 22:16 a-onl Start_Run_3445,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355180 04/12/11 22:14 S. Sirca spot++
  • 355179 04/12/11 22:14 a-onl End_of_Run_3444
  • 355178 04/12/11 21:28 miham Problem with Pion Rejector on HRS-L
  • 355177 04/12/11 21:02 miham Screen shot of HAND scalers for run #3444
  • 355176 04/12/11 21:01 miham Screen shot of BB and HRS scalers for #3444
  • 355175 04/12/11 20:59 miham Screen shot of Coda GUI for #3444
  • 355174 04/12/11 20:57 a-onl Start_Run_3444,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355173 04/12/11 20:56 a-onl End_of_Run_3443
  • 355172 04/12/11 20:46 auto_snap Radiation and Ozone snapshot
  • 355171 04/12/11 20:46 auto_snap Miscellaneous snapshot
  • 355170 04/12/11 20:46 auto_snap Hall A Tools snapshot
  • 355169 04/12/11 20:46 auto_snap Compton Polarimeter snapshot
  • 355168 04/12/11 20:46 auto_snap BigBox Magnet snapshot
  • 355167 04/12/11 20:46 auto_snap Beam Position and Current snapshot
  • 355166 04/12/11 19:55 miham Channel 6 in plane N4 tripped again
  • 355165 04/12/11 19:49 miham HAND scalers screen shot for run #3443
  • 355164 04/12/11 19:46 miham BB and HRS Scalers screen shot for run #3443
  • 355163 04/12/11 19:43 miham DAQ GUI screen shot for prod. run #3443
  • 355162 04/12/11 19:39 a-onl Start_Run_3443,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355161 04/12/11 19:38 JinGe target is moved to loop 1 4He 20 cm
  • 355160 04/12/11 19:31 a-onl End_of_Run_3442
  • 355159 04/12/11 18:47 miham DAQ GUI screenshot for Al dummy run #3442
  • 355158 04/12/11 18:40 miham HAND scaler screen shot for Al 15cm dummy run #3442
  • 355157 04/12/11 18:32 miham BB and HRS-L scalers for run #3442 (Al 15cm Dummy)
  • 355156 04/12/11 18:28 a-onl Start_Run_3442,Run_type=Production SRC,target_type=20cm dummy,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Aluminum 15cm dummy target ,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355155 04/12/11 18:26 JinGe target is moved to dummy 15 cm
  • 355154 04/12/11 18:18 a-onl End_of_Run_3441
  • 355153 04/12/11 17:48 yez Yes, it is normal.
  • 355152 04/12/11 17:34 miham Cherenkov spectrum different than the golden run
  • 355151 04/12/11 17:13 adaq Beam Time Accounting
  • 355150 04/12/11 17:01 a-onl Start_Run_3441,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355149 04/12/11 17:00 a-onl End_of_Run_3440
  • 355148 04/12/11 16:52 miham Problem with HAND, Plane N6, Bar 3
  • 355147 04/12/11 16:41 Jack (Techs) Hall probe readbacks reset
  • 355146 04/12/11 16:31 yez R-HRS Optics
  • 355145 04/12/11 16:29 S. Sirca HRS magnet IOCs rebooted
  • 355144 04/12/11 15:52 BoZhao The issue of iocha22 software reboot
  • 355143 04/12/11 15:41 a-onl Start_Run_3440,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355142 04/12/11 15:40 a-onl End_of_Run_3439
  • 355141 04/12/11 14:24 a-onl Start_Run_3439,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355140 04/12/11 14:24 a-onl End_of_Run_3438
  • 355139 04/12/11 14:21 Watson Spot++
  • 355138 04/12/11 13:30 Annand Hall-C auto steer
  • 355137 04/12/11 13:28 posik target screen shot
  • 355136 04/12/11 13:21 Annand 1/2 wave plate Out
  • 355135 04/12/11 13:03 a-onl Start_Run_3438,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355134 04/12/11 13:02 a-onl End_of_Run_3437
  • 355133 04/12/11 12:46 auto_snap Radiation and Ozone snapshot
  • 355132 04/12/11 12:46 auto_snap Miscellaneous snapshot
  • 355131 04/12/11 12:46 auto_snap Hall A Tools snapshot
  • 355130 04/12/11 12:46 auto_snap Compton Polarimeter snapshot
  • 355129 04/12/11 12:46 auto_snap BigBox Magnet snapshot
  • 355128 04/12/11 12:46 auto_snap Beam Position and Current snapshot
  • 355127 04/12/11 11:42 a-onl Start_Run_3437,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355126 04/12/11 11:41 a-onl End_of_Run_3436
  • 355125 04/12/11 11:20 Annand BigBite scaler display
  • 355124 04/12/11 10:52 Annand 3435 replayT5.C OK
  • 355123 04/12/11 10:26 Annand Run 3435 end crash
  • 355122 04/12/11 10:23 a-onl Start_Run_3436,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355121 04/12/11 10:12 a-onl End_of_Run_3435
  • 355120 04/12/11 10:12 a-onl End_of_Run_3435
  • 355119 04/12/11 08:57 a-onl Start_Run_3435,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355118 04/12/11 08:55 a-onl End_of_Run_3434
  • 355117 04/12/11 08:52 R. Michaels Run plan for this week and beyond
  • 355116 04/12/11 08:09 kai target runs pretty well in whole shift
  • 355115 04/12/11 08:08 adaq Beam Time Accounting
  • 355114 04/12/11 07:55 nmuangma owl shift summary
  • 355113 04/12/11 07:40 a-onl Start_Run_3434,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355112 04/12/11 07:39 a-onl End_of_Run_3433
  • 355111 04/12/11 06:51 nmuangma QE scan (requested from MCC)
  • 355110 04/12/11 06:23 nmuangma notice Zero reading for HAND Left N6 bar 2. reset the HV
  • 355109 04/12/11 06:22 a-onl Start_Run_3433,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355108 04/12/11 06:17 a-onl End_of_Run_3431
  • 355107 04/12/11 05:23 a-onl Start_Run_3431,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355106 04/12/11 05:22 a-onl End_of_Run_3430
  • 355105 04/12/11 04:46 auto_snap Radiation and Ozone snapshot
  • 355104 04/12/11 04:46 auto_snap Miscellaneous snapshot
  • 355103 04/12/11 04:46 auto_snap Hall A Tools snapshot
  • 355102 04/12/11 04:46 auto_snap Compton Polarimeter snapshot
  • 355101 04/12/11 04:45 auto_snap BigBox Magnet snapshot
  • 355100 04/12/11 04:45 auto_snap Beam Position and Current snapshot
  • 355099 04/12/11 04:06 a-onl Start_Run_3430,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355098 04/12/11 04:02 a-onl End_of_Run_3429
  • 355097 04/12/11 04:02 a-onl End_of_Run_3429
  • 355096 04/12/11 03:49 kai target snapshot
  • 355095 04/12/11 03:00 Leckey Neutron Detector Snapshot
  • 355094 04/12/11 02:58 Leckey Bigbite + Scaler Snapshot
  • 355093 04/12/11 02:46 a-onl Start_Run_3429,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355092 04/12/11 02:45 a-onl End_of_Run_3428
  • 355091 04/12/11 01:30 a-onl Start_Run_3428,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355090 04/12/11 01:30 a-onl End_of_Run_3427
  • 355089 04/12/11 01:24 nmuangma half wave plate in (requested by hall C)
  • 355088 04/12/11 00:14 a-onl Start_Run_3427,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355087 04/12/11 00:13 a-onl End_of_Run_3426
  • 355086 04/12/11 00:05 adaq Beam Time Accounting
  • 355085 04/12/11 00:01 Norum Shift summary
  • Monday

  • 355084 04/11/11 22:57 a-onl Start_Run_3426,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355083 04/11/11 22:56 a-onl End_of_Run_3425
  • 355082 04/11/11 21:41 a-onl Start_Run_3425,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355081 04/11/11 21:39 a-onl End_of_Run_3424
  • 355080 04/11/11 20:45 auto_snap Radiation and Ozone snapshot
  • 355079 04/11/11 20:45 auto_snap Miscellaneous snapshot
  • 355078 04/11/11 20:45 auto_snap Hall A Tools snapshot
  • 355077 04/11/11 20:45 auto_snap Compton Polarimeter snapshot
  • 355076 04/11/11 20:45 auto_snap BigBox Magnet snapshot
  • 355075 04/11/11 20:45 auto_snap Beam Position and Current snapshot
  • 355074 04/11/11 20:25 a-onl Start_Run_3424,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355073 04/11/11 20:24 a-onl End_of_Run_3423
  • 355072 04/11/11 20:16 JinGe target snapshot
  • 355071 04/11/11 19:06 a-onl Start_Run_3423,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355070 04/11/11 19:04 a-onl End_of_Run_3422
  • 355069 04/11/11 18:50 V. Sulkosky Golden Rootfile now 3199
  • 355068 04/11/11 17:57 V. Sulkosky /adaql10/work1 partial clean up
  • 355067 04/11/11 17:45 a-onl Start_Run_3422,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355066 04/11/11 17:44 a-onl End_of_Run_3421
  • 355065 04/11/11 16:27 a-onl Start_Run_3421,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355064 04/11/11 16:24 a-onl End_of_Run_3420
  • 355063 04/11/11 16:13 adaq Beam Time Accounting
  • 355062 04/11/11 16:02 a-onl Start_Run_3420,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355061 04/11/11 16:01 a-onl End_of_Run_3419
  • 355060 04/11/11 15:53 Annand Day Shift Summary
  • 355059 04/11/11 15:37 yez target screen
  • 355058 04/11/11 15:31 Fansler LQ1
  • 355057 04/11/11 14:55 Annand 1/2 wave plate out
  • 355056 04/11/11 14:46 a-onl Start_Run_3419,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355055 04/11/11 14:45 a-onl End_of_Run_3418
  • 355054 04/11/11 13:27 a-onl Start_Run_3418,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355053 04/11/11 13:26 Annand HAND HV trip
  • 355052 04/11/11 13:23 a-onl End_of_Run_3417
  • 355051 04/11/11 13:08 Annand spot++
  • 355050 04/11/11 12:45 auto_snap Radiation and Ozone snapshot
  • 355049 04/11/11 12:45 auto_snap Miscellaneous snapshot
  • 355048 04/11/11 12:45 auto_snap Hall A Tools snapshot
  • 355047 04/11/11 12:45 auto_snap Compton Polarimeter snapshot
  • 355046 04/11/11 12:45 auto_snap BigBox Magnet snapshot
  • 355045 04/11/11 12:45 auto_snap Beam Position and Current snapshot
  • 355044 04/11/11 12:44 a-onl Start_Run_3417,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355043 04/11/11 12:43 a-onl End_of_Run_3416
  • 355042 04/11/11 12:23 dhamilton /adaql10/work1 full - moved some old data
  • 355041 04/11/11 12:22 yez target screen
  • 355040 04/11/11 11:22 a-onl Start_Run_3416,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355039 04/11/11 11:21 a-onl End_of_Run_3415
  • 355038 04/11/11 10:08 Annand spot++ after beam restore
  • 355037 04/11/11 10:06 Annand Q1 Left current restored
  • 355036 04/11/11 10:04 Annand Beam returned
  • 355035 04/11/11 10:02 a-onl Start_Run_3415,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355034 04/11/11 10:01 a-onl End_of_Run_3414
  • 355033 04/11/11 09:54 a-onl Start_Run_3414,Run_type=Spot++,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=spot++,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355032 04/11/11 09:07 Annand Hall-A access check He system
  • 355031 04/11/11 08:22 yez target screen
  • 355030 04/11/11 08:17 Annand He level Q1 HRS L
  • 355029 04/11/11 08:14 Annand Lost Beam around 45 min
  • 355028 04/11/11 08:13 a-onl End_of_Run_3413
  • 355027 04/11/11 08:11 adaq Beam Time Accounting
  • 355026 04/11/11 08:05 a-onl Start_Run_3413,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355025 04/11/11 08:03 a-onl End_of_Run_3412
  • 355024 04/11/11 08:00 nmuangma owl shift summary
  • 355023 04/11/11 07:43 nmuangma Liquid level on Left Q1 is low: Shift Leader must keep an eye for it.
  • 355022 04/11/11 06:47 a-onl Start_Run_3412,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355021 04/11/11 06:46 a-onl End_of_Run_3411
  • 355020 04/11/11 05:29 a-onl Start_Run_3411,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355019 04/11/11 05:25 a-onl End_of_Run_3410
  • 355018 04/11/11 05:07 kelleher BigBite Singles and scalers display
  • 355017 04/11/11 05:04 kelleher Neutron Singles Snapshot
  • 355016 04/11/11 05:02 a-onl Start_Run_3410,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355015 04/11/11 05:01 a-onl End_of_Run_3409
  • 355014 04/11/11 04:45 auto_snap Radiation and Ozone snapshot
  • 355013 04/11/11 04:45 auto_snap Miscellaneous snapshot
  • 355012 04/11/11 04:45 auto_snap Hall A Tools snapshot
  • 355011 04/11/11 04:45 auto_snap Compton Polarimeter snapshot
  • 355010 04/11/11 04:44 auto_snap BigBox Magnet snapshot
  • 355009 04/11/11 04:44 auto_snap Beam Position and Current snapshot
  • 355008 04/11/11 04:14 Leckey Target Snapshot
  • 355007 04/11/11 03:47 a-onl Start_Run_3409,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355006 04/11/11 03:46 a-onl End_of_Run_3408
  • 355005 04/11/11 02:31 a-onl Start_Run_3408,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 355004 04/11/11 02:30 a-onl End_of_Run_3407
  • 355003 04/11/11 02:23 orchen Triple coincedence status, 750MeV/c kinematics, 2.7C
  • 355002 04/11/11 02:18 nmuangma spot ++ 2x2mm
  • 355001 04/11/11 02:10 orchen Pmiss coverage.
  • 355000 04/11/11 01:14 a-onl Start_Run_3407,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354999 04/11/11 01:13 a-onl End_of_Run_3406
  • 354998 04/11/11 00:25 nmuangma HANDs HV trip at 2008: L9.6 twice
  • 354997 04/11/11 00:03 adaq Beam Time Accounting
  • Sunday

  • 354996 04/10/11 23:57 Ramesh Shift Summary
  • 354995 04/10/11 23:54 a-onl Start_Run_3406,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354994 04/10/11 23:52 a-onl End_of_Run_3405
  • 354993 04/10/11 23:33 wdconinc Target alarm handler: mute...
  • 354992 04/10/11 22:38 a-onl Start_Run_3405,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354991 04/10/11 22:37 a-onl End_of_Run_3404
  • 354990 04/10/11 21:58 Fatiha Neutron Detector Single rates
  • 354989 04/10/11 21:56 Fatiha Scaler Data screen and BigBite single rates
  • 354988 04/10/11 21:29 Ramesh spot check for checklist
  • 354987 04/10/11 21:22 a-onl Start_Run_3404,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354986 04/10/11 21:12 a-onl End_of_Run_3403
  • 354985 04/10/11 20:44 auto_snap Radiation and Ozone snapshot
  • 354984 04/10/11 20:44 auto_snap Miscellaneous snapshot
  • 354983 04/10/11 20:44 auto_snap Hall A Tools snapshot
  • 354982 04/10/11 20:44 auto_snap Compton Polarimeter snapshot
  • 354981 04/10/11 20:44 auto_snap BigBox Magnet snapshot
  • 354980 04/10/11 20:44 auto_snap Beam Position and Current snapshot
  • 354979 04/10/11 19:56 a-onl Start_Run_3403,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354978 04/10/11 19:55 a-onl End_of_Run_3402
  • 354977 04/10/11 18:59 wdconinc Target is alive and well (incl screen shots)
  • 354976 04/10/11 18:40 a-onl Start_Run_3402,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354975 04/10/11 18:28 a-onl End_of_Run_3401
  • 354974 04/10/11 17:38 Ramesh BSY dropped to power permit, now back to business
  • 354973 04/10/11 17:35 a-onl Start_Run_3401,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354972 04/10/11 17:31 a-onl End_of_Run_3400
  • 354971 04/10/11 16:11 Annand Day Shift Summary
  • 354970 04/10/11 16:03 adaq Beam Time Accounting
  • 354969 04/10/11 16:02 a-onl Start_Run_3400,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354968 04/10/11 16:01 a-onl End_of_Run_3399
  • 354967 04/10/11 14:57 johans ROC sync plots
  • 354966 04/10/11 14:45 a-onl Start_Run_3399,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354965 04/10/11 14:44 a-onl End_of_Run_3398
  • 354964 04/10/11 14:39 R. Michaels re: run 3394 data file
  • 354963 04/10/11 13:26 a-onl Start_Run_3398,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354962 04/10/11 13:26 a-onl End_of_Run_3397
  • 354961 04/10/11 13:07 Annand spot++ for run 3397
  • 354960 04/10/11 12:53 Annand 1/2 Wave plate IN
  • 354959 04/10/11 12:44 auto_snap Radiation and Ozone snapshot
  • 354958 04/10/11 12:44 auto_snap Miscellaneous snapshot
  • 354957 04/10/11 12:44 auto_snap Hall A Tools snapshot
  • 354956 04/10/11 12:44 auto_snap Compton Polarimeter snapshot
  • 354955 04/10/11 12:44 auto_snap BigBox Magnet snapshot
  • 354954 04/10/11 12:44 auto_snap Beam Position and Current snapshot
  • 354953 04/10/11 12:06 a-onl Start_Run_3397,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354952 04/10/11 12:06 a-onl End_of_Run_3396
  • 354951 04/10/11 11:12 johans Datafiles for runs 3393 and 3394
  • 354950 04/10/11 10:49 a-onl Start_Run_3396,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354949 04/10/11 10:48 a-onl End_of_Run_3395
  • 354948 04/10/11 09:57 Nikos Target screens
  • 354947 04/10/11 09:50 johans Coda run number synch problem
  • 354946 04/10/11 09:30 a-onl Start_Run_3395,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354945 04/10/11 09:29 a-onl End_of_Run_3394
  • 354944 04/10/11 09:13 a-onl Start_Run_3394,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354943 04/10/11 09:09 Annand Short beam interruption
  • 354942 04/10/11 09:08 a-onl End_of_Run_3393
  • 354941 04/10/11 08:06 adaq Beam Time Accounting
  • 354940 04/10/11 07:59 nmuangma owl shift summary
  • 354939 04/10/11 07:55 a-onl Start_Run_3393,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354938 04/10/11 07:49 a-onl End_of_Run_3392
  • 354937 04/10/11 07:39 nmuangma nmuangma
  • 354936 04/10/11 07:39 nmuangma spot++
  • 354935 04/10/11 06:31 a-onl Start_Run_3392,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354934 04/10/11 06:29 a-onl End_of_Run_3391
  • 354933 04/10/11 05:14 a-onl Start_Run_3391,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354932 04/10/11 05:13 a-onl End_of_Run_3390
  • 354931 04/10/11 04:44 auto_snap Radiation and Ozone snapshot
  • 354930 04/10/11 04:44 auto_snap Miscellaneous snapshot
  • 354929 04/10/11 04:44 auto_snap Hall A Tools snapshot
  • 354928 04/10/11 04:44 auto_snap Compton Polarimeter snapshot
  • 354927 04/10/11 04:44 auto_snap BigBox Magnet snapshot
  • 354926 04/10/11 04:43 auto_snap Beam Position and Current snapshot
  • 354925 04/10/11 03:59 a-onl Start_Run_3390,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354924 04/10/11 03:58 a-onl End_of_Run_3389
  • 354923 04/10/11 02:43 a-onl Start_Run_3389,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354922 04/10/11 02:42 a-onl End_of_Run_3388
  • 354921 04/10/11 01:27 a-onl Start_Run_3388,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354920 04/10/11 01:25 a-onl End_of_Run_3387
  • 354919 04/10/11 00:31 nmuangma bbps1:2001: L10.2 is trip Again at 12:31
  • 354918 04/10/11 00:24 nmuangma bbps1:2001: L10.2 is trip
  • 354917 04/10/11 00:10 a-onl Start_Run_3387,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354916 04/10/11 00:09 a-onl End_of_Run_3386
  • 354915 04/10/11 00:04 adaq Beam Time Accounting
  • 354914 04/10/11 00:03 Ramesh Shift summary
  • Saturday

  • 354913 04/09/11 23:33 Fatiha Neutron detector single rates
  • 354912 04/09/11 23:33 Fatiha Scaler screen
  • 354911 04/09/11 23:22 Nuruzzaman Target screens.
  • 354910 04/09/11 23:17 Ramesh general tools
  • 354909 04/09/11 23:14 Ramesh spot check for checklist
  • 354908 04/09/11 22:54 a-onl Start_Run_3386,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354907 04/09/11 22:53 a-onl End_of_Run_3385
  • 354906 04/09/11 21:38 a-onl Start_Run_3385,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354905 04/09/11 21:36 a-onl End_of_Run_3384
  • 354904 04/09/11 20:43 auto_snap Radiation and Ozone snapshot
  • 354903 04/09/11 20:43 auto_snap Miscellaneous snapshot
  • 354902 04/09/11 20:43 auto_snap Hall A Tools snapshot
  • 354901 04/09/11 20:43 auto_snap Compton Polarimeter snapshot
  • 354900 04/09/11 20:43 auto_snap BigBox Magnet snapshot
  • 354899 04/09/11 20:43 auto_snap Beam Position and Current snapshot
  • 354898 04/09/11 20:29 Nuruzzaman Target screens.
  • 354897 04/09/11 19:03 a-onl Start_Run_3384,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354896 04/09/11 18:09 a-onl Start_Run_3383,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354895 04/09/11 18:08 a-onl End_of_Run_3382
  • 354894 04/09/11 16:54 a-onl Start_Run_3382,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354893 04/09/11 16:53 a-onl End_of_Run_3381
  • 354892 04/09/11 16:29 Nuruzzaman Target screens at begining of swing shift.
  • 354891 04/09/11 16:03 adaq Beam Time Accounting
  • 354890 04/09/11 15:57 Annand Shift Summary
  • 354889 04/09/11 15:37 a-onl Start_Run_3381,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354888 04/09/11 15:37 a-onl End_of_Run_3380
  • 354887 04/09/11 14:16 a-onl Start_Run_3380,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354886 04/09/11 14:15 a-onl End_of_Run_3379
  • 354885 04/09/11 13:18 Annand 1/2 wave plate out
  • 354884 04/09/11 12:55 a-onl Start_Run_3379,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354883 04/09/11 12:54 a-onl End_of_Run_3378
  • 354882 04/09/11 12:48 johans End of run report appeared.
  • 354881 04/09/11 12:43 auto_snap Radiation and Ozone snapshot
  • 354880 04/09/11 12:43 auto_snap Miscellaneous snapshot
  • 354879 04/09/11 12:43 auto_snap Hall A Tools snapshot
  • 354878 04/09/11 12:43 auto_snap Compton Polarimeter snapshot
  • 354877 04/09/11 12:43 auto_snap BigBox Magnet snapshot
  • 354876 04/09/11 12:43 auto_snap Beam Position and Current snapshot
  • 354875 04/09/11 12:40 a-onl End_of_Run_3377
  • 354874 04/09/11 12:24 Johans End of run post to log error
  • 354873 04/09/11 12:17 Annand Spott++
  • 354872 04/09/11 11:39 a-onl Start_Run_3378,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354871 04/09/11 10:22 a-onl Start_Run_3377,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354870 04/09/11 10:21 a-onl End_of_Run_3376
  • 354869 04/09/11 09:37 Nikos Target screens
  • 354868 04/09/11 09:02 a-onl Start_Run_3376,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354867 04/09/11 09:01 a-onl End_of_Run_3375
  • 354866 04/09/11 08:07 adaq Beam Time Accounting
  • 354865 04/09/11 08:04 nmuangma owl shift summary
  • 354864 04/09/11 07:52 A. Shahinyan Target status
  • 354863 04/09/11 07:44 a-onl Start_Run_3375,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354862 04/09/11 07:44 a-onl End_of_Run_3374
  • 354861 04/09/11 06:58 nmuangma One bottle of the Argon gas is almost out. The other is still full.
  • 354860 04/09/11 06:51 nmuangma Argon pressure is LOW
  • 354859 04/09/11 06:28 a-onl Start_Run_3374,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354858 04/09/11 06:26 a-onl End_of_Run_3373
  • 354857 04/09/11 06:18 a-onl Start_Run_3373,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354856 04/09/11 06:17 a-onl End_of_Run_3372
  • 354855 04/09/11 05:00 a-onl Start_Run_3372,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354854 04/09/11 05:00 nmuangma hv trip for HANDs L10.2
  • 354853 04/09/11 04:59 a-onl End_of_Run_3371
  • 354852 04/09/11 04:53 nmuangma half-wave plate is now in
  • 354851 04/09/11 04:43 auto_snap Radiation and Ozone snapshot
  • 354850 04/09/11 04:43 auto_snap Miscellaneous snapshot
  • 354849 04/09/11 04:43 auto_snap Hall A Tools snapshot
  • 354848 04/09/11 04:43 auto_snap Compton Polarimeter snapshot
  • 354847 04/09/11 04:43 auto_snap BigBox Magnet snapshot
  • 354846 04/09/11 04:43 auto_snap Beam Position and Current snapshot
  • 354845 04/09/11 04:42 orchen The quest for more chrge at 750MeV/c kinematics.
  • 354844 04/09/11 03:44 a-onl Start_Run_3371,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354843 04/09/11 03:43 a-onl End_of_Run_3370
  • 354842 04/09/11 02:51 A. Shahinyan Target status
  • 354841 04/09/11 02:27 a-onl Start_Run_3370,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354840 04/09/11 02:26 a-onl End_of_Run_3369
  • 354839 04/09/11 01:10 a-onl Start_Run_3369,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354838 04/09/11 01:08 a-onl End_of_Run_3368
  • 354837 04/09/11 00:25 A. Shahinyan Target status
  • 354836 04/09/11 00:04 adaq Beam Time Accounting
  • 354835 04/09/11 00:02 V. Sulkosky Swing Shift Summary (April 8, 2011)
  • Friday

  • 354834 04/08/11 23:53 a-onl Start_Run_3368,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354833 04/08/11 23:52 a-onl End_of_Run_3367
  • 354832 04/08/11 23:22 V. Sulkosky bbps1:2008 L9:6 tripped again.
  • 354831 04/08/11 23:21 V. Sulkosky N5 R3 Nosier
  • 354830 04/08/11 23:16 C. Gu Screenshot for scintillator and scaler
  • 354829 04/08/11 23:15 C. Gu screenshot for netron decector
  • 354828 04/08/11 23:07 V. Sulkosky Spot++ Run 3367
  • 354827 04/08/11 22:40 V. Sulkosky Beam Back
  • 354826 04/08/11 22:38 a-onl Start_Run_3367,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354825 04/08/11 22:28 V. Sulkosky Beam Off for Cavity Bypass
  • 354824 04/08/11 22:26 a-onl End_of_Run_3366
  • 354823 04/08/11 22:02 Nuruzzaman Target Screens.
  • 354822 04/08/11 21:48 a-onl Start_Run_3366,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354821 04/08/11 21:45 a-onl End_of_Run_3365
  • 354820 04/08/11 21:15 V. Sulkosky Left HRS HV GUI
  • 354819 04/08/11 21:05 C. Gu Screenshot for scintillator and scaler
  • 354818 04/08/11 21:04 C. Gu Screenshot for neutron detector
  • 354817 04/08/11 20:42 auto_snap Radiation and Ozone snapshot
  • 354816 04/08/11 20:42 auto_snap Miscellaneous snapshot
  • 354815 04/08/11 20:42 auto_snap Hall A Tools snapshot
  • 354814 04/08/11 20:42 auto_snap Compton Polarimeter snapshot
  • 354813 04/08/11 20:42 auto_snap BigBox Magnet snapshot
  • 354812 04/08/11 20:42 auto_snap Beam Position and Current snapshot
  • 354811 04/08/11 20:27 a-onl Start_Run_3365,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354810 04/08/11 20:26 a-onl End_of_Run_3364
  • 354809 04/08/11 20:14 Nuruzzaman bbps1:2001 L10:2 trip again.
  • 354808 04/08/11 20:09 Nuruzzaman bbps1:2001 L10:2 trip again.
  • 354807 04/08/11 19:53 C. Gu Screenshot for scintillator and scaler
  • 354806 04/08/11 19:52 Nuruzzaman bbps1:2001 L10:2 tripped during run 3364.
  • 354805 04/08/11 19:51 C. Gu Screenshot for neutron detector
  • 354804 04/08/11 19:11 Nuruzzaman Target Screens.
  • 354803 04/08/11 19:10 a-onl Start_Run_3364,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354802 04/08/11 19:08 a-onl End_of_Run_3363
  • 354801 04/08/11 18:43 C. Gu screenshot for scintillator and scaler
  • 354800 04/08/11 18:42 C. Gu screenshot for neutron detector
  • 354799 04/08/11 18:23 V. Sulkosky Hall A Tools
  • 354798 04/08/11 18:00 V. Sulkosky bbps1:2008 L9.6 Trips
  • 354797 04/08/11 17:58 V. Sulkosky Left-Side HAND HV tripped bbps1:2008
  • 354796 04/08/11 17:54 a-onl Start_Run_3363,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354795 04/08/11 17:52 a-onl End_of_Run_3362
  • 354794 04/08/11 17:32 a-onl End_of_Run_3361
  • 354793 04/08/11 16:59 C. Gu screenshot for scintillator and scaler
  • 354792 04/08/11 16:56 C. Gu screenshot for neutron detector
  • 354791 04/08/11 16:17 a-onl Start_Run_3361,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354790 04/08/11 16:16 a-onl End_of_Run_3360
  • 354789 04/08/11 16:05 adaq Beam Time Accounting
  • 354788 04/08/11 16:05 dhamilton Day Shift Summary
  • 354787 04/08/11 15:57 Nuruzzaman Target screens at begining of swing shift.
  • 354786 04/08/11 15:42 dhamilton Spot++ Screenshot
  • 354785 04/08/11 15:40 dhamilton Hall A Tools Screenshot
  • 354784 04/08/11 15:35 yez target screen
  • 354783 04/08/11 15:25 orchen Triple Coincedence for 1.92C, 750MeV/c Kinematics, Friday April 8th.
  • 354782 04/08/11 14:49 a-onl Start_Run_3360,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354781 04/08/11 14:48 a-onl End_of_Run_3359
  • 354780 04/08/11 13:19 a-onl Start_Run_3359,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354779 04/08/11 13:01 a-onl End_of_Run_3358
  • 354778 04/08/11 12:42 auto_snap Radiation and Ozone snapshot
  • 354777 04/08/11 12:42 auto_snap Miscellaneous snapshot
  • 354776 04/08/11 12:42 auto_snap Hall A Tools snapshot
  • 354775 04/08/11 12:42 auto_snap Compton Polarimeter snapshot
  • 354774 04/08/11 12:42 auto_snap BigBox Magnet snapshot
  • 354773 04/08/11 12:42 auto_snap Beam Position and Current snapshot
  • 354772 04/08/11 12:24 Annand HV trips HAND N5-3 R
  • 354771 04/08/11 12:01 Annand HV trip during run 3358
  • 354770 04/08/11 11:58 yez Target Screen
  • 354769 04/08/11 11:45 a-onl Start_Run_3358,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354768 04/08/11 11:43 a-onl End_of_Run_3357
  • 354767 04/08/11 10:26 a-onl Start_Run_3357,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354766 04/08/11 10:24 a-onl End_of_Run_3356
  • 354765 04/08/11 09:58 V. Sulkosky The Problem was BBTS1 not TS11
  • 354764 04/08/11 09:36 adaq N6 Position 3 Missing
  • 354763 04/08/11 09:23 BoZhao A few inefficient channels for N6-L
  • 354762 04/08/11 09:17 BoZhao A few inefficient channels for N6-L
  • 354761 04/08/11 09:10 dhamilton Reset Hand HV N6 left
  • 354760 04/08/11 09:07 a-onl Start_Run_3356,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354759 04/08/11 09:06 adaq a few channels for N6-L
  • 354758 04/08/11 09:05 a-onl End_of_Run_3355
  • 354757 04/08/11 08:40 a-onl Start_Run_3355,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354756 04/08/11 08:39 a-onl End_of_Run_3354
  • 354755 04/08/11 08:13 yez target screen
  • 354754 04/08/11 08:11 adaq Beam Time Accounting
  • 354753 04/08/11 07:56 nmuangma tool screen shot
  • 354752 04/08/11 07:48 nmuangma owl shift summary
  • 354751 04/08/11 07:24 a-onl Start_Run_3354,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354750 04/08/11 07:23 a-onl End_of_Run_3353
  • 354749 04/08/11 06:07 a-onl Start_Run_3353,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354748 04/08/11 06:01 nmuangma beam off: QE scan at MCC
  • 354747 04/08/11 06:01 a-onl End_of_Run_3352
  • 354746 04/08/11 04:42 auto_snap Radiation and Ozone snapshot
  • 354745 04/08/11 04:42 auto_snap Miscellaneous snapshot
  • 354744 04/08/11 04:42 auto_snap Hall A Tools snapshot
  • 354743 04/08/11 04:42 auto_snap Compton Polarimeter snapshot
  • 354742 04/08/11 04:42 auto_snap BigBox Magnet snapshot
  • 354741 04/08/11 04:42 auto_snap Beam Position and Current snapshot
  • 354740 04/08/11 04:39 a-onl Start_Run_3352,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354739 04/08/11 04:39 a-onl End_of_Run_3351
  • 354738 04/08/11 04:32 israely Target screenshot
  • 354737 04/08/11 03:40 orchen No data in the root tree -> There are histograms in the file.
  • 354736 04/08/11 03:22 a-onl Start_Run_3351,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354735 04/08/11 03:21 a-onl End_of_Run_3350
  • 354734 04/08/11 03:20 orchen Pion rejuctor problem was a DB change
  • 354733 04/08/11 03:06 yez Don't see PRL info in the SRC replayed data. Update PRL2 info in database from 20110222/~20110326/
  • 354732 04/08/11 02:48 yez PRLs work normal.
  • 354731 04/08/11 02:31 hkkang ND HV
  • 354730 04/08/11 01:56 a-onl Start_Run_3350,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354729 04/08/11 01:55 israely Hand left side is triped.
  • 354728 04/08/11 01:51 a-onl End_of_Run_3349
  • 354727 04/08/11 01:40 orchen CHANGE IN HRS PION REJECTOR?!?!?!
  • 354726 04/08/11 01:08 hkkang ND N6 ADC plot
  • 354725 04/08/11 01:06 a-onl Start_Run_3349,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354724 04/08/11 01:04 a-onl End_of_Run_3348
  • 354723 04/08/11 00:21 a-onl Start_Run_3348,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354721 04/08/11 00:03 adaq Beam Time Accounting
  • 354720 04/08/11 00:02 jpchen Swing shift summary
  • Thursday

  • 354719 04/07/11 23:56 a-onl Start_Run_3347,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354718 04/07/11 23:56 israely Target screenshot
  • 354717 04/07/11 23:47 a-onl End_of_Run_3346
  • 354716 04/07/11 23:31 jpchen DAQ probelm. Then beam off
  • 354715 04/07/11 23:27 Mitra CODA crashed several times.
  • 354714 04/07/11 23:22 a-onl Start_Run_3346,Run_type=Spot++,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Raster check,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354713 04/07/11 22:47 a-onl End_of_Run_3344
  • 354712 04/07/11 22:44 a-onl Start_Run_3344,Run_type=Spot++,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Raster check,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354711 04/07/11 22:19 a-onl Start_Run_3342,Run_type=Spot++,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Raster check,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354710 04/07/11 22:15 a-onl End_of_Run_3341
  • 354709 04/07/11 22:15 Nuruzzaman Target moved to He4.
  • 354708 04/07/11 22:05 a-onl Start_Run_3341,Run_type=Spot++,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Raster check,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354707 04/07/11 21:44 Nuruzzaman Target moved to BeO.
  • 354706 04/07/11 21:02 jpchen Beam status
  • 354705 04/07/11 20:41 auto_snap Radiation and Ozone snapshot
  • 354704 04/07/11 20:41 auto_snap Miscellaneous snapshot
  • 354703 04/07/11 20:41 auto_snap Hall A Tools snapshot
  • 354702 04/07/11 20:41 auto_snap Compton Polarimeter snapshot
  • 354701 04/07/11 20:41 auto_snap BigBox Magnet snapshot
  • 354700 04/07/11 20:41 auto_snap Beam Position and Current snapshot
  • 354699 04/07/11 20:12 Nuruzzaman Target moved to Empty.
  • 354698 04/07/11 20:12 a-onl End_of_Run_3340
  • 354697 04/07/11 19:48 jpchen Beam status
  • 354696 04/07/11 17:57 yez Set its HV back to orignal value -908V
  • 354695 04/07/11 17:23 yez Get the righ HV map of PRL and reduce HV of PRL2#14 from -908V to -850V
  • 354694 04/07/11 16:13 a-onl Start_Run_3340,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354693 04/07/11 16:12 a-onl End_of_Run_3339
  • 354692 04/07/11 16:03 adaq Beam Time Accounting
  • 354691 04/07/11 16:02 dhamilton Day shift summary
  • 354690 04/07/11 15:54 Fansler LQ2
  • 354688 04/07/11 15:41 a-onl Start_Run_3339,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354687 04/07/11 15:40 a-onl End_of_Run_3338
  • 354686 04/07/11 15:00 A. Camsonne Happex receiver surveyed and taken out from the Hall
  • 354685 04/07/11 14:33 a-onl Start_Run_3338,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354684 04/07/11 14:32 a-onl End_of_Run_3337
  • 354683 04/07/11 13:57 A. Camsonne Amplifier and attenuator in for HAPPEX BCM cavities parasitic test
  • 354682 04/07/11 12:41 auto_snap Radiation and Ozone snapshot
  • 354681 04/07/11 12:41 auto_snap Miscellaneous snapshot
  • 354680 04/07/11 12:41 auto_snap Hall A Tools snapshot
  • 354679 04/07/11 12:41 auto_snap Compton Polarimeter snapshot
  • 354678 04/07/11 12:41 auto_snap BigBox Magnet snapshot
  • 354677 04/07/11 12:41 auto_snap Beam Position and Current snapshot
  • 354676 04/07/11 10:27 yez L-HRS: Cerenkov data base updated /DB/20110326/
  • 354675 04/07/11 10:15 Solvignon RQ3 test
  • 354674 04/07/11 09:57 V. Sulkosky HAND L6 4th PMT
  • 354673 04/07/11 09:00 V. Sulkosky Triples
  • 354672 04/07/11 08:11 a-onl Start_Run_3337,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354671 04/07/11 08:11 adaq Beam Time Accounting
  • 354670 04/07/11 07:22 nmuangma shift summary
  • 354669 04/07/11 06:40 nmuangma beam off
  • 354668 04/07/11 06:34 a-onl End_of_Run_3336
  • 354667 04/07/11 06:23 a-onl Start_Run_3336,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354666 04/07/11 06:22 a-onl End_of_Run_3335
  • 354665 04/07/11 06:21 hkkang scintillator and scaler monitor
  • 354664 04/07/11 06:20 hkkang neutron detector
  • 354663 04/07/11 06:18 nmuangma hall A tool screen
  • 354662 04/07/11 06:01 nmuangma beam off at 6:30, 8-10 hrs to get the beam back
  • 354661 04/07/11 05:05 a-onl Start_Run_3335,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354660 04/07/11 05:04 a-onl End_of_Run_3334
  • 354659 04/07/11 04:41 auto_snap Radiation and Ozone snapshot
  • 354658 04/07/11 04:41 auto_snap Miscellaneous snapshot
  • 354657 04/07/11 04:41 auto_snap Hall A Tools snapshot
  • 354656 04/07/11 04:41 auto_snap Compton Polarimeter snapshot
  • 354655 04/07/11 04:41 auto_snap BigBox Magnet snapshot
  • 354654 04/07/11 04:41 auto_snap Beam Position and Current snapshot
  • 354653 04/07/11 03:54 israely Target screenshot
  • 354652 04/07/11 03:50 hkkang Neutron TDC
  • 354651 04/07/11 03:47 a-onl Start_Run_3334,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354650 04/07/11 03:46 a-onl End_of_Run_3333
  • 354649 04/07/11 02:45 a-onl Start_Run_3333,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354648 04/07/11 02:44 a-onl End_of_Run_3332
  • 354647 04/07/11 02:44 hkkang scintillator and scaler
  • 354646 04/07/11 02:42 hkkang screenshot for neutron detector
  • 354645 04/07/11 01:30 a-onl Start_Run_3332,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354644 04/07/11 01:29 a-onl End_of_Run_3331
  • 354643 04/07/11 01:28 hkkang scaler, scintillator
  • 354642 04/07/11 01:27 hkkang neutron detector
  • 354641 04/07/11 00:46 eip triple coinc events
  • 354640 04/07/11 00:10 a-onl Start_Run_3331,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354639 04/07/11 00:09 a-onl End_of_Run_3330
  • 354638 04/07/11 00:06 adaq Beam Time Accounting
  • 354637 04/07/11 00:05 Ole Hansen Disabled LHRS NMR nuisance alarm
  • Wednesday

  • 354636 04/06/11 23:58 Ole Hansen Swing Shift Summary
  • 354635 04/06/11 23:49 israely Target screenshot
  • 354634 04/06/11 23:05 Nuruzzaman Target Screens.
  • 354633 04/06/11 22:51 a-onl Start_Run_3330,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354632 04/06/11 22:50 a-onl End_of_Run_3329
  • 354631 04/06/11 21:52 yez It could be due to new pedestal values are not updated
  • 354630 04/06/11 21:34 Ole Hansen spot++ run 3329 - looks normal
  • 354629 04/06/11 21:30 a-onl Start_Run_3329,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354628 04/06/11 21:29 a-onl End_of_Run_3328
  • 354627 04/06/11 21:28 Ole Hansen spot++ run 3328 plots
  • 354626 04/06/11 21:27 Ole Hansen spot++ run 3328 events 50k-100k
  • 354625 04/06/11 21:22 orchen (e,e'N_recoil) from 2 and 8 uA runs in the new kinematics
  • 354624 04/06/11 21:08 Ole Hansen spot++ run 3328
  • 354623 04/06/11 20:44 a-onl Start_Run_3328,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354622 04/06/11 20:41 auto_snap Radiation and Ozone snapshot
  • 354621 04/06/11 20:40 auto_snap Miscellaneous snapshot
  • 354620 04/06/11 20:40 auto_snap Hall A Tools snapshot
  • 354619 04/06/11 20:40 auto_snap Compton Polarimeter snapshot
  • 354618 04/06/11 20:40 auto_snap BigBox Magnet snapshot
  • 354617 04/06/11 20:40 auto_snap Beam Position and Current snapshot
  • 354616 04/06/11 20:11 Nuruzzaman Target Screens.
  • 354615 04/06/11 19:33 Ole Hansen Beam off to work on position instability problem
  • 354614 04/06/11 19:32 a-onl End_of_Run_3327
  • 354613 04/06/11 19:11 a-onl Start_Run_3327,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354612 04/06/11 19:10 a-onl End_of_Run_3326
  • 354611 04/06/11 18:49 orchen BB (e,e'p_recoil) from verius runs in the 750MeV/c kinematics
  • 354610 04/06/11 18:34 orchen decrease(?) in HAND (e,e'n_recoil) from verius runs in the 750MeV/c kinematics
  • 354609 04/06/11 18:13 Carlos Cerekov replay plot
  • 354608 04/06/11 17:49 Ole Hansen Unstable beam position
  • 354607 04/06/11 17:45 a-onl Start_Run_3326,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354606 04/06/11 17:45 Nuruzzaman Target Screens.
  • 354605 04/06/11 17:44 a-onl End_of_Run_3325
  • 354604 04/06/11 16:34 Ole Hansen BB HV trip
  • 354603 04/06/11 16:29 orchen 4He(e,e'pN) , 1.52C , 750MeV/c kinematics
  • 354602 04/06/11 16:28 V. Sulkosky Re: bbts1 rebooted and power cycled
  • 354601 04/06/11 16:26 V. Sulkosky bbts1 rebooted and power cycled
  • 354600 04/06/11 16:20 a-onl Start_Run_3325,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354599 04/06/11 16:18 saw HAND scalers updated to account for EDTM pulses
  • 354598 04/06/11 16:14 a-onl End_of_Run_3324
  • 354597 04/06/11 16:03 adaq Beam Time Accounting
  • 354596 04/06/11 16:02 dhamilton Day Shift Summary
  • 354595 04/06/11 15:58 a-onl Start_Run_3324,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354594 04/06/11 15:58 Mitra S Target Screen
  • 354593 04/06/11 15:57 a-onl End_of_Run_3323
  • 354592 04/06/11 15:38 dhamilton Spot++ Screenshot
  • 354591 04/06/11 15:37 dhamilton Hall A Tools Screenshot
  • 354590 04/06/11 15:27 V. Sulkosky Sync roc plots for Run 3314
  • 354589 04/06/11 14:57 V. Sulkosky BigBite Normalization Scalers Runs 3312-3322
  • 354588 04/06/11 14:46 dhamilton TS power-cycled
  • 354587 04/06/11 14:44 lselvy SawDTMon back online
  • 354586 04/06/11 14:41 a-onl Start_Run_3323,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354585 04/06/11 14:38 a-onl End_of_Run_3322
  • 354584 04/06/11 13:20 a-onl Start_Run_3322,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354583 04/06/11 13:19 a-onl End_of_Run_3321
  • 354582 04/06/11 12:40 auto_snap Radiation and Ozone snapshot
  • 354581 04/06/11 12:40 auto_snap Miscellaneous snapshot
  • 354580 04/06/11 12:40 auto_snap Hall A Tools snapshot
  • 354579 04/06/11 12:40 auto_snap Compton Polarimeter snapshot
  • 354578 04/06/11 12:40 auto_snap BigBox Magnet snapshot
  • 354577 04/06/11 12:40 auto_snap Beam Position and Current snapshot
  • 354576 04/06/11 11:58 a-onl Start_Run_3321,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354575 04/06/11 11:57 a-onl End_of_Run_3320
  • 354574 04/06/11 11:48 Mitra S Target screens snap shots
  • 354573 04/06/11 10:40 a-onl Start_Run_3320,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354572 04/06/11 10:39 a-onl End_of_Run_3319
  • 354571 04/06/11 10:19 yez LHRS:Pion Rejectors Map look correct, so is the HV map wrong?
  • 354570 04/06/11 09:13 a-onl Start_Run_3319,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354569 04/06/11 09:09 a-onl End_of_Run_3317
  • 354568 04/06/11 08:02 adaq Beam Time Accounting
  • 354567 04/06/11 07:58 Carlos Owl shift summary
  • 354566 04/06/11 07:52 a-onl Start_Run_3317,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354565 04/06/11 07:51 a-onl End_of_Run_3316
  • 354564 04/06/11 07:51 israely Target screenshot
  • 354563 04/06/11 06:40 Carlos Hall A tools
  • 354562 04/06/11 06:36 Carlos spot++
  • 354561 04/06/11 06:31 Carlos Beam is back
  • 354560 04/06/11 06:31 a-onl Start_Run_3316,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354559 04/06/11 05:43 hkkang Again, ROC syncs
  • 354558 04/06/11 05:24 Carlos No beam for a while
  • 354557 04/06/11 05:17 a-onl End_of_Run_3315
  • 354556 04/06/11 04:52 igorko ROC sync
  • 354555 04/06/11 04:50 israely Target screenshot
  • 354554 04/06/11 04:40 auto_snap Radiation and Ozone snapshot
  • 354553 04/06/11 04:40 auto_snap Miscellaneous snapshot
  • 354552 04/06/11 04:40 auto_snap Hall A Tools snapshot
  • 354551 04/06/11 04:40 auto_snap Compton Polarimeter snapshot
  • 354550 04/06/11 04:40 auto_snap BigBox Magnet snapshot
  • 354549 04/06/11 04:40 auto_snap Beam Position and Current snapshot
  • 354548 04/06/11 04:33 hkkang ROC syncs in HRS
  • 354547 04/06/11 04:28 hkkang ROC syncs plot
  • 354546 04/06/11 04:01 a-onl Start_Run_3315,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354545 04/06/11 04:00 a-onl End_of_Run_3314
  • 354544 04/06/11 02:45 a-onl Start_Run_3314,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354543 04/06/11 02:44 a-onl End_of_Run_3313
  • 354542 04/06/11 02:10 israely Target screenshot
  • 354541 04/06/11 02:09 adaq Bigbite Scintillator screenshot
  • 354540 04/06/11 02:08 israely Neutron detector screenshot
  • 354539 04/06/11 01:57 nmuangma do we need to set t0 for L&R vdc?
  • 354538 04/06/11 01:44 hkkang L-HRS VDC
  • 354537 04/06/11 01:43 orchen L-HRS cerenkov sum
  • 354536 04/06/11 01:29 orchen Back to production
  • 354535 04/06/11 01:27 lselvy SawDTMon Offline
  • 354534 04/06/11 01:26 a-onl Start_Run_3313,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354533 04/06/11 01:10 a-onl End_of_Run_3312
  • 354532 04/06/11 01:08 lselvy iocha22 and iocha11 reboot
  • 354531 04/06/11 01:02 lselvy ROC 5 Reset Problem
  • 354530 04/06/11 01:02 a-onl Start_Run_3312,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354529 04/06/11 01:01 a-onl End_of_Run_3310
  • 354528 04/06/11 00:58 Carlos Half-wave plate IN
  • 354527 04/06/11 00:42 a-onl End_of_Run_3310
  • 354526 04/06/11 00:39 a-onl Start_Run_3310,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354525 04/06/11 00:07 adaq Beam Time Accounting
  • 354524 04/06/11 00:05 Ramesh Shift Summary
  • 354523 04/06/11 00:00 Ramesh Snapshot of Hall A Tools with problems, also with IOCs
  • Tuesday

  • 354522 04/05/11 23:58 a-onl Start_Run_3309,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354521 04/05/11 23:53 Ramesh Asked for controlled access to reset LHRS HV
  • 354520 04/05/11 23:51 a-onl Start_Run_3308,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354519 04/05/11 23:45 a-onl Start_Run_3307,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354518 04/05/11 23:41 a-onl Start_Run_3306,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354517 04/05/11 23:36 Nuruzzaman Target Screens at end of swing shift.
  • 354516 04/05/11 23:28 Ramesh VDC and other High Voltages lost
  • 354515 04/05/11 23:19 a-onl Start_Run_3305,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354514 04/05/11 23:14 a-onl Start_Run_3304,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354513 04/05/11 22:56 a-onl Start_Run_3303,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354512 04/05/11 22:52 a-onl Start_Run_3302,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354511 04/05/11 22:46 a-onl Start_Run_3301,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354510 04/05/11 22:46 orchen BB current timing resolution
  • 354509 04/05/11 22:45 a-onl End_of_Run_3300
  • 354508 04/05/11 21:49 Nuruzzaman Target Screens.
  • 354507 04/05/11 21:29 a-onl Start_Run_3300,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354506 04/05/11 21:29 a-onl End_of_Run_3299
  • 354505 04/05/11 20:40 auto_snap Radiation and Ozone snapshot
  • 354504 04/05/11 20:40 auto_snap Miscellaneous snapshot
  • 354503 04/05/11 20:39 auto_snap Hall A Tools snapshot
  • 354502 04/05/11 20:39 auto_snap Compton Polarimeter snapshot
  • 354501 04/05/11 20:39 auto_snap BigBox Magnet snapshot
  • 354500 04/05/11 20:39 auto_snap Beam Position and Current snapshot
  • 354499 04/05/11 20:14 a-onl Start_Run_3299,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354498 04/05/11 20:13 a-onl End_of_Run_3298
  • 354497 04/05/11 20:07 V. Sulkosky Added Pathlength Coefficients to Right VDC DB
  • 354496 04/05/11 18:58 a-onl Start_Run_3298,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354495 04/05/11 18:57 a-onl End_of_Run_3297
  • 354494 04/05/11 18:44 Nuruzzaman Target Screens.
  • 354493 04/05/11 18:35 V. Sulkosky Vertex Variables Now Working
  • 354492 04/05/11 17:29 a-onl Start_Run_3297,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354491 04/05/11 17:28 a-onl End_of_Run_3296
  • 354490 04/05/11 17:13 Ramesh Production started
  • 354489 04/05/11 17:06 a-onl Start_Run_3296,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354488 04/05/11 17:03 a-onl End_of_Run_3295
  • 354487 04/05/11 17:01 a-onl Start_Run_3295,Run_type=Check out,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=1,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Beam recovery,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354486 04/05/11 16:57 Nuruzzaman Target Moved to He4.
  • 354485 04/05/11 16:50 a-onl End_of_Run_3294
  • 354484 04/05/11 16:49 Ramesh spot check with BeO target when in beam recovery procedure
  • 354483 04/05/11 16:47 a-onl Start_Run_3294,Run_type=Check out,target_type=BeO,beam_e0=4.46,beam_raster=on,beam_current=5,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354482 04/05/11 16:45 a-onl End_of_Run_3293
  • 354481 04/05/11 16:35 Nuruzzaman Target moved to BeO.
  • 354480 04/05/11 16:26 Jack (Techs) RQ1 current regulation issue
  • 354479 04/05/11 16:23 saw Funky HAND scaler GUI readings
  • 354478 04/05/11 16:21 Ramesh Hall in beam permit
  • 354477 04/05/11 16:17 A. Camsonne I meant DAC instead of DAQ in the title
  • 354476 04/05/11 16:08 A. Camsonne Upstream Happex Cavity receiver DAQ dead
  • 354475 04/05/11 16:02 adaq Beam Time Accounting
  • 354474 04/05/11 15:58 dhamilton Day shift summary
  • 354473 04/05/11 15:48 B Anderson Target screens. Beam still off.
  • 354472 04/05/11 15:28 dhamilton Low pressure alarm on Detector Gas System
  • 354471 04/05/11 14:48 a-onl Start_Run_3293,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354470 04/05/11 14:47 a-onl End_of_Run_3292
  • 354469 04/05/11 13:50 B Anderson Target screens. Beam still off.
  • 354468 04/05/11 13:06 a-onl Start_Run_3292,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354467 04/05/11 13:04 a-onl End_of_Run_3291
  • 354466 04/05/11 12:39 auto_snap Radiation and Ozone snapshot
  • 354465 04/05/11 12:39 auto_snap Miscellaneous snapshot
  • 354464 04/05/11 12:39 auto_snap Hall A Tools snapshot
  • 354463 04/05/11 12:39 auto_snap Compton Polarimeter snapshot
  • 354462 04/05/11 12:39 auto_snap BigBox Magnet snapshot
  • 354461 04/05/11 12:39 auto_snap Beam Position and Current snapshot
  • 354460 04/05/11 11:50 a-onl Start_Run_3291,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354459 04/05/11 11:49 a-onl End_of_Run_3290
  • 354458 04/05/11 11:26 yez HV which I turned of now is on and its value is put back to -1020
  • 354457 04/05/11 11:26 V. Sulkosky Hall A Visual Inspection
  • 354456 04/05/11 11:24 yez PRL2 HV map is wrong (or data map)
  • 354455 04/05/11 11:00 B Anderson Target screens. Beam still off.
  • 354454 04/05/11 09:53 yez Issue on Pion Rejector 2#14: ADC saturates before showing a shap peak of overfload
  • 354453 04/05/11 09:42 a-onl Start_Run_3290,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Cosmics,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354452 04/05/11 09:40 yez Turn off HV of PRL2#14
  • 354451 04/05/11 08:59 dhamilton Hall status moving to controlled access
  • 354450 04/05/11 08:53 B Anderson Target screens. Beam off.
  • 354449 04/05/11 08:41 dhamilton Lost beam due to arc dipole problem
  • 354448 04/05/11 08:38 a-onl End_of_Run_3289
  • 354447 04/05/11 08:03 adaq Beam Time Accounting
  • 354446 04/05/11 07:55 Carlos Munoz Owl shift summary
  • 354445 04/05/11 07:32 a-onl Start_Run_3289,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354444 04/05/11 07:30 a-onl End_of_Run_3288
  • 354443 04/05/11 07:28 Carlos Hall A tools
  • 354442 04/05/11 07:27 Carlos spot++
  • 354441 04/05/11 07:13 Hisham Target status
  • 354440 04/05/11 06:16 a-onl Start_Run_3288,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354439 04/05/11 06:15 a-onl End_of_Run_3287
  • 354438 04/05/11 04:57 Carlos DAQ hung
  • 354437 04/05/11 04:55 a-onl Start_Run_3287,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354436 04/05/11 04:43 a-onl Start_Run_3286,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354435 04/05/11 04:42 a-onl End_of_Run_3285
  • 354434 04/05/11 04:39 auto_snap Radiation and Ozone snapshot
  • 354433 04/05/11 04:39 auto_snap Miscellaneous snapshot
  • 354432 04/05/11 04:39 auto_snap Hall A Tools snapshot
  • 354431 04/05/11 04:39 auto_snap Compton Polarimeter snapshot
  • 354430 04/05/11 04:39 auto_snap BigBox Magnet snapshot
  • 354429 04/05/11 04:39 auto_snap Beam Position and Current snapshot
  • 354428 04/05/11 04:36 Hisham Target status
  • 354427 04/05/11 03:25 a-onl Start_Run_3285,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354426 04/05/11 03:24 a-onl End_of_Run_3284
  • 354425 04/05/11 02:07 a-onl Start_Run_3284,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354424 04/05/11 02:06 Hisham Target status
  • 354423 04/05/11 02:06 a-onl End_of_Run_3283
  • 354422 04/05/11 00:52 a-onl Start_Run_3283,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354421 04/05/11 00:51 a-onl End_of_Run_3281
  • 354420 04/05/11 00:50 a-onl End_of_Run_3282
  • 354419 04/05/11 00:02 adaq Beam Time Accounting
  • Monday

  • 354418 04/04/11 23:59 Ramesh Shift Summary
  • 354417 04/04/11 23:57 Nuruzzaman Target Screens at end of swing shift.
  • 354416 04/04/11 23:34 a-onl Start_Run_3282,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354415 04/04/11 23:14 Ramesh spot 2x2
  • 354414 04/04/11 22:55 a-onl End_of_Run_3280
  • 354413 04/04/11 22:21 Ramesh Re: HV reduction on HAND noisy bar, N6-4R
  • 354412 04/04/11 22:20 a-onl Start_Run_3281,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354411 04/04/11 22:04 Nuruzzaman Target Screens.
  • 354410 04/04/11 22:01 Norum N6 Plane Right Raw TDC 3
  • 354409 04/04/11 21:15 orchen HV reduction on HAND noisy bar, N6-4R
  • 354408 04/04/11 21:05 a-onl Start_Run_3280,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354407 04/04/11 21:04 a-onl End_of_Run_3279
  • 354406 04/04/11 20:39 auto_snap Radiation and Ozone snapshot
  • 354405 04/04/11 20:39 auto_snap Miscellaneous snapshot
  • 354404 04/04/11 20:38 auto_snap Hall A Tools snapshot
  • 354403 04/04/11 20:38 auto_snap Compton Polarimeter snapshot
  • 354402 04/04/11 20:38 auto_snap BigBox Magnet snapshot
  • 354401 04/04/11 20:38 auto_snap Beam Position and Current snapshot
  • 354400 04/04/11 20:31 orchen 4He(e,e'pN) , 1.156C , 750MeV/c kinematics
  • 354399 04/04/11 20:03 Nuruzzaman Target Screens.
  • 354398 04/04/11 19:55 Norum ND N6 TDC's noisy
  • 354397 04/04/11 19:27 Norum ND N6 TDC's noisy
  • 354396 04/04/11 19:25 a-onl Start_Run_3279,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354395 04/04/11 19:22 a-onl End_of_Run_3278
  • 354394 04/04/11 18:09 Ramesh gas shed alarms (two red lights for about 2 minutes)
  • 354393 04/04/11 18:09 Norum ND N6 TDC's noisy
  • 354392 04/04/11 18:05 Nuruzzaman Target Screens.
  • 354391 04/04/11 18:02 a-onl Start_Run_3278,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354390 04/04/11 17:58 a-onl End_of_Run_3277
  • 354389 04/04/11 17:10 a-onl Start_Run_3277,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354388 04/04/11 17:08 a-onl End_of_Run_3276
  • 354387 04/04/11 17:07 a-onl Start_Run_3276,Run_type=Other,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=test run for HAND scalers,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354386 04/04/11 17:04 a-onl End_of_Run_3275
  • 354385 04/04/11 16:50 a-onl Start_Run_3275,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354384 04/04/11 16:49 a-onl End_of_Run_3274
  • 354383 04/04/11 16:07 Nuruzzaman Target Screens, begining of swing shift.
  • 354382 04/04/11 16:03 adaq Beam Time Accounting
  • 354381 04/04/11 15:56 dhamilton Day shift summary
  • 354380 04/04/11 15:28 B Anderson Target screens
  • 354379 04/04/11 15:27 dhamilton Beam restored
  • 354378 04/04/11 15:25 a-onl Start_Run_3274,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354377 04/04/11 15:12 dhamilton Beam away for injector tests
  • 354376 04/04/11 15:12 dhamilton Spot++
  • 354375 04/04/11 15:10 dhamilton Hall A Tools
  • 354374 04/04/11 15:05 a-onl End_of_Run_3273
  • 354373 04/04/11 14:48 a-onl Start_Run_3273,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354372 04/04/11 14:46 a-onl End_of_Run_3272
  • 354371 04/04/11 13:47 B Anderson Target screens
  • 354370 04/04/11 13:24 a-onl Start_Run_3272,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354369 04/04/11 13:23 a-onl End_of_Run_3271
  • 354368 04/04/11 12:47 dhamilton Gas system interlock status alarm
  • 354367 04/04/11 12:38 auto_snap Radiation and Ozone snapshot
  • 354366 04/04/11 12:38 auto_snap Miscellaneous snapshot
  • 354365 04/04/11 12:38 auto_snap Hall A Tools snapshot
  • 354364 04/04/11 12:38 auto_snap Compton Polarimeter snapshot
  • 354363 04/04/11 12:38 auto_snap BigBox Magnet snapshot
  • 354362 04/04/11 12:38 auto_snap Beam Position and Current snapshot
  • 354361 04/04/11 12:07 a-onl Start_Run_3271,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354360 04/04/11 12:06 a-onl End_of_Run_3270
  • 354359 04/04/11 11:48 B Anderson Target screens
  • 354358 04/04/11 10:47 a-onl Start_Run_3270,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354357 04/04/11 10:46 a-onl End_of_Run_3269
  • 354356 04/04/11 09:46 B Anderson Target screens
  • 354355 04/04/11 09:43 JohanS Big Hand N6 left qdc
  • 354354 04/04/11 09:35 dhamilton BigHand layer 6 left HV power-cycle
  • 354353 04/04/11 09:26 a-onl Start_Run_3269,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354352 04/04/11 09:24 a-onl End_of_Run_3268
  • 354351 04/04/11 08:45 a-onl Start_Run_3268,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354350 04/04/11 08:45 a-onl End_of_Run_3267
  • 354349 04/04/11 08:38 dhamilton Gas Alarm Reset
  • 354348 04/04/11 08:16 dhamilton BB HV alarm
  • 354347 04/04/11 08:03 adaq Beam Time Accounting
  • 354346 04/04/11 07:55 Carlos Munoz Owl shift summary
  • 354345 04/04/11 07:46 Alejandro Ma N6 Left screenshot
  • 354344 04/04/11 07:43 israely Target screenshot
  • 354343 04/04/11 07:25 Carlos Hall A tools
  • 354342 04/04/11 07:24 a-onl Start_Run_3267,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354341 04/04/11 07:24 Carlos spot++
  • 354340 04/04/11 07:24 a-onl End_of_Run_3266
  • 354339 04/04/11 07:10 israely The was alarm in the following channels:
  • 354338 04/04/11 06:05 israely Target screenshot
  • 354337 04/04/11 06:05 a-onl Start_Run_3266,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354336 04/04/11 06:04 a-onl End_of_Run_3265
  • 354335 04/04/11 04:50 a-onl Start_Run_3265,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354334 04/04/11 04:49 a-onl End_of_Run_3264
  • 354333 04/04/11 04:38 auto_snap Radiation and Ozone snapshot
  • 354332 04/04/11 04:38 auto_snap Miscellaneous snapshot
  • 354331 04/04/11 04:38 auto_snap Hall A Tools snapshot
  • 354330 04/04/11 04:38 auto_snap Compton Polarimeter snapshot
  • 354329 04/04/11 04:38 auto_snap BigBox Magnet snapshot
  • 354328 04/04/11 04:38 auto_snap Beam Position and Current snapshot
  • 354327 04/04/11 03:35 a-onl Start_Run_3264,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354326 04/04/11 03:33 a-onl End_of_Run_3263
  • 354325 04/04/11 02:54 alejandro ma 3261 replay screenshot
  • 354324 04/04/11 02:50 israely Igor - I tried to reset
  • 354323 04/04/11 02:45 alejandro ma neutron detector screen
  • 354322 04/04/11 02:44 israely Bigbite screenshot
  • 354321 04/04/11 02:41 israely Target screenshot
  • 354320 04/04/11 02:31 igorko plane 6 bar 3
  • 354319 04/04/11 02:30 a-onl Start_Run_3263,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354318 04/04/11 02:16 a-onl End_of_Run_3262
  • 354317 04/04/11 01:01 a-onl Start_Run_3262,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354316 04/04/11 00:58 a-onl End_of_Run_3261
  • 354315 04/04/11 00:11 Alejandro Ma 3261 Replay
  • 354314 04/04/11 00:01 adaq Beam Time Accounting
  • Sunday

  • 354313 04/03/11 23:56 israely Target screenshot
  • 354312 04/03/11 23:55 Ramesh Shift Summary
  • 354311 04/03/11 23:53 Sarah Swing shift run list (analysis)
  • 354310 04/03/11 23:48 Sarah Screen capture of the neutron detector single rates during run 3261
  • 354309 04/03/11 23:47 Sarah Screen captures of the bigbite and scaler data screens during run 3261
  • 354308 04/03/11 23:41 a-onl Start_Run_3261,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354307 04/03/11 23:40 a-onl End_of_Run_3260
  • 354306 04/03/11 22:39 Sarah Screen capture of the neutron detector single rates during run 3260
  • 354305 04/03/11 22:38 Sarah Screen captures for the bigbite and scaler data screens during run 3260
  • 354304 04/03/11 22:29 Ramesh Hall A general tools screen
  • 354303 04/03/11 22:26 Ramesh spot check for shift checklist
  • 354302 04/03/11 22:24 a-onl Start_Run_3260,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354301 04/03/11 22:24 a-onl End_of_Run_3259
  • 354300 04/03/11 22:13 LeRose Target Screens
  • 354299 04/03/11 22:00 Sarah Screen capture of the neutron detector single rates during run 3259
  • 354298 04/03/11 21:59 Sarah Screen captures of the bigbite and the scaler data screens during run 3259
  • 354297 04/03/11 21:08 a-onl Start_Run_3259,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354296 04/03/11 21:07 a-onl End_of_Run_3258
  • 354295 04/03/11 20:38 auto_snap Radiation and Ozone snapshot
  • 354294 04/03/11 20:38 auto_snap Miscellaneous snapshot
  • 354293 04/03/11 20:38 auto_snap Hall A Tools snapshot
  • 354292 04/03/11 20:37 auto_snap Compton Polarimeter snapshot
  • 354291 04/03/11 20:37 auto_snap BigBox Magnet snapshot
  • 354290 04/03/11 20:37 auto_snap Beam Position and Current snapshot
  • 354289 04/03/11 20:24 Sarah Screen capture of neutron detector single rates during run 3258
  • 354288 04/03/11 20:21 Sarah Screen captures of bigbite and scaler data screens during run 3258
  • 354287 04/03/11 19:54 LeRose Target Screens
  • 354286 04/03/11 19:52 a-onl Start_Run_3258,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354285 04/03/11 19:51 a-onl End_of_Run_3257
  • 354284 04/03/11 19:17 Sarah Screen capture of the neutron detector single rates during run 3257
  • 354283 04/03/11 19:16 Sarah Screen captures of the bigbite and scaler data screens during run 3257
  • 354282 04/03/11 18:28 a-onl Start_Run_3257,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354281 04/03/11 18:27 a-onl End_of_Run_3256
  • 354280 04/03/11 17:45 Sarah Screen capture of the neutron detector single rates during run 3256
  • 354279 04/03/11 17:44 Sarah Screen captures for the bigbite and scaler screens during run 3256
  • 354278 04/03/11 17:40 LeRose Target Screens
  • 354277 04/03/11 17:13 a-onl Start_Run_3256,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354276 04/03/11 17:12 a-onl End_of_Run_3255
  • 354275 04/03/11 16:33 Sarah Screen capture of the neutron detector single rates during run 3255
  • 354274 04/03/11 16:32 Sarah Screen captures of the bigbite and scaler data screens during run 3255
  • 354273 04/03/11 16:04 adaq Beam Time Accounting
  • 354272 04/03/11 15:55 Carlos Munoz Day shift summary
  • 354271 04/03/11 15:54 a-onl Start_Run_3255,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354270 04/03/11 15:47 Carlos Transmission problems
  • 354269 04/03/11 15:46 a-onl End_of_Run_3254
  • 354268 04/03/11 15:22 Carlos Hall A tools
  • 354267 04/03/11 15:20 Carlos spot++
  • 354266 04/03/11 15:16 B Anderson Target screens. Beam at 4uA.
  • 354265 04/03/11 15:04 Carlos Flammable gas alarm
  • 354264 04/03/11 14:55 Carlos Coda restarted
  • 354263 04/03/11 14:54 a-onl Start_Run_3254,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354262 04/03/11 14:48 Carlos Back to 4uA
  • 354261 04/03/11 14:48 a-onl Start_Run_3253,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354260 04/03/11 14:47 a-onl End_of_Run_3252
  • 354259 04/03/11 14:02 B Anderson Target screens. Beam at 8uA
  • 354258 04/03/11 13:50 Carlos 8 uA run
  • 354257 04/03/11 13:49 a-onl Start_Run_3252,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=8,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354256 04/03/11 13:48 a-onl End_of_Run_3251
  • 354255 04/03/11 12:57 B Anderson Target screens
  • 354254 04/03/11 12:37 auto_snap Radiation and Ozone snapshot
  • 354253 04/03/11 12:37 auto_snap Miscellaneous snapshot
  • 354252 04/03/11 12:37 auto_snap Hall A Tools snapshot
  • 354251 04/03/11 12:37 auto_snap Compton Polarimeter snapshot
  • 354250 04/03/11 12:37 auto_snap BigBox Magnet snapshot
  • 354249 04/03/11 12:37 auto_snap Beam Position and Current snapshot
  • 354248 04/03/11 12:30 Carlos 2 uA run
  • 354247 04/03/11 12:28 a-onl Start_Run_3251,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=2,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354246 04/03/11 12:26 a-onl End_of_Run_3250
  • 354245 04/03/11 11:32 orchen HAND adc bars problem fixed
  • 354244 04/03/11 11:16 Carlos HV powercycled
  • 354243 04/03/11 11:12 a-onl Start_Run_3250,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354242 04/03/11 11:11 a-onl End_of_Run_3249
  • 354241 04/03/11 10:55 B Anderson Target screens; beam back on
  • 354240 04/03/11 10:50 Carlos HV trip (2001 S13 L10.2)
  • 354239 04/03/11 10:47 Carlos Beam is back
  • 354238 04/03/11 10:46 a-onl Start_Run_3249,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354237 04/03/11 10:45 a-onl End_of_Run_3248
  • 354236 04/03/11 10:40 a-onl Start_Run_3248,Run_type=Spot++,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354235 04/03/11 10:10 B Anderson Target screens; beam off
  • 354234 04/03/11 10:02 Carlos Munoz Injector work
  • 354233 04/03/11 10:01 a-onl End_of_Run_3247
  • 354232 04/03/11 09:41 James Bittne Replay of run 3247
  • 354231 04/03/11 09:14 Carlos Munoz Half-wave plate is now OUT
  • 354230 04/03/11 09:13 a-onl Start_Run_3247,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354229 04/03/11 09:12 a-onl End_of_Run_3246
  • 354228 04/03/11 09:06 B Anderson Target screens
  • 354227 04/03/11 08:05 adaq Beam Time Accounting
  • 354226 04/03/11 08:00 D.Wang owl shift summary
  • 354225 04/03/11 07:55 a-onl Start_Run_3246,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354224 04/03/11 07:54 a-onl End_of_Run_3245
  • 354223 04/03/11 06:39 a-onl Start_Run_3245,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354222 04/03/11 06:38 a-onl End_of_Run_3244
  • 354221 04/03/11 05:20 a-onl Start_Run_3244,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354220 04/03/11 05:19 a-onl End_of_Run_3243
  • 354219 04/03/11 04:37 auto_snap Radiation and Ozone snapshot
  • 354218 04/03/11 04:37 auto_snap Miscellaneous snapshot
  • 354217 04/03/11 04:37 auto_snap Hall A Tools snapshot
  • 354216 04/03/11 04:37 auto_snap Compton Polarimeter snapshot
  • 354215 04/03/11 04:37 auto_snap BigBox Magnet snapshot
  • 354214 04/03/11 04:37 auto_snap Beam Position and Current snapshot
  • 354213 04/03/11 04:03 a-onl Start_Run_3243,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354212 04/03/11 04:01 a-onl End_of_Run_3242
  • 354211 04/03/11 02:43 a-onl Start_Run_3242,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354210 04/03/11 02:42 a-onl End_of_Run_3241
  • 354209 04/03/11 01:28 a-onl Start_Run_3241,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354208 04/03/11 01:27 a-onl End_of_Run_3240
  • 354207 04/03/11 00:12 a-onl Start_Run_3240,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354206 04/03/11 00:12 a-onl End_of_Run_3239
  • 354205 04/03/11 00:07 wenbiao target status
  • 354204 04/03/11 00:02 adaq Beam Time Accounting
  • Saturday

  • 354203 04/02/11 23:53 Ramesh Shift Summary
  • 354202 04/02/11 23:48 Sarah Swing shift run list (analysis)
  • 354201 04/02/11 23:47 Mitra Target screens
  • 354200 04/02/11 22:58 a-onl Start_Run_3239,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354199 04/02/11 22:55 a-onl End_of_Run_3238
  • 354198 04/02/11 22:21 Ramesh General Tools snapshot for the shift checklist
  • 354197 04/02/11 22:19 adaq spot check for the shift checklist
  • 354196 04/02/11 22:17 Mitra S Target screens after beam return
  • 354195 04/02/11 22:15 Sarah Screen capture of the neutron detector single rates during run 3238
  • 354194 04/02/11 22:14 Sarah Screen capture of the bigbite and scaler data screens during run 3238
  • 354192 04/02/11 22:06 V. Sulkosky Right Dipole Regulator Should be off
  • 354191 04/02/11 22:02 Sarah Odd histogram in Run 3238
  • 354190 04/02/11 21:50 orchen HAND (e.e'pn) for 0.6C from 625MeV/c Kinematics
  • 354189 04/02/11 21:45 Ramesh Left dipole hall probe alarm is beam related
  • 354188 04/02/11 21:42 orchen HAND (e.e'pn) for 0.6C from 750MeV/c Kinematics
  • 354187 04/02/11 21:38 Sarah HACWEB1 rebooted
  • 354186 04/02/11 21:36 Sarah This run is mostly cosmics
  • 354185 04/02/11 21:35 Ramesh beam back
  • 354184 04/02/11 21:35 a-onl Start_Run_3238,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354183 04/02/11 21:34 a-onl End_of_Run_3237
  • 354182 04/02/11 21:19 a-onl Start_Run_3237,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Cosmics, No beam!,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354181 04/02/11 21:06 a-onl End_of_Run_3236
  • 354180 04/02/11 21:03 Ramesh no beam status update
  • 354179 04/02/11 20:37 auto_snap Radiation and Ozone snapshot
  • 354178 04/02/11 20:37 auto_snap Miscellaneous snapshot
  • 354177 04/02/11 20:37 auto_snap Hall A Tools snapshot
  • 354176 04/02/11 20:37 auto_snap Compton Polarimeter snapshot
  • 354175 04/02/11 20:37 auto_snap BigBox Magnet snapshot
  • 354174 04/02/11 20:36 auto_snap Beam Position and Current snapshot
  • 354173 04/02/11 19:01 Sarah HACWEB1 is locked up
  • 354172 04/02/11 18:48 Ramesh beam down since 18:26 again due to an injector problem again
  • 354171 04/02/11 18:17 a-onl Start_Run_3236,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354170 04/02/11 18:16 Ramesh beam back
  • 354169 04/02/11 18:16 a-onl End_of_Run_3235
  • 354168 04/02/11 18:10 a-onl Start_Run_3235,Run_type=Cosmics,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=0,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Cosmics, no beam,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354167 04/02/11 18:08 Ramesh Beam down for a while due to the injector problem
  • 354166 04/02/11 18:06 a-onl End_of_Run_3234
  • 354165 04/02/11 17:40 Ramesh Right HRS dipole regulator is OFF, but why?
  • 354164 04/02/11 17:30 Sarah Screen capture of neutron detector single rates screen during run 3234
  • 354163 04/02/11 17:29 Sarah Screen capture of bigbite and scaler data screens during run 3234
  • 354162 04/02/11 17:23 Ramesh Spot move on Sunday 10-11 AM
  • 354161 04/02/11 17:23 a-onl Start_Run_3234,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354160 04/02/11 17:22 a-onl End_of_Run_3233
  • 354159 04/02/11 16:52 orchen (e,e'pp) and (e,e'p) for 0.55C in the 6250MeV/c kinematics
  • 354158 04/02/11 16:46 Sarah Screen snapshot of neutron detector single rates during run 3233
  • 354157 04/02/11 16:45 Sarah Screen snapshots of bigbite and scaler data screens during run 3233
  • 354156 04/02/11 16:24 Sarah Weird histogram in run 3232
  • 354155 04/02/11 16:18 orchen (e,e'pp) and (e,e'p) for 0.55C in the 750MeV/c kinematics
  • 354154 04/02/11 16:09 a-onl Start_Run_3233,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354153 04/02/11 16:08 a-onl End_of_Run_3232
  • 354152 04/02/11 16:06 Mitra S Target screens
  • 354151 04/02/11 16:03 adaq Beam Time Accounting
  • 354150 04/02/11 15:55 C. Gu Day Shift Summary
  • 354149 04/02/11 15:24 dalton Flamable gas alarm
  • 354148 04/02/11 14:51 a-onl Start_Run_3232,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354147 04/02/11 14:50 a-onl End_of_Run_3231
  • 354146 04/02/11 13:40 C. Gu Spot++
  • 354145 04/02/11 13:34 a-onl Start_Run_3231,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354144 04/02/11 13:33 a-onl End_of_Run_3230
  • 354143 04/02/11 12:51 C. Gu Tools screen saved
  • 354142 04/02/11 12:36 auto_snap Radiation and Ozone snapshot
  • 354141 04/02/11 12:36 auto_snap Miscellaneous snapshot
  • 354140 04/02/11 12:36 auto_snap Hall A Tools snapshot
  • 354139 04/02/11 12:36 auto_snap Compton Polarimeter snapshot
  • 354138 04/02/11 12:36 auto_snap BigBox Magnet snapshot
  • 354137 04/02/11 12:36 auto_snap Beam Position and Current snapshot
  • 354136 04/02/11 12:19 a-onl Start_Run_3230,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354135 04/02/11 12:18 a-onl End_of_Run_3229
  • 354134 04/02/11 12:12 C. Gu Gas Shed Airflow Alarm
  • 354133 04/02/11 10:56 a-onl Start_Run_3229,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354132 04/02/11 10:56 a-onl End_of_Run_3228
  • 354131 04/02/11 09:32 a-onl Start_Run_3228,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354130 04/02/11 09:31 a-onl End_of_Run_3227
  • 354129 04/02/11 08:17 arie correction for owl shift summary - normal format
  • 354128 04/02/11 08:14 a-onl Start_Run_3227,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354127 04/02/11 08:13 a-onl End_of_Run_3226
  • 354126 04/02/11 08:04 adaq Beam Time Accounting
  • 354125 04/02/11 08:04 wenbiao target status
  • 354124 04/02/11 08:04 arie shift summary
  • 354123 04/02/11 07:30 D.Wang target status
  • 354122 04/02/11 06:55 a-onl Start_Run_3226,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354121 04/02/11 06:53 a-onl End_of_Run_3225
  • 354120 04/02/11 06:02 a-onl Start_Run_3225,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354119 04/02/11 05:24 a-onl End_of_Run_3224
  • 354118 04/02/11 05:19 D.Wang target status
  • 354117 04/02/11 04:36 auto_snap Radiation and Ozone snapshot
  • 354116 04/02/11 04:36 auto_snap Miscellaneous snapshot
  • 354115 04/02/11 04:36 auto_snap Hall A Tools snapshot
  • 354114 04/02/11 04:36 auto_snap Compton Polarimeter snapshot
  • 354113 04/02/11 04:36 auto_snap BigBox Magnet snapshot
  • 354112 04/02/11 04:36 auto_snap Beam Position and Current snapshot
  • 354111 04/02/11 04:14 arie general tools
  • 354110 04/02/11 04:10 arie spot++
  • 354109 04/02/11 04:07 a-onl Start_Run_3224,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354108 04/02/11 04:06 a-onl End_of_Run_3223
  • 354107 04/02/11 03:15 D.Wang target status
  • 354106 04/02/11 02:47 a-onl Start_Run_3223,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354105 04/02/11 02:47 a-onl End_of_Run_3222
  • 354104 04/02/11 02:14 israely Screenshots of neutron detector single rate
  • 354103 04/02/11 02:13 israely Bigbite scintillator screenshot
  • 354102 04/02/11 01:32 a-onl Start_Run_3222,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354101 04/02/11 01:31 a-onl End_of_Run_3221
  • 354100 04/02/11 01:17 orchen HAND (e,e'pn) BG for 750MeV/c and 625MeV/c kinematics
  • 354099 04/02/11 00:59 orchen BB triple coincedence 750MeV/c vs. 625MeV/c kinematics
  • 354098 04/02/11 00:33 adaq Beam Time Accounting
  • 354097 04/02/11 00:15 a-onl Start_Run_3221,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354096 04/02/11 00:12 a-onl End_of_Run_3220
  • 354095 04/02/11 00:09 a-onl Start_Run_3220,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354094 04/02/11 00:08 a-onl End_of_Run_3219
  • Friday

  • 354093 04/01/11 23:56 Fatiha Shift Summary
  • 354092 04/01/11 23:54 Sarah Swing shift run list (analysis)
  • 354091 04/01/11 23:43 zwzhao target screen
  • 354090 04/01/11 23:16 Sarah Screen capture of the neutron detector single rates during run 3219
  • 354089 04/01/11 23:15 Sarah Screen capture of bigbite and scaler data screens during run 3219
  • 354088 04/01/11 23:09 Sarah HV trip
  • 354087 04/01/11 22:42 a-onl Start_Run_3219,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354086 04/01/11 22:41 a-onl End_of_Run_3218
  • 354085 04/01/11 22:15 Fatiha Beam current
  • 354084 04/01/11 22:02 Sarah Screen capture of the neutron detector single rates during run 3218
  • 354083 04/01/11 22:00 Sarah Screen capture of bigbite and scaler data screens during run 3218
  • 354082 04/01/11 21:39 Fatiha Spot++
  • 354081 04/01/11 21:27 Sarah CODA restarted
  • 354080 04/01/11 21:26 a-onl Start_Run_3218,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354079 04/01/11 21:22 a-onl Start_Run_3217,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354078 04/01/11 21:20 a-onl End_of_Run_3216
  • 354077 04/01/11 20:39 Sarah Screen capture of the neutron detector single rates during run 3216
  • 354076 04/01/11 20:36 Sarah Screen shots of the bigbite and scaler data screens during run 3216
  • 354075 04/01/11 20:36 auto_snap Radiation and Ozone snapshot
  • 354074 04/01/11 20:36 auto_snap Miscellaneous snapshot
  • 354073 04/01/11 20:36 auto_snap Hall A Tools snapshot
  • 354072 04/01/11 20:36 auto_snap Compton Polarimeter snapshot
  • 354071 04/01/11 20:36 auto_snap BigBox Magnet snapshot
  • 354070 04/01/11 20:35 auto_snap Beam Position and Current snapshot
  • 354069 04/01/11 20:00 a-onl Start_Run_3216,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354068 04/01/11 19:59 a-onl End_of_Run_3215
  • 354067 04/01/11 19:38 BeckyW My First Day
  • 354066 04/01/11 18:59 orchen Improved (?) HRS coincedence for 625MeV/c and 750MeV/c kinematics
  • 354065 04/01/11 18:45 a-onl Start_Run_3215,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354064 04/01/11 18:44 a-onl End_of_Run_3214
  • 354063 04/01/11 17:59 Sarah Screen snapshot of neutron detector scaler rates during run 3214
  • 354062 04/01/11 17:58 Sarah Screen snapshots of the bigbite and scaler data screens during run 3214
  • 354061 04/01/11 17:28 a-onl Start_Run_3214,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354060 04/01/11 17:27 a-onl End_of_Run_3213
  • 354059 04/01/11 16:21 Sarah Screen snapshot of neutron detector single rates during run 3213
  • 354058 04/01/11 16:20 Sarah Screen snapshot of BigBite and scaler data during run 3213
  • 354057 04/01/11 16:18 Fatiha Hall-A General tool screen
  • 354056 04/01/11 16:15 Sarah Day shift run list (analysis)
  • 354055 04/01/11 16:11 a-onl Start_Run_3213,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354054 04/01/11 16:10 a-onl End_of_Run_3212
  • 354053 04/01/11 16:02 adaq Beam Time Accounting
  • 354052 04/01/11 15:59 Leckey Day Shift Summary
  • 354051 04/01/11 15:50 V. Sulkosky Pile-up Effect in BB TOF @ 625 MeV/c
  • 354050 04/01/11 15:47 orchen BB and HAND (e,e'N_recoil) for the 750MeV/c and 625MeV/c kinematics
  • 354049 04/01/11 15:42 Higinbotham Right Arm at Lower Momentum and Smaller Angle
  • 354048 04/01/11 15:37 wenbiao Right HRS dipole problem Run Information
  • 354047 04/01/11 15:30 Higinbotham Right HRS set current to get 410.77A readback
  • 354046 04/01/11 15:03 eip HRS coinc background
  • 354045 04/01/11 15:02 Leckey Right Arm Dipole Stable
  • 354044 04/01/11 14:51 orchen HRS coincedence for 625MeV/c and 750MeV/c kinematics
  • 354043 04/01/11 14:50 a-onl Start_Run_3212,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Right arm dipole now at correct momentum, regulation is now off,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354042 04/01/11 14:48 Higinbotham Right HRS Dipole Problems Runs #3204 through #3210
  • 354041 04/01/11 14:48 a-onl End_of_Run_3211
  • 354040 04/01/11 14:07 LeRose Left NMR
  • 354039 04/01/11 13:52 a-onl Start_Run_3211,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production 4He 20 cm Target, 4 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354038 04/01/11 13:51 a-onl End_of_Run_3210
  • 354037 04/01/11 13:30 orchen Pmiss for the new 750MeV/c kinematics
  • 354036 04/01/11 13:09 Leckey Gas Shed Airflow Alarm
  • 354035 04/01/11 12:35 auto_snap Radiation and Ozone snapshot
  • 354034 04/01/11 12:35 auto_snap Miscellaneous snapshot
  • 354033 04/01/11 12:35 auto_snap Hall A Tools snapshot
  • 354032 04/01/11 12:35 auto_snap Compton Polarimeter snapshot
  • 354031 04/01/11 12:35 auto_snap BigBox Magnet snapshot
  • 354030 04/01/11 12:35 auto_snap Beam Position and Current snapshot
  • 354029 04/01/11 12:35 Mitra S Screen shots of BigBite and Scaler data for run 3210
  • 354028 04/01/11 12:31 Mitra S Screen shot of neutron detector single rates for run 3210
  • 354027 04/01/11 12:23 a-onl Start_Run_3210,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production 4He 20 cm Target, 4 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354026 04/01/11 12:19 a-onl End_of_Run_3209
  • 354025 04/01/11 12:06 Mitra S Screen shot of neutron detector single rate for run 3209
  • 354024 04/01/11 12:02 Mitra S Screen shots of BigBite and Scaler data at the start of run 3209
  • 354023 04/01/11 11:57 Leckey Right Arm Dipole Ramping
  • 354022 04/01/11 11:54 Leckey Right Arm Dipole Flat at 50A for 60mins
  • 354021 04/01/11 11:46 a-onl Start_Run_3209,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production 4He 20 cm Target, 4 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354020 04/01/11 11:45 a-onl End_of_Run_3208
  • 354019 04/01/11 11:45 a-onl Start_Run_3208,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production 4He 20 cm Target, 4 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354018 04/01/11 11:35 a-onl End_of_Run_3206
  • 354017 04/01/11 11:10 Leckey Raster Pics
  • 354016 04/01/11 11:06 Leckey Right Arm Dipole Current
  • 354015 04/01/11 10:43 Jack (Techs) Right Dipole current setting issue
  • 354014 04/01/11 10:37 wenbiao target status
  • 354013 04/01/11 09:55 Mitra S Screen shots of BigBite and Scalaer data during run 3206
  • 354012 04/01/11 09:51 Mitra S Screen shots of neutron detector
  • 354011 04/01/11 09:50 Leckey BBPS1:2001:Module L10.2 Trip
  • 354010 04/01/11 09:46 Mitra S No acknowledge box at the begining of the run
  • 354009 04/01/11 09:40 a-onl Start_Run_3206,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production 4He 20 cm Target, 4 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354008 04/01/11 09:16 Mitra S CODA crashed
  • 354007 04/01/11 09:14 Leckey Beam Away for 20 mins
  • 354006 04/01/11 08:57 Leckey Right Arm Dipole Problem
  • 354005 04/01/11 08:56 a-onl Start_Run_3205,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production 4He 20 cm Target, 4 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354004 04/01/11 08:53 a-onl End_of_Run_3204
  • 354003 04/01/11 08:32 Mitra s Screen shots of neutron detector single rate
  • 354002 04/01/11 08:28 Mitra S Screen shots of BigBite and scaler data during run 3204
  • 354001 04/01/11 08:14 a-onl Start_Run_3204,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production 4He 20 cm Target, 4 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 354000 04/01/11 08:12 a-onl End_of_Run_3203
  • 353999 04/01/11 08:03 adaq Beam Time Accounting
  • 353998 04/01/11 08:00 arie shift summary
  • 353997 04/01/11 07:57 D.Wang target status
  • 353996 04/01/11 06:56 a-onl Start_Run_3203,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production 4He 20 cm Target, 4 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353995 04/01/11 06:55 a-onl End_of_Run_3202
  • 353994 04/01/11 06:27 D.Wang target status
  • 353993 04/01/11 06:07 a-onl Start_Run_3202,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production 4He 20 cm Target, 4 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353992 04/01/11 06:06 a-onl End_of_Run_3201
  • 353991 04/01/11 04:52 a-onl Start_Run_3201,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production 4He 20 cm Target, 4 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353990 04/01/11 04:51 a-onl End_of_Run_3200
  • 353989 04/01/11 04:35 auto_snap Radiation and Ozone snapshot
  • 353988 04/01/11 04:35 auto_snap Miscellaneous snapshot
  • 353987 04/01/11 04:35 auto_snap Hall A Tools snapshot
  • 353986 04/01/11 04:35 auto_snap Compton Polarimeter snapshot
  • 353985 04/01/11 04:35 auto_snap BigBox Magnet snapshot
  • 353984 04/01/11 04:35 auto_snap Beam Position and Current snapshot
  • 353983 04/01/11 04:23 arie general tools
  • 353982 04/01/11 04:19 arie spot_R
  • 353981 04/01/11 04:04 D.Wang target status
  • 353980 04/01/11 03:31 a-onl Start_Run_3200,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production 4He 20 cm Target, 4 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353979 04/01/11 03:30 a-onl End_of_Run_3199
  • 353978 04/01/11 02:15 a-onl Start_Run_3199,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production 4He 20 cm Target, 4 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353977 04/01/11 02:14 a-onl End_of_Run_3198
  • 353976 04/01/11 02:06 D.Wang target status
  • 353975 04/01/11 01:01 a-onl Start_Run_3198,Run_type=Production SRC,target_type=4He 20cm cell,beam_e0=4.46,beam_raster=on,beam_current=4,leftarm_p=3.6014,leftarm_theta=20.3,leftarm_sieve=out,rightarm_p=1.19635,rightarm_theta=24.5,rightarm_sieve=out,comment_text=Production 4He 20 cm Target, 4 muA,bbite_current=518,bbite_theta=92,hand_angle=92,bbite_sieve=out,hand_lead=on
  • 353974 04/01/11 01:00 a-onl End_of_Run_3197
  • 353973 04/01/11 00:06 adaq Beam Time Accounting
  • 353972 04/01/11 00:05 Ramesh Shift Summary (Swing Shift)

  • Earlier months at hallaweb.jlab.org/adaq (archive).  The HALOG was generated by version HA1.48
    If you have any comments or problems, please contact:
    Robert Michaels
    rom@jlab.org
    (757) 269-7410